### v1

# On the volume of K-semistable Fano manifolds

Abstract.
We prove that the volume of an $n$-dimensional K-semistable Fano manifold that is not ${\mathbb{P}}^{n}$ is at most $2n^{n}$. Moreover, the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$. Our proof is based on a new connection between K-semistability and minimal rational curves. More generally, we show that the volume of a K-semistable Fano manifold with a minimal rational curve of degree $d$ is bounded above by the volume of ${\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1}$.
1. Introduction
The study of K-(poly)stable Fano varieties is a very active research area. Thanks to the resolution of the Yau-Tian-Donaldson conjecture, we know that a Fano variety admits a Kähler-Einstein metric (resp. a unique Kähler-Einstein metric) if and only if it is K-polystable (resp. K-stable). A Fano variety is K-semistable if and only if it degenerates (via a special test configuration) to a K-polystable Fano variety. There are well-established valuative criterion for K-(semi)stability ([Fuj19, Li17]) and powerful methods to test it (see [AZ22]).
In this paper, we study the anticanonical volume (or degree) of K-semistable Fano varieties. For a Fano variety $X$, its (anticanonical) volume ${\rm vol}(X)$ is defined to be the self-intersection number $(-K_{X})^{n}$. It is known that the volume of an $n-$dimensional Fano manifold can exceed the volume of ${\mathbb{P}}^{n}$, even among toric Fano manifolds (see [Deb01, Section 5.11]). In a major advance in the study of K-stability, Fujita [Fuj18] proved that the volume of an $n$-dimensional K-semistable Fano manifolds $X$ satisfies ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$ and the first equality holds if and only if $X\cong{\mathbb{P}}^{n}$. It was later generalized to possibly singular ${\mathbb{Q}}$-Fano varieties by Liu [Liu18]. The toric Fano case was proved earlier in [BB17]. Moreover, from the boundedness of K-semistable ${\mathbb{Q}}$-Fano varieties ([Jia20, Corollary 1.2]), we know that the set of volumes of $n$-dimensional K-semistable ${\mathbb{Q}}$-Fano varieties is finite away from zero. In this paper, we solve the conjecture on characterizing the second-largest volume for K-semistable Fano manifolds,
Theorem 1.1.
Any K-semistable Fano manifold $X$ that is not ${\mathbb{P}}^{n}$ satisfies ${\rm vol}(X)=(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$. In particular, this holds for any Kähler-Einstein metric manifold with positive Ricci curvature.
Like the result of Fujita, the last statement could be seen as a result in Kähler differential geometry. However, our proof uses purely algebraic geometry. The above result seems to be first conjectured in [AIM20, Problem 2.6], and is stated as a folklore conjecture in [Zhu25, Conjecture 6.8] which highlights the interesting (but also mysterious) feature that there are two Fano manifolds with second largest volume. Our proof will indeed give a satisfactory explanation of this feature by connecting it to the theory of minimal rational curves on Fano manifolds.
There is a local analogue of this conjecture sometimes called the ODP conjecture (see Conjecture 2.3) for the volume of klt singularity (see [SS17, LX19]). The ODP conjecture would imply that the volume of singular K-semistable Fano varieties is strictly less than $2n^{n}$ (by Liu’s local-to-global volume comparison, see Remark 3.9). Conversely, using Theorem 1.1, we immediately verify the ODP conjecture for the Fano cone over a K-semistable smooth Fano manifold (see Theorem 4.5).
The toric version of Theorem 1.1 has its own interest and is called the “gap hypothesis” in [AB24, Conjecture 3.10]. It implies the sharp bound of the canonical height of canonical model of toric Fano varieties over ${\rm Spec}\,{\mathbb{Z}}$ by [AB24, Lemma 3.8]. Moreover, in the toric case, the ODP conjecture is solved by Moraga-Süß’s result ([MS24]) using convex geometric methods. So combined with Theorem 1.1, we indeed have the sharp upper bound for all possibly singular K-semistable toric Fano varieties.
Theorem 1.2.
The second largest volume of $n$-dimensional K-semistable toric Fano varieties is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
We remark that Theorem 1.2 was proved in [AB24] for smooth toric Fano manifolds of dimension $n\leqslant 6$ by using the classification of [Obr07] and also for certain singular toric Fano varieties by convex geometric arguments (see [AB24, Lemma 3.9]). Furthermore, using the well-known one-to-one correspondence between Gorenstein toric Fano varieties and reflexive lattice polytopes in ${\mathbb{R}}^{n}$, Theorem 1.2 immediately implies a convex geometric statement.
Corollary 1.3.
Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with barycenter at $0\in{\mathbb{R}}^{n}$. Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $[0,2]\times(n\Delta_{n-1})$.
We will prove Theorem 1.1 by exploring a new connection between K-stability and minimal rational curves. Roughly speaking, we are going to estimate the volume by using the valuative criterion to test K-semistability via valuations associated to weighted blowups along minimal rational curves. An interesting new phenomenon we find is that the volume estimates become sharp when we let certain weights go to infinity, and we indeed get the following general result, which to our knowledge is the first result connecting explicitly the theory of K-stability and the theory of minimal rational curves on Fano manifolds.
Theorem 1.4.
Assume that a K-semistable Fano manifold $X$ admits a minimal rational curve $f:{\mathbb{P}}^{1}\rightarrow X$ such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ for $12\leq d\leq n+$ 1. Then we have ${\rm vol}(X)\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$.
The smooth toric case is special in the sense that there are always smoothly embedded projective spaces with trivial normal bundles (see [CFH14, Ara06]). So the toric Fano case can also be considered as an example of the following direct generalization of Fujita’s result.
Theorem 1.5.
Let $X$ be an $n$-dimensional K-semistable Fano manifold, and $Z\subset X$ a codimension $r$ smooth complex submanifold. Assume that the normal bundle of $Z$ inside $X$ is trivial and set $d=(-K_{X})^{n-r}\cdot Z$. Then we have the inequality
(1) $(-K_{X})^{n}\leqslant(r+1)^{r}\cdot\binom{n}{r}\cdot d=(-K_{{\mathbb{P}}^{r}\times Z})^{n}.$
Moreover, the equality holds if and only if $X$ is biholomorphic to ${\mathbb{P}}^{r}\times Z$. In particular $Z$ is also a K-semistable Fano manifold.
We sketch the organization of this paper. In the next section, we recall basic knowledge about some key concepts used in later sections: K-stability, Seshadri constant, and minimal rational curves. In section 3, we prove Theorem 1.5, and Theorem 1.1 in the case the Fano manifold contains a minimal rational curve with a trivial normal bundle. We also deduce Theorem 1.2 and Corollary 1.3. In section 4, we prove Theorem 1.4 to achieve the full proof of Theorem 1.1. We end the paper by giving some related examples and finding (K-semistable) Fano manifolds with minimal volumes.
Acknowledgments
C. Li is partially supported by NSF (Grant No. DMS-2305296). M. Miao would like to thank his advisor Gang Tian for his constant support, encouragement, and guidance. M. Miao would like to thank Jiyuan Han, Xiaowei Jiang, Yang Liu, Yalong Shi, and Linsheng Wang for helpful discussions. We thank Robert Berman, Kento Fujita, Yuchen Liu, Chenyang Xu and Ziquan Zhuang for their interest and helpful comments, and Kento Fujita for pointing to us the Fact 1. This work is carried out during M. Miao’s visit to Rutgers University. He would like to thank the graduate school of Nanjing University for providing funding support and thank the Department of Mathematics at Rutgers University for its hospitality.
2. Preliminaries
We will work over ${\mathbb{C}}$. Unless otherwise specified, all varieties are assumed to be normal and projective. A ${\mathbb{Q}}$-Fano variety $X$ is a variety with at worst klt singularities such that the anti-canonical divisor $-K_{X}$ is an ample ${\mathbb{Q}}$-Cartier divisor. A singularity $x\in X$ consists of a variety $X$ and a closed point $x\in X$. A singularity $x\in X$ is called klt if $X$ is klt near a neighborhood of $x$. A ${\mathbb{R}}$-valuation over singularity $x$ is a valuation $v\colon K(X)^{*}\rightarrow{\mathbb{R}}$ center at $x$ (namely, for all $f\in{\mathcal{O}}_{X,x}$, we have $v(f)\geqslant 0$ and $v(f)>0$ if and only if $f\in\mathfrak{m}_{x}$) and $v|_{{\mathbb{C}}^{*}}=0$. The set of all such valuations is denoted by ${\rm Val}_{X,x}$.
2.1. K-stability
K-stability, first introduced by Tian ([Tia97]) and later reformulated algebraically by Donaldson ([Don02]), is an algebro-geometric notion to characterize the existence of Kähler-Einstein metrics on Fano varieties. In this subsection, we recall some notions in K-stability theory that are relevant to our paper, and refer to [Xu24] for a detailed exposition of K-stability theory.
We say that a prime divisor $E$ is over $X$ if there exists a proper birational morphism $\mu\colon Y\rightarrow X$ such that $Y$ is normal and $E$ is a prime divisor on $Y$. We define the log discrepancy of the divisor $E$ over $X$ as
$A(E)=A_{X}(E)\coloneqq 1+{\rm coeff}_{E}(K_{Y}-\mu^{*}K_{X}).$
The volume of an ${\mathbb{R}}$-Cartier divisor $D$ is defined as ${\rm vol}_{X}(D)\coloneqq\mathop{{\rm lim\,sup}}_{m\rightarrow+\infty}\frac{h^{0}(X,{\mathcal{O}}_{X}(\lfloor mD\rfloor))}{m^{n}/n!}$. Note that the limsup is actually a limit and the volume function is continuous in the big cone $\text{Big}(X)$ ([Laz04, Section 2.2.C]). Define the $S$-invariant
$S(E):=S(-K_{X};E)\coloneqq\frac{1}{(-K_{X})^{n}}\int_{0}^{\infty}{\rm vol}(-K_{X})-xE)\,{\rm d}x,$
where, for the simplicity of notation, we just write ${\rm vol}(-K_{X}-xE)$ for ${\rm vol}(\mu^{*}(-K_{X})-xE)$. We have the following valuative criterion for K-semistability,
Theorem 2.1.
[Fuj19, Li17] A Fano variety $X$ is K-semistable if and only if $A(E)-S(E)\geqslant 0$ for every divisor $E$ over $X$.
Another equivalent way to characterize K-semistability is via the $\delta$-invariant (also known as stability threshold), we recall $\delta(X)\coloneqq{\rm inf}_{E/X}\frac{A(E)}{S(E)}$ where the infinum is taking over all divisors $E$ over $X$ ([FO18, BJ20]). Then $X$ is K-semistable if and only if $\delta(X)\geqslant 1$ by the valuative criterion.
One may define the log discrepancy $A_{X}(v)$ for any valuation $v\in{\rm Val}_{X,x}$, see [JM12, Section 5.1]. For a klt singularity $x\in(X,\Delta)$, we always have $A_{X}(v)>0$ for any valuation $v\in{\rm Val}_{X,x}$. Denote ${\rm Val}^{*}_{X,x}=\{v\in{\rm Val}_{X,x}\mid A_{(X,\Delta)}(v)<+\infty\}$. The volume of a valuation $v\in{\rm Val}_{X,x}$ is defined as
${\rm vol}_{X,x}(v)\coloneqq\mathop{{\rm lim\,sup}}_{m\rightarrow+\infty}\frac{l({\mathcal{O}}_{X,x}/\mathfrak{a}_{m}(v))}{m^{n}/n!},$
where $\mathfrak{a}_{m}(v)$ denotes the valuation ideals: $\mathfrak{a}_{m}(v)\coloneqq\{f\in{\mathcal{O}}_{X,x}\mid v(f)\geq m\}$. The first named author introduced the following invariant for singularity ([Li18]), which plays a key role in the study of local K-stability:
Definition 2.2.
Suppose $x\in X$ is a klt singularity, for a valuation $v\in{\rm Val}_{X,x}$, we define the normalized volume
${\widehat{\rm vol}}_{X}(v)\coloneqq\left\{\begin{aligned} &A_{X}(v)^{n}\cdot{\rm vol}(v)&\text{if}\ &A_{X}(v)<+\infty\\ &+\infty&\text{if}\ &A_{X}(v)=+\infty.\end{aligned}\right.$
The local volume of $x\in X$ is defined as ${\widehat{\rm vol}}(x,X)\coloneqq{\rm inf}_{v\in{\rm Val}_{X,x}}{\widehat{\rm vol}}_{X}(v)$.
It was proved that for any $n$-dimensional klt singularity $x\in X$, we have ${\widehat{\rm vol}}(x,X)\leqslant n^{n}$, with the equality holds if and only if $x\in X$ is a smooth point ([LX19]). The following conjecture is known as the ODP conjecture ([SS17, LX19]).
Conjecture 2.3.
(ODP conjecture) The second largest local volume of an $n$-dimensional klt singularity is $2(n-1)^{n}$, with equality if and only if $x\in X$ is an ordinary double point.
For later purposes, we formulate a calculus lemma and an estimate of volume. Let $\mathscr{F}$ denote the collection of continuous functions $\phi:[0,\infty)\rightarrow[0,\infty)$ that is piecewise smooth, strictly increasing and surjective. In particular, such a function $\phi$ satisfies
(2) $\mathop{{\rm lim}}_{x\rightarrow 0}\phi(x)=0,\quad\mathop{{\rm lim}}_{x\rightarrow+\infty}\phi(x)=+\infty.$
For any $\phi\in\mathscr{F}$, consider the following function of $V\in[0,\infty)$:
$F^{\phi}(V):=\frac{1}{V}\int_{0}^{\phi^{-1}(V)}\left(V-\phi(x)\right){\rm d}x.$
Lemma 2.4.
(1) For any $\phi\in\mathscr{F}$, the function $F^{\phi}(V)$ is also an element of $\mathscr{F}$.
(2) If $\psi\in\mathscr{F}$ satisfies $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ for any $V\in(0,\infty)$, then $F^{\phi}(V)\geq F^{\psi}(V)$ for any $V\in(0,\infty)$. If moreover the first strict inequality holds for $V\geq V_{1}$, then $F^{\phi}(V)>F^{\psi}(V)$ for $V\geq V_{1}$. As a consequence, we get $(F^{\phi})^{-1}(A)<(F^{\psi})^{-1}(A)$ for any $A\in(0,\infty)$ with strict inequality if $A>F^{\psi}(V_{1})$.
Proof.
First assume that $\phi(x)$ is smooth and strictly increasing on $[0,\infty)$. Then its inverse function $\phi^{-1}(x)$ is also smooth, strictly increasing and satisfies (2). Set $G(V)=\int_{0}^{\phi^{-1}(V)}(V-\phi(x)){\rm d}x$ so that $F(V)=F^{\phi}(V)=G(V)/V$. Then $G$ is a smooth function satisfying $G(0)=0$. Its derivative is equal to $G^{\prime}(V)=\int_{0}^{\phi^{-1}(V)}1\,{\rm d}x=\phi^{-1}(V)$. We can then calculate the derivative of $F(V)$ at any $V\in(0,\infty)$:
$F^{\prime}(V)=\frac{G^{\prime}\cdot V-G}{V^{2}}=\frac{\int_{0}^{\phi^{-1}(V)}\phi(x){\rm d}x}{V^{2}}=\frac{H(V)}{V^{2}}>0$
with $H(V)=H^{\phi}(V)=\int_{0}^{\phi^{-1}(V)}\phi(x){\rm d}x$. So we know that $F(V)$ is a strictly increasing function of $V\in[0,\infty)$. Using L’Hospital’s rule, we easily see that $F$ satisfies the limit conditions in (2). So $F\in\mathscr{F}$.
Note that $H(0)=0$ and $H^{\prime}(V)=\phi(\phi^{-1}(V))(\phi^{-1})^{\prime}(V)=V(\phi^{-1})^{\prime}(V)$. By integrating the assumed inequality twice, we get the inequality $F^{\phi}(V)\geq F^{\psi}(V)$. For the last statement, note that under the assumption of strict inequality, we know that if $A>F^{\psi}(V_{1})$ then $(F^{\psi})^{-1}(A)>V_{1}$ which implies: $F^{\phi}((F^{\psi})^{-1}(A))>F^{\psi}((F^{\psi})^{-1}(A))=A$. So we conclude that $(F^{\psi})^{-1}(A)>(F^{\phi})^{-1}(A)$ when $A>F^{\psi}(V_{1})$.
If $\phi$ is only piecewise smooth, we can just carry out the above argument piecewise on each interval of smoothness. The same argument applies to the second statement too. ∎
Corollary 2.5.
Assume $(X,L:=-K_{X})$ is a K-semistable ${\mathbb{Q}}$-Fano variety. Let $E$ be a divisor over $X$ that satisfies the estimate: there exists $\phi\in\mathscr{F}$ such that for any $x\in[0,\infty)$,
(3) ${\rm vol}(L-xE)\geqslant V-\phi(x).$
There exists a unique solution $T=T(A_{X}(E))$ to the equation
(4) $(T-A_{X}(E))\phi(T)=\Phi(T)$
where $\Phi(x)=\int_{0}^{x}\phi(t){\rm d}t$ is the primitive function of $\phi(x)$ with $\Phi(0)=0$. Moreover, we have an estimate of the volume:
(5) $V=(-K_{X})^{n}\leqslant\phi(T)=(F^{\phi})^{-1}(A(E)).$
The equality $V=\phi(T)$ holds if and only if the equality holds in (3) for $x\in[0,T]$. In this case, $T$ is the pseudo-effective threshold of the valuation ${\rm ord}_{E}$ and in particular ${\rm vol}(L-TE)=0$.
Proof.
We have an estimate:
[…]
$=$ $A_{X}(E)-F^{\phi}(V).$
By the above lemma, $F^{\phi}(V)$ is a strictly increasing function that diverges to +$\infty$+ ∞ as $V\rightarrow+\infty$. So there exists a unique $V^{*}$ that satisfies $A(E)=F^{\phi}(V^{*})$ which is equivalent to the equality:
$(\phi^{-1}(V^{*})-A(E))V^{*}=\int_{0}^{\phi^{-1}(V^{*})}\phi(x){\rm d}x.$
Setting $T=\phi^{-1}(V^{*})$, we see that $T$ satisfies $(T-A(E))\phi(T)=\Phi(T)$. ∎
Example 2.6 (Fujita [Fuj18]).
Assume that $\phi(x)=x^{n}$ and $A(E)=n$. The $\Phi(x)=\frac{x^{n+1}}{n+1}$. The equation (4) has the solution $T=n+1$ and $\phi(T)=(n+1)^{n}$.
Example 2.7.
Assume that $\phi(x)=2nx^{n-1}$ and $A(E)=n-1$. The $\Phi(x)=2x^{n}$. The equation (4) has the solution $T=n$ and $\phi(T)=2n^{n}$.
2.2. Seshadri Constant
Let $X$ be a normal projective variety and $L$ an ample ${\mathbb{Q}}$-Cartier divisor on $X$. Let $Z\subset X$ be a nonsingular closed subvariety of $X$ of codimension $r$. The Seshadri constant of $L$ at $Z$ is defined as $\epsilon(L,Z)\coloneqq{\rm sup}\{t\in{\mathbb{R}}_{>0}\mid\pi^{*}L-tE\,\text{\,is ample}\}$, where $\pi\colon{\rm Bl}_{Z}X\rightarrow X$ is the blowup of $X$ along $Z$ and $E\cong{\mathbb{P}}(N_{Z/X}^{\vee})$ is the exceptional divisor. When $X={\mathbb{P}}^{r}\times Z$ and $Z$ is identified with $\{p\}\times Z$ for a fixed $p\in{\mathbb{P}}^{r}$, $(\mathrm{Bl}_{Z}X,E)$ are the same as $((\mathrm{Bl}_{p}{\mathbb{P}}^{r})\times Z,E_{p}\times Z)$ where $E_{p}\cong{\mathbb{P}}^{r-1}$ is the exceptional divisor of the blow up $\mathrm{Bl}_{p}{\mathbb{P}}^{r}\rightarrow{\mathbb{P}}^{r}$. We see that $\epsilon(-K_{{\mathbb{P}}^{r}\times Z},Z)=r+1$.
The next proposition generalizes a result of [LZ18] to higher dimensional subvariety $Z$ with trivial normal bundle, which says there is a gap between $r$ and $r+1$ for all the possible value of $\epsilon(-K_{X},Z)$.
Proposition 2.8.
Suppose $X$ is a ${\mathbb{Q}}$-Fano variety of dimension $n$. If there exists a nonsingular subvariety $Z\subset X$ of codimension $r\geqslant 2$ with trivial normal bundle $N_{Z/X}={\mathcal{O}}_{Z}^{\oplus r}$ and $\epsilon(-K_{X},Z)>r$, then $X\cong{\mathbb{P}}^{r}\times Z$.
Proof.
We follow the argument as in [LZ18, Theorem 2]. For simplicity, we denote $\epsilon\coloneqq\epsilon(-K_{X},Z)>r$. Let $\pi\colon\hat{X}\coloneqq{\rm Bl}_{Z}X\rightarrow X$ be the blowup of $X$ along $Z$ and $E$ is the exceptional divisor. We denote $B\coloneqq\pi^{*}(-K_{X})-\epsilon E$. From the definition of $\epsilon(-K_{X},Z)$, we know $B$ is a nef divisor. We have $K_{\hat{X}}=\pi^{*}K_{X}+(r-1)E$, so $B-K_{\hat{X}}=2(\pi^{*}(-K_{X})-\frac{\epsilon+r-1}{2}E)$ is nef and big since $(\epsilon+r-1)/2<\epsilon$. Then by Kawamata’s basepoint-free theorem ([KM98, Theorem 3.3]), we know $B$ is indeed semiample. Then there exists a fibration $\Phi\colon\hat{X}\rightarrow Y\subseteq{\mathbb{P}}(H^{0}(\hat{X},kB))$ induced by the complete linear series $|kB|$ for some $k\gg 0$ and $Y$ is a closed subvariety. Next, let $m$ be an integer such that $mB$ is Cartier. Note that
$mB-E-K_{\hat{X}}=(m+1)\left(\pi^{*}(-K_{X})-\frac{m\epsilon+r}{m+1}E\right)$
is an ample divisor since $(m\epsilon+r)/(m+1)<\epsilon$. Then, by Kodaira’s vanishing theorem, we get $H^{1}(\hat{X},mB-E)=0$. Therefore, the natural map $H^{0}(\hat{X},mB)\rightarrow H^{0}(E,mB|_{E})$ is surjective for all $m>0$ such that $mB$ is Cartier. We conclude that $\Phi|_{E}\colon E\rightarrow Y$ is a closed embedding. Since $\hat{X}$ has klt singularites, then it has rational singularities ([KM98, Theorem 5.22]), so in particular it is Cohen-Macaulay. Since $B=-K_{\hat{X}}+(r-1-\epsilon)E\sim_{\Phi,{\mathbb{Q}}}0$, we get $-K_{\hat{X}}\sim_{\Phi,{\mathbb{Q}}}\lambda E$ where $\lambda-r+1+\epsilon>1$. By [LZ18, Lemma 8], we have $\Phi\colon\hat{X}\rightarrow Y$ is not birational. Thus, $\Phi$ must be a fiber type contraction.
Since we have already shown $\Phi|_{E}\colon E\rightarrow Y$ is a closed embedding, we conclude that $\Phi|_{E}\colon E\rightarrow Y$ is in fact an isomorphism, so $Y\cong E\cong{\mathbb{P}}(N_{Z/X}^{\vee})\cong Z\times{\mathbb{P}}^{r-1}$. The general fiber of $\Phi$ is a smooth rational curve. By a similar argument as in [LZ18, Lemma 6], we can show $\Phi\colon\hat{X}\rightarrow Y$ is a smooth ${\mathbb{P}}^{1}$-fibration. Note that $s=\Phi|_{E}^{-1}\colon Y\rightarrow E$ gives a section of $\Phi$, then there exists a rank 2 vector bundle ${\mathcal{E}}$ over $Y$ such that $\hat{X}={\mathbb{P}}_{Y}({\mathcal{E}})$. By [Har77, Proposition V.2.6], there exists an invertible sheaf ${\mathcal{L}}$ on Y and a surjective morphism ${\mathcal{E}}\rightarrow{\mathcal{L}}$. We denote ${\mathcal{K}}=\ker({\mathcal{E}}\rightarrow{\mathcal{L}})$. Then ${\mathcal{O}}_{Y}(-1)\cong s^{*}N_{E/\hat{X}}\cong{\mathcal{L}}\otimes{\mathcal{K}}^{-1}$. We know that the choice of ${\mathcal{E}}$ is unique up to twisting of an invertible sheaf. So we may assume ${\mathcal{K}}={\mathcal{O}}_{Y}$, then ${\mathcal{L}}={\mathcal{O}}_{Y}(-1)\cong p_{2}^{*}{\mathcal{O}}_{{\mathbb{P}}^{r-1}}(-1)$ with the projection $p_{2}\colon Y\rightarrow{\mathbb{P}}^{r-1}$ and we have the following short exact sequence
$0\rightarrow{\mathcal{O}}_{Y}\rightarrow{\mathcal{E}}\rightarrow{\mathcal{O}}_{Y}(-1)\rightarrow 0$
By [Har77, Prop III.6.3], we have
${\rm Ext}^{1}({\mathcal{L}},{\mathcal{O}}_{Y})=H^{1}(Y,{\mathcal{L}}^{\vee})=\bigoplus_{i+j=1}H^{i}({\mathbb{P}}^{r-1},{\mathcal{O}}_{{\mathbb{P}}^{r-1}}(1))\otimes H^{j}(Z,{\mathcal{O}}_{Z})=0,$
so the short exact sequence actually splits. Therefore, ${\mathcal{E}}\cong{\mathcal{O}}_{Y}\oplus{\mathcal{O}}_{Y}(-1)$ and $\hat{X}\cong{\mathbb{P}}({\mathcal{O}}_{Y}\oplus{\mathcal{O}}_{Y}(-1))\cong{\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{r-1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{r-1}}(-1))\times Z\cong{\rm Bl}_{p}{\mathbb{P}}^{r}\times Z$ is isomorphic to ${\mathbb{P}}^{r}\times Z$ blowup along $\{p\}\times Z$. So $X\cong{\mathbb{P}}^{r}\times Z$. ∎
2.3. Minimal rational curves
Since the celebrated work of Mori introducing the bend-and-break method for constructing rational curves on Fano manifolds ([Mor79]), the theory of rational curves has been developed extensively and has found numerous applications in algebraic geometry, especially in the classification of Fano manifolds with special properties. Here we recall some basic knowledge of rational curves and refer to [Kol96] for detailed expositions.
A free rational curve is represented by a morphism $f:{\mathbb{P}}^{1}\rightarrow X$ that satisfies $f^{*}TX=\oplus_{i=1}^{n}{\mathcal{O}}_{{\mathbb{P}}^{1}}(a_{i})$ with $a_{i}\geq 0$. A free rational curve is called minimal (or standard) if
$f^{*}TX={\mathcal{O}}_{{\mathbb{P}}^{1}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)^{d-2}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus(n-d+1)}.$
Such a morphism $f$ must be an immersion and its image $f({\mathbb{P}}^{1})$ has at worst nodal singularities. We will also be interested in its normal bundle:
$N_{f/X}:=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
Any Fano manifold always admits minimal rational curves, which can be obtained from the bend-and-break process starting with a free rational curve (see [Kol96, IV.Theorem 2.10]). Denote by $\mathrm{RatCurves}^{n}(X)$ the normalization of open subset of $\mathrm{Ch}(X)$ parametrizing integral rational curves. An irreducible component ${\mathcal{M}}$ of $\mathrm{RatCurves}^{n}(X)$ is referred to as a family of rational curves on $X$. The anticanonical degree $\deg({\mathcal{M}})$ of the family ${\mathcal{M}}$ is defined to be $-K_{X}\cdot C$ for any curve $C$ belonging to the family. This family ${\mathcal{M}}$ is equipped with a ${\mathbb{P}}^{1}$-bundle $p:{\mathcal{U}}\rightarrow{\mathcal{M}}$ and an evaluation morphism $q:{\mathcal{U}}\rightarrow X$. The family ${\mathcal{M}}$ is a dominating family if the evaluation morphism $q:{\mathcal{U}}\rightarrow X$ is dominant (i.e. has a dense image). This is equivalent to the condition that a general rational curve in ${\mathcal{M}}$ is free. A dominating family ${\mathcal{M}}$ is locally unsplit if, for a general point $x\in X$, the subfamily ${\mathcal{M}}_{x}=p(q^{-1}(x))$ parametrizing curves through $x$ is proper (i.e. compact). Note that ${\mathcal{M}}$ is a dominating family of rational curves on $X$ that has a minimal anticanonical degree, then ${\mathcal{M}}$ is a dominating locally unsplit family. But not all dominating locally unsplit family has a minimal anticanonical degree.
Let ${\mathcal{M}}$ be a dominating locally unsplit family of rational curves. By Mori’s bend-and-break argument, we know that for a general point $x\in X$, a general curve in ${\mathcal{M}}_{x}$ is minimal. In particular, $\deg({\mathcal{M}})\in\{2,3,\cdots,n+1\}$. Indeed, if a general free rational curve $C=[f]$ in ${\mathcal{M}}_{x}$ is not minimal, then there are at least two ${\mathcal{O}}(2)$ summands in the splitting of $f^{*}TX$. We can then fix two points on the curve $C$ and bend-and-break rational curve into a non-integral (reducible) curve which would contradict the locally unsplit property. Following [Miy04, CD15], we set
(6) $l_{X}:={\rm min}\{\deg{\mathcal{M}};{\mathcal{M}}\text{ is a dominating locally unsplit family of rational curves on }X\}.$
By the above discussion, it is easy to see that (see [CD15, Remark 4.2]):
(7) $l_{X}$ $=$ ${\rm min}\{-K_{X}\cdot C;C\subset X\text{ is a free rational curve on }X\}$
$=$ ${\rm min}\{-K_{X}\cdot C;C\subset X\text{ is a minimal rational curve on }X\}.$
Example 2.9.
To prove our main results, we will need the following important results on classification of Fano manifolds with dominating locally unsplit families of large anticanonical degrees.
  1. (1)
    ([CMSB02]) $l_{X}=n+1$ if and only if $X\cong{\mathbb{P}}^{n}$.
  2. (2)
    ([Miy04, CD15, DH17]) $l_{X}=n$ if and only if $X\cong Q^{n}$ or $X$ is the blowup of ${\mathbb{P}}^{n}$ along subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ that is contained in a hyperplane.
Example 2.10 ([HM03]).
On the other extreme, the condition $l_{X}=2$ is equivalent to the following condition:
  1. (1)
    There exists a minimal rational curve with trivial normal bundle.
  2. (2)
    For a general point $x\in X$, there are only finitely many rational curves through $x$ which have minimal degree with respect to $K_{X}^{-1}$.
  3. (3)
    for a general point $x\in X$, there exists a rational curve which has degree 2 with respect to $K_{X}^{-1}$.
There are many examples of such Fano manifolds. For example, except for the projective space ${\mathbb{P}}^{3}$ and 3-dimensional hyperquadric $Q^{3}$, all Fano 3-folds of Picard number 1 satisfy this condition. Hypersurfaces of ${\mathbb{P}}^{n+1}$ of degree $n$ or $n+1$ are also examples of such Fano manifolds when $n\geqslant 3$.
3. Warm-up: case of trivial normal bundles
In this section, we prove the following theorem, which is a special case of Theorem 1.1.
Theorem 3.1.
Let $X$ be a K-semistable Fano manifold that contains a rational curve with a trivial normal bundle. Then $(-K_{X})^{n}\leqslant 2n^{n}$ and the equality holds if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
The rest of this section is devoted to the proof of Theorem 3.1. We start with a proposition that proves a more general version of estimate (1) by incorporating the $\delta$-invariant.
Proposition 3.2.
Assume $X$ is a $n$-dimensional Fano manifold, and $Z\subset X$ is a codimension $r$ non-singular subvariety with trivial normal bundle $N_{Z/X}={\mathcal{O}}_{Z}^{\oplus r}$, and we set $d=(-K_{X})^{n-r}\cdot Z$, then
$(-K_{X})^{n}\leqslant\delta(X)^{-r}\cdot(r+1)^{r}\binom{n}{r}\cdot d=\delta(X)^{-r}\cdot(-K_{{\mathbb{P}}^{r}\times Z})^{n}.$
Proof.
Let $\pi\colon\hat{X}\coloneqq{\rm Bl}_{Z}X\rightarrow X$ be the blowup of $X$ along $Z$ with the exceptional divisor $E$. First, it is clear that the log discrepancy
$A_{X}(E)=1+{\rm coeff}_{E}(K_{\hat{X}}-\pi^{*}K_{X})=1+(r-1)=r.$
For the simplicity of notation, set $L=-K_{X}$. We can assume $x\in{\mathbb{Q}}$ since the volume function ${\rm vol}_{\hat{X}}(\pi^{*}L-xE)$ is continuous. Then we take $k\in{\mathbb{N}}^{*}$ sufficiently large such that $kx\in{\mathbb{Z}}_{>0}$. Note that we have the exact sequence:
$0\rightarrow H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{xk})\rightarrow H^{0}(X,kL)\rightarrow H^{0}(X,kL\otimes{\mathcal{O}}_{xkZ})\rightarrow\cdots,$
which implies:
$h^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{xk})\geqslant h^{0}(X,kL)-h^{0}(X,kL\otimes{\mathcal{O}}_{xkZ}).$
Note that the higher direct images $R^{i}\pi_{*}{\mathcal{O}}_{\hat{X}}=0$ for $i>0$. By the Leray spectral sequences, we get $H^{0}(\hat{X},\pi^{*}(kL)-xkE)=H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{xk})$ and $H^{0}(\hat{X},\pi^{*}(kL))=H^{0}(X,kL)$. Thus,
${\rm vol}_{\hat{X}}(\pi^{*}L-xE)=\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{xk})}{k^{n}/n!}\geqslant L^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,kL\otimes{\mathcal{O}}_{xkZ})}{k^{n}/n!}.$
For $j\in{\mathbb{Z}}_{\geqslant 0}$, we use the exact sequence
$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0$
And $H^{1}(X,kL\otimes({\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}))=\oplus H^{1}(Z,-kK_{Z})=0$ from Kodaira vanishing and the assumption that normal bundle $N_{Z/X}$ is trivial. It is easy to prove by induction that:
(8) $h^{0}(X,kL\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{xk})$ $=$ $h^{0}(X,kL\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{xk-1})+h^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{xk-1}/{\mathcal{I}}^{xk}_{Z})$
$=$ $\sum_{j=0}^{xk-1}h^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}^{j+1}_{Z}).$
Using the assumption that $N_{Z/X}$ is trivial and $Z$ is non-singular, by [Har77, Thm II. 8.24], we get:
${\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}$$\cong$ $\mathrm{Sym}^{j}({\mathcal{I}}_{Z}/{\mathcal{I}}_{Z}^{2})=\mathrm{Sym}^{j}(N_{Z/X}^{\vee})={\mathcal{O}}_{Z}^{\oplus\binom{r-1+j}{r-1}}.$
Then the right-hand-side of (8) is given by
$\sum_{j=0}^{xk-1}h^{0}(Z,{\mathcal{O}}_{X}(kL)\otimes{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}^{j+1}_{Z})=\sum_{j=0}^{xk-1}\binom{j+r-1}{r-1}\cdot h^{0}(Z,-kK_{Z})=\binom{r+xk-1}{r}\cdot h^{0}(Z,-kK_{Z}).$
Then,
(9) $\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,kL)\otimes{\mathcal{O}}_{xkZ}}{k^{n}/n!}$ $=$ $\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{n!}{r!(n-r)!}\cdot\frac{(r+xk-1)!}{(xk-1)!\cdot k^{r}}\cdot\frac{h^{0}(Z,-kK_{Z})}{k^{n-r}/(n-r)!}$
$=$ $\binom{n}{r}(-K_{Z})^{n-r}\cdot x^{r}.$
We set $d=(-K_{Z})^{n-r}=(-K_{X})^{n-r}\cdot Z$. So we get ${\rm vol}_{\hat{X}}(\pi^{*}L-xE)\geqslant L^{n}-d\binom{n}{r}x^{r}$. Since $E\cong Z\times{\mathbb{P}}^{r-1}$ and ${\mathcal{O}}(-E)|_{E}=p_{2}^{*}{\mathcal{O}}_{{\mathbb{P}}^{r-1}}(1)$, one can easily see that the right-hand side equals to the top-intersection number
$(\pi^{*}L-xE)^{n}$ $=$ $\sum_{k=0}^{n}(\pi^{*}L)^{n-k}\cdot x^{k}(-E)^{k}$
$=$ $(\pi^{*}L)^{n}+(-1)^{r}\binom{n}{r}dx^{r}\cdot(-1)^{r-1}=L^{n}-d\binom{n}{r}x^{r}.$
Moreover, we have:
$S(-K_{X};E)$ $=$ $\frac{1}{(-K_{X})^{n}}\int_{0}^{T_{X}(E)}{\rm vol}_{\hat{X}}(\pi^{*}(-K_{X})-xE)\,{\rm d}x$
$\geqslant$ $\frac{1}{(-K_{X})^{n}}\int_{0}^{\epsilon}\left((-K_{X})^{n}-d\binom{n}{r}x^{r}\right)\,{\rm d}x$
$=$ $\epsilon-\frac{1}{(-K_{X})^{n}}\binom{n}{r}\frac{d\cdot\epsilon^{r+1}}{r+1}=\frac{r}{r+1}\epsilon,$
where $\epsilon=\left((-K_{X})^{n}/(\binom{n}{r}\cdot d)\right)^{1/r}$. Then,
$\delta(X)$$\leqslant$ $\frac{A_{X}(E)}{S(-K_{X};E)}\leqslant\frac{r}{\frac{r}{r+1}\epsilon}=\frac{r+1}{\left((-K_{X})^{n}/(\binom{n}{r}\cdot d)\right)^{1/r}}.$
Therefore, $(-K_{X})^{n}\leqslant\delta(X)^{-r}(r+1)^{r}\cdot\binom{n}{r}\cdot d=\delta(X)^{-r}\cdot(-K_{{\mathbb{P}}^{r}\times Z})^{n}$. ∎
Lemma 3.3.
Under the same notation as the above proposition, we set
(10) $\Lambda_{Z}(L)\coloneqq\{x\in{\mathbb{R}}_{\geqslant 0}\mid{\rm vol}_{\hat{X}}(\pi^{*}L-xE)=(\pi^{*}L-xE)^{n}\}.$
Then we have $\epsilon(L,Z)={\rm max}\{t\in{\mathbb{R}}_{\geqslant 0}\mid x\in\Lambda_{Z}(L)\,\,\text{for all}\,x\in[0,t]\}$.
Proof.
The argument is similar with [Fuj18, Theorem 2.3(2)]. We denote
(11) $\gamma=\gamma_{Z}(L)\coloneqq{\rm max}\{t\in{\mathbb{R}}_{\geqslant 0}\mid x\in\Lambda_{Z}(L)\,\,\text{for all}\,x\in[0,t]\}.$
When $0\leqslant t\leqslant\epsilon(L,Z)$, we have $\pi^{*}L-tE$ is nef, then ${\rm vol}_{\hat{X}}(\pi^{*}L-tE)=(\pi^{*}L-tE)^{n}$. Therefore, $\epsilon(L,Z)\leqslant\gamma$. In particular, $\gamma>0$. Now, in order to show $\epsilon(L,Z)\geqslant\gamma$, it suffices to show for any $\eta>0$ sufficiently small such that $\gamma-\eta\in{\mathbb{Q}}_{>0}$ and $\pi^{*}L-(\gamma-\eta)E$ is ample. Fix an $\delta\in{\mathbb{Q}}_{>0}$ such that $\pi^{*}L-\delta E$ is ample. Take any $t\in{\mathbb{Q}}_{>0}$ satisfying $t\leqslant{\rm min}\{1,(\gamma-\eta)/\delta,\,\eta/(\gamma-\delta)\}$. Then,
$\left(\pi^{*}L-(\gamma-\eta)E\right)-t(\pi^{*}L-\delta E)=(1-t)\left(\pi^{*}L-\frac{\gamma-\eta-t\delta}{1-t}E\right).$
We set $x_{t}\coloneqq(\gamma-\eta-t\delta)/(1-t)$. We take sufficiently large $k\in{\mathbb{N}}^{*}$ such that $kx_{t}\in{\mathbb{Z}}_{>0}$. From Kodaira’s vanishing theorem, we have $H^{i}(X,kL)=0$ for $i\geqslant 1$. Then the exact sequence is given by
$0\rightarrow H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow H^{0}(X,kL)\rightarrow H^{0}(X,kL\otimes{\mathcal{O}}_{kx_{t}Z})\rightarrow H^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow 0$
And $H^{i}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})=H^{i-1}(X,kL\otimes({\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{kx_{t}}))$ for $i\geqslant 2$. On the one hand, since $x_{t}\in\Lambda_{Z}(L)$ and by the equation (9),
$\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})}{k^{n}/n!}=-L^{n}+{\rm vol}_{\hat{X}}(\pi^{*}L-x_{t}E)+\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,kL\otimes{\mathcal{O}}_{kx_{t}Z})}{k^{n}/n!}=0.$
Then, $h^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})=o(k^{n})$. On the other hand, for $j\in{\mathbb{Z}}_{\geqslant 0}$, by the exact sequence
$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0$
We have $H^{i}(X,kL\otimes{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1})=0$ for $i\geqslant 1$ and $j\geqslant 1$ since $N_{Z/X}$ is trivial. So we get
$H^{i}(X,kL\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1})\cong H^{i}(X,kL\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j})$
for all $i\geqslant 1$ and $j\geqslant 1$. Then, for $i\geqslant 2$,
$h^{i}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})=h^{i-1}(X,kL\otimes({\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{kx_{t}}))=h^{i-1}(X,kL\otimes({\mathcal{O}}_{X}/{\mathcal{I}}_{Z}))=0.$
And the higher direct images $R^{i}\pi_{*}{\mathcal{O}}_{\hat{X}}=0$ for $i>0$, then by the Leray spectral sequence, we have $H^{i}(\hat{X},\pi^{*}(kL)-kx_{t}E)\cong H^{i}(X,kL\otimes{\mathcal{I}}_{Z}^{xk_{t}})$ for $i\geqslant 0$. In particular, we get $h^{i}(\hat{X},kL-kx_{t}E)=o(k^{n})$ for $i\geqslant 1$. Then by [dFKL07, Theorem A], we conclude that $\pi^{*}L-(\gamma-\eta)E$ is ample. This shows $\epsilon(L,Z)\geqslant\gamma$. ∎
Proof of Theorem 1.5.
Since $X$ is assumed to be K-semistable, we know that $\delta(X)\geq 1$ by the valuative criterion (Theorem 2.1). Without loss of generality, we assume the codimension $r$ of $Z$ in $X$ is positive. By the proof of Proposition 3.2, the equality (1) holds if and only if $\delta(X)=1$ and ${\rm vol}_{\hat{X}}(\pi^{*}(-K_{X})-xE)=(\pi^{*}(-K_{X})-xE)^{n}$ for all $x\in[0,r+1]$, so $\gamma_{Z}(-K_{X})=r+1$ (see (11)). Then by Lemma 3.3, we have $\epsilon(-K_{X},Z)=\gamma_{Z}(-K_{X})=r+1$. By Proposition 2.8, we get $X\cong Z\times{\mathbb{P}}^{r}$. ∎
Theorem 3.4.
The second largest volume of $n$-dimensional K-semistable toric Fano manifold is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
Proof.
When $X$ is a smooth toric Fano manifold, by the work of Chen-Fu-Hwang [CFH14, Corollary 2.5] which is partly based on the work of Araujo [Ara06], we know that there exists submanifold $Z\cong{\mathbb{P}}^{n-r}$ that is contained in $X$ with trivial normal bundle. By Proposition 3.2 and $\delta(X)\geqslant 1$, we have $(-K_{X})^{n}\leqslant(-K_{{\mathbb{P}}^{r}\times Z})^{n}=(-K_{{\mathbb{P}}^{r}\times{\mathbb{P}}^{n-r}})^{n}=\binom{n}{r}\cdot(r+1)^{r}\cdot(n-r+1)^{n-r}=:c_{r}$. By the following lemma 3.5, for $12\leq r\leq n-$ 1, $c_{r}\leq 2n^{n}$ with equality holds if and only if $r=2$ or $r=n-1$ in which case $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
Lemma 3.5.
The sequence $c_{r}:=\binom{n}{r}(r+1)^{r}(n-r+1)^{n-r}$ with $0\leq r\leq n$ satisfies $c_{r}=c_{n-r}$ and $c_{0}=(n+1)^{n}>c_{1}=2n^{n}>c_{2}>\cdots>c_{\lfloor n/2\rfloor}$.
Proof.
It is clear that $c_{r}=c_{n-r}$. We calculate:
$\frac{c_{r+1}}{c_{r}}$ $=$ $\left(\frac{r+2}{r+1}\right)^{r+1}\cdot\left(\frac{n-r}{n-r+1}\right)^{n-r}=a_{r+1}/a_{n-r}$
[…]
Proposition 3.6.
Let $X$ be a K-semistable Fano manifold. Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersed rational curve with a non-empty subset of nodal points such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}^{n-1}$. Then we have an estimate of the volume $(-K_{X})^{n}<2n^{n}$. As a consequence, if there exists a minimal rational curve with a trivial normal bundle on $X$ and $(-K_{X})^{n}=2n^{n}$, then the minimal rational curve has no nodal points and $X$ must be isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
Proof.
The last statement follows from the strict volume inequality and Theorem 1.5. So we just need to prove the inequality $(-K_{X})^{n}<2n^{n}$. Denote the ideal sheaf of the nodal curve $C=f({\mathbb{P}}^{1})$ in $X$ by ${\mathcal{I}}={\mathcal{I}}_{C}$. Consider the blow up $\mu:\tilde{X}\rightarrow X$ along $C$. By using the same argument as in the proof of Proposition 3.2, we get the estimate:
${\rm vol}({\mu}^{*}(-K_{X})-xE)$$\geqslant$ $(-K_{X})^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,L^{k}\otimes{\mathcal{O}}_{kxC})}{k^{n}/n!}$
$\geqslant$ $(-K_{X})^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{\mathop{\sum}_{y=0}^{xk-1}h^{0}(C,kL\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})}{k^{n}/n!}.$
Since $C$ is a local complete intersection, its conormal sheaf $N^{\vee}_{C/X}={\mathcal{I}}/{\mathcal{I}}^{2}$ is locally free and ${\mathcal{I}}^{y}/{\mathcal{I}}^{y+1}={\rm Sym}^{y}(N_{C/X}^{\vee})$. So we get:
$h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})=h^{0}(C,L^{k}\otimes{\rm Sym}^{y}({\mathcal{I}}/{\mathcal{I}}^{2}))\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(dk)\otimes{\rm Sym}^{y}(f^{*}({\mathcal{I}}/{\mathcal{I}}^{2}))).$
There is an inclusion ${\mathcal{F}}:=f^{*}({\mathcal{I}}/{\mathcal{I}}^{2})\hookrightarrow N_{f/X}^{\vee}=(f^{*}TX/T{\mathbb{P}}^{1})^{\vee}$. We know that $\deg(N^{\vee}_{f/X})=2-d$ and $\deg({\mathcal{F}})=\deg(K_{X})-\deg\omega_{C}=-d-2m+2$ where $d=-K_{X}\cdot C$. Since $N^{\vee}_{f/X}$ is trivial, $d=2$ and ${\mathcal{F}}={\mathcal{O}}(-a_{1})\oplus{\mathcal{O}}(-a_{2})\oplus\cdots\oplus{\mathcal{O}}(-a_{n-1})$ with $a_{1}\geq a_{2}\geq\cdots\geq a_{n-1}\geq 0$ and $\sum_{i=1}^{n-1}a_{i}=2m$. It is then easy to get the estimate:
(12) $h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-2)\oplus{\mathcal{O}}^{\oplus(n-2)}))$
or
(13) $h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-1)^{\oplus 2}\oplus{\mathcal{O}}^{\oplus(n-3)})).$
For the first case (12), we get:
(14) $\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leq\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-2i+1)a_{y,i}$
where $a_{y,i}=\binom{n-3+y-i}{n-3}$. If $x\leqslant 1$, then by using the explicit expression of binomial coefficient, the right-hand-side of (14) is equal to $b_{1}k^{n}/n!+O(k^{n-1})$ where $b_{1}$ is given by the following integral (see (19) for a derivation of a more complicated formula):
$b_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-2z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z=\frac{2n!}{(n-2)!}\int_{0}^{x}(1-z)(x-z)^{n-2}{\rm d}z$
$=$ $2nx^{n-1}-2x^{n}<2nx^{n-1}.$
When $x>1$, the right-hand-side of (14) splits into the sum: $\sum_{y=0}^{k}\sum_{i=0}^{y}+\sum_{y=k+1}^{xk-1}\sum_{i=0}^{k}$ and the right-hand-side of (14) is equal to $b_{2}k^{n}/n!+O(k^{n-1})$ with
$b_{2}$ $=$ $(2n-2)+n!\int_{1}^{x}{\rm d}t\int_{0}^{1}(2-2z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z$
$=$ $(2n-2)+\frac{2n!}{(n-2)!}\int_{0}^{1}(1-z)((x-z)^{n-2}-(1-z)^{n-2}){\rm d}z$
$=$ $\frac{2n!}{(n-2)!}\int_{0}^{1}u(x-1+u)^{n-2}{\rm d}u+(2n-2)-(2n-2)$
$=$ $2\displaystyle 2(nx^{n-1}-x^{n}+(x-1)^{n})<2(n-1)x^{n-1}+$ 2 .
We set $\psi=2nx^{n-1}$ and
(15) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 1\\ 2(n-1)x^{n-1}+2&\text{if }\;x>1,\end{array}\right.$
so that $\phi(x)<\psi(x)$. It is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and the strict inequality holds when $V>2n$. So we can apply Lemma 2.4 to conclude that $F^{\phi}(V)\geq F^{\psi}(V)$ and the strict inequality holds for $V>2n$. From Lemma 2.4 and Corollary 2.5, we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$.
In the second case (13), we get:
$\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leq\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-i+1)(i+1)\binom{n-4+y-i}{n-4}.$
Estimating as before, when $x\leqslant 2$, the right-hand-side of (14) is equal to $c_{1}k^{n}/n!+O(k^{n-1})$. The coefficient $c_{1}$ is equal to:
$c_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-z)z\frac{(t-z)^{n-4}}{(n-4)!}{\rm d}z$
$=$ $\frac{n!}{(n-3)!}\int_{0}^{x}(2-z)z(x-z)^{n-3}{\rm d}z=2x^{n-1}(n-x)<2nx^{n-1}.$
When $x\geqslant 2$, the right-hand-side of (14) splits into the sum: $\sum_{y=0}^{2k}\sum_{i=0}^{y}+\sum_{y=2k+1}^{xk-1}\sum_{i=0}^{2k}$ and the right-hand-side of (14) is equal to $c_{2}k^{n}/n!+O(k^{n-1})$ with the coefficient $c_{2}$ is equal to:
$c_{2}$ $=$ $2^{n}(n-2)+n!\int_{2}^{x}{\rm d}t\int_{0}^{2}(2-z)z\frac{(t-z)^{n-4}}{(n-4)!}{\rm d}z$
$=$ $2^{n}(n-2)+\frac{n!}{(n-3)!}\int_{0}^{2}(2-z)z((x-z)^{n-3}-(2-z)^{n-3}){\rm d}z$
$=$ $\frac{n!}{(n-3)!}\int_{0}^{2}(2-z)z(x-z)^{n-3}{\rm d}z$
$=$ $2nx^{n-1}-2x^{n}+(x-2)^{n}+n(x-2)^{n-1}<(2n-1)x^{n-1}$
$<$ $(2n-1)x^{n-1}+2^{n-1}.$
We set $\psi(x)=2nx^{n-1}$ and
(16) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 2\\ (2n-1)x^{n-1}+2^{n-1}&\text{if }\;x\geq 2.\end{array}\right.$
It is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and strict inequality holds for $V>2^{n}\cdot n$. Then by Lemma 2.4 we get $F^{\phi}(V)\geq F^{\psi}(V)$ with strict inequality when $V>2^{n}\cdot n$. So we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$. ∎
Proof of Theorem 3.1.
If $X$ contains a rational curve with non-empty nodal points, then by Proposition 3.6, we have $(-K_{X})^{n}<2n^{n}$. So it suffices to consider the case when $C$ is an embedded rational curve, in which case the result follows immediately from Theorem 1.5. ∎
Proposition 3.7.
Suppose $X$ is a singular $n$-dimensional K-semistable toric ${\mathbb{Q}}$-Fano variety. When $n=2$, then $(-K_{X})^{2}\leqslant\frac{9}{2}$. When $n\geqslant 3$, then $(-K_{X})^{n}\leqslant\frac{16}{27}(n+1)^{n}$. In particular, we have $(-K_{X})^{n}<(-K_{{\mathbb{P}}^{1}\times{\mathbb{P}}^{n}})^{n}=2n^{n}$ holds for all positive integer $n\geqslant 2$.
Proof.
Suppose $x\in X$ lies in the singular locus and $(X,x)$ is a toric singularity. It was proved in [MS24, Theorem 2] that if $(X,x)$ is a $n$-dimensional ${\mathbb{Q}}$-Gorenstein toric singularity, then
  1. (1)
    If $n=2$, then ${\widehat{\rm vol}}(X,x)\leqslant 2$;
  2. (2)
    If $n\leqslant 3$, then ${\widehat{\rm vol}}(X,x)\leqslant\frac{16}{27}n^{n}$.
By Liu’s local-to-global volume comparison ([Liu18]), we have $(-K_{X})^{n}\leqslant(\frac{n+1}{n})^{n}\cdot{\widehat{\rm vol}}(X,x)$. Then when $n=2$, $(-K_{X})^{2}\leqslant(\frac{3}{2})^{2}\cdot 2=\frac{9}{2}<2\cdot 2^{2}=8=(-K_{{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}})^{2}$. When $n\geqslant 3$,
$(-K_{X})^{n}\leqslant(\frac{n+1}{n})^{n}\cdot\frac{16}{27}n^{n}=\frac{16}{27}\cdot(n+1)^{n}<2n^{n}=(-K_{{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}})^{n}.$
Theorem 3.8.
The second largest volume of $n$-dimensional K-semistable toric ${\mathbb{Q}}$-Fano varieties is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
Proof.
When $X$ is a smooth toric Fano manifold, this follows from Theorem 3.4. When $X$ is a strictly singular ${\mathbb{Q}}$-Fano toric variety, this follows from Proposition 3.7. When the equality $(-K_{X})^{n}=2n^{n}$ holds, from Proposition 3.7, we know $X$ must be nonsingular. Then by Theorem 3.4 again, we know $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
Remark 3.9.
We remark that the singular version of Theorem 1.1 (that is, consider all ${\mathbb{Q}}$-Fano varieties, not just smooth Fano manifolds) follows from ODP conjecture (Conjecture 2.3). Let $X$ be a K-semistable ${\mathbb{Q}}$-Fano variety and $x\in X$ be a singular point. Assume the ODP conjecture holds, by the same argument of Proposition 3.7, we have $(-K_{X})^{n}\leqslant(\frac{n+1}{n})^{n}\cdot{\widehat{\rm vol}}(X,x)\leqslant(\frac{n+1}{n})^{n}\cdot 2(n-1)^{n}=(\frac{n+1}{n})^{n}\cdot(\frac{n-1}{n})^{n}\cdot 2n^{n}<2n^{n}$. Since the inequality is strict, it is expected that the equality of Theorem 1.1 would hold only if $X$ is a smooth Fano manifold.
We conclude this section by a convex-geometric application.
Corollary 3.10.
Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with barycenter at $0\in{\mathbb{R}}^{n}$. Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $[0,2]\times(n\Delta_{n-1})$.
Proof.
Here $\Delta_{n}$ denote the standard $n$-dimensional simplex with vertices $0,e_{1},\cdots,e_{n}$ where $e_{i}$ are standard lattice basis of ${\mathbb{Z}}^{n}$ for $i=1,\cdots,n$. And two convex bodies $P$ and $Q$ in ${\mathbb{R}}^{n}$ are called unimodularly equivalent if there exists an affine lattice automorphism of ${\mathbb{Z}}^{n}$ mapping $P$ onto $Q$. Let $X_{P}$ be the projective toric variety associated to a reflexive polytope $P$. Then $X_{P}$ is a Gorenstein Fano variety ([CLS11, Theorem 8.3.4]). By the result of [WZ04, SZ12, BB13], we know that the barycenter of $P$ is zero if and only if $X_{P}$ is K-semistable. Since $P$ is not unimodularly equivalent to $(n+1)\Delta_{n}$, we have $X_{P}\ncong{\mathbb{P}}^{n}$, then Theorem 3.8 immediately implies $(-K_{X_{P}})^{n}\leqslant 2n^{n}$ and the equality holds only if $X_{P}\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. Then the reflexive polytope $P$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and $P\cong[0,2]\times(n\Delta_{n-1})$ up to unimodularly equivalent. ∎
4. Volume estimates and minimal rational curves
4.1. Proof of Theorem 1.4 and Theorem 1.1
The calculation in the previous section is based on the test of K-stability via standard blow-ups. In this section, we prove Theorem 1.4 and Theorem 1.1 by considering the test of K-stability via weighted blow-ups along (Zariski open sets of) minimal rational curves.
Proof of Theorem 1.4.
Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ be minimal rational curve, or equivalently an immersed rational curve that satisfies $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ with $12\leq d\leq n+$ 1 and $d=f^{*}(-K_{X})\cdot{\mathbb{P}}^{1}=(-K_{X})\cdot f({\mathbb{P}}^{1})$. The (relative) normal bundle has the splitting:
$N_{f/X}=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
Correspondingly, we have the splitting of the conormal bundle:
$N^{\vee}_{f/X}={\mathcal{O}}(-1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
Set $Z=f({\mathbb{P}}^{1})$ which is an irreducible curve with possible nodal singularities. Fix a regular point $x\in Z$, there exists a Zariski open neighborhood $U$ and coordinates $\{z_{i};i=1,\dots,n\}$ such that
  1. (1)
    $Z\cap U=\{z_{1}=\cdots=z_{n-1}=0\}$;
  2. (2)
    For any $21\leq i\leq d-$ 2, ${\rm d}z_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}$ over $Z\cap U$;
  3. (3)
    For any $d-1\leq j\leq n-1$, ${\rm d}z_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}$ over $Z\cap U$.
For any $\ell\in{\mathbb{Z}}_{>0}$, we can define a divisorial valuation $v={\rm ord}_{E}$: for any $f\in{\mathcal{O}}(U)$,
(17) $v(f)={\rm min}\{|I|+\ell|J|;f=\sum_{I,J}{a_{IJ}}z^{\prime I}z^{\prime\prime J},a_{IJ}=a_{IJ}(z_{n})\neq 0\}$
where $z^{\prime}=\{z_{1},\dots,z_{d-2}\}$ (resp. $z^{\prime\prime}=\{z_{d-1},\cdots,z_{n-1}\}$), and $z^{\prime I}=z_{1}^{i_{1}}\cdots z_{d-2}^{i_{d-2}}$ for $I=(i_{1},\dots,i_{d-2})\in{\mathbb{N}}^{d-2}$ with $|I|=i_{1}+\cdots+i_{d-2}$ (resp. $z^{\prime\prime J}=z_{d-1}^{j_{1}}\cdots z_{n-1}^{j_{n-d+1}}$ for $J=(j_{1},\cdots,j_{n-d+1})\in{\mathbb{N}}^{n-d+1}$ with $|J|=j_{1}+\cdots+j_{n-d+1}$).
When restricted to $U$, the valuation $v$ corresponds to the exceptional divisor $E$ of the weighted blowup of $X\cap U$ along $Z\cap U$ with weights $(\underbrace{1,\dots,1}_{d-2},\underbrace{\ell,\cdots,\ell}_{n-d+1})$ (See [QR12] for more a general set-up of weighted blowups). Set ${\mathcal{I}}_{x}={\mathcal{I}}_{x}({\rm ord}_{E})=\{f\in{\mathcal{O}}_{X};v(f)\geq x\}$. Then we have the exact sequence:
$0\rightarrow H^{0}(X,L^{k}\otimes{\mathcal{I}}_{xk})\rightarrow H^{0}(X,L^{k})\rightarrow H^{0}({X},L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk})\rightarrow\cdots,$
which implies:
${\rm vol}(L-xE)\geq L^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}({X},L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk})}{k^{n}/n!}.$
By using the exact sequence:
$0\rightarrow{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{y+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{y}\rightarrow 0,$
we get inductively the estimates:
(18) $h^{0}(X,L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk})$$\leq$ $\sum_{y=0}^{xk-1}h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1}).$
For any $[s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, we know that with respect to a local trivialization of $L^{k}$, the section $s$ is represented locally over $U$ by a holomorphic function with the following expression modulo terms with $|I|+\ell|J|>y$,
$s$ $=$ $\sum_{|I|+\ell|J|=y}a_{IJ}z^{\prime I}z^{\prime\prime J}=\sum_{k=0}^{\lfloor y/\ell\rfloor}\sum_{|I|=y-\ell k,|J|=k}a_{IJ}z^{\prime I}z^{\prime\prime J}$
$=$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor+m(\ell-1)\\ |I|=y-\ell\lfloor y/\ell\rfloor+m\ell\\ |J|=\lfloor y/\ell\rfloor-m\end{subarray}}a_{IJ}z^{\prime I}z^{\prime\prime J}.$
Assume that $W$ is another Zariski open set such that $W\cap Z$ is non-empty and has at most one nodal point. In particular, there are at most two irreducible components of $W\cap Z$. Fix one irreducible component $Z_{1}$ which is smooth, we can choose coordinates (after possibly shrinking $W$) such that
  1. (1)
    $W\cap Z_{1}=\{w_{1}=\cdots=w_{n-1}=0\}$
  2. (2)
    For any $21\leq i\leq d-$ 2, ${\rm d}w_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}$ over $W\cap Z_{1}$.
  3. (3)
    For any $d-1\leq j\leq n-1$, ${\rm d}w_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}$ over $W\cap Z_{1}$.
Near $U\cap W\cap Z$, the transition functions from $\mathbf{z}$ and $\mathbf{w}$ must be of the form:
$z_{k}=a_{k}(w_{n})w_{k}+R_{2},\quad 1\leq k\leq n-1$
where $a_{k}(w_{n})$ is a transition function of ${\mathcal{O}}_{{\mathbb{P}}^{1}}(-1)$ for $21\leq k\leq d-$ 2 and of ${\mathcal{O}}_{{\mathbb{P}}^{1}}$ for $d-1\leq k\leq n-1$, and $R_{2}$ consists of terms of degree at least 2 in $w_{1},\dots,w_{n-1}$. So near $W\cap Z_{1}$, we have $s=\sum_{|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor}b_{IJ}(w_{n})w^{\prime I}w^{\prime\prime J}+R$ where $R$ are terms of (unweighted) total degree in $(w^{\prime},w^{\prime\prime})$ strictly greater than $y-(\ell-1)\lfloor y/\ell\rfloor$. We can deal with the other possible component of $W\cap Z$ using the same argument. So we see that $s$ induces a well-defined section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor}({\mathcal{O}}^{\oplus(n-d+1)}))$. This defines a linear map of vector spaces:
$H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\rightarrow H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor}({\mathcal{O}}^{\oplus(n-d+1)})),$
whose kernel consists of elements $s\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ that has expansion ($m\geq 1$ now):
$s=\sum_{m=1}^{\lfloor y/\ell\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor+m(\ell-1)\\ |I|=y-\ell\lfloor y/\ell\rfloor+m\ell\\ |J|=\lfloor y/\ell\rfloor-m\end{subarray}}a_{IJ}z^{\prime I}z^{\prime\prime J},$
which by the same argument as above induces a section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor+\ell}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-1}({\mathcal{O}}^{\oplus(n-d+1)}))$. So we inductively get the estimate:
$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$$\leq$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor+m\ell}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)}))$
$\leq$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{\ell m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
For the last inequality, we used the facts that $y-\ell\lfloor y/\ell\rfloor\geq 0$ and ${\mathcal{O}}(-1)$ is negative. Then the right-hand-side of (18) is bounded above by
$I_{k}(x):=\sum_{y=0}^{xk-1}\sum_{m=0}^{\lfloor y/\ell\rfloor}a_{y,m}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd-\ell m)),$
where
$a_{y,m}=\binom{d-3+\ell m}{d-3}\binom{n-d+\lfloor y/\ell\rfloor-m}{n-d}.$
We will calculate the leading coefficient $b_{1}={\rm lim}_{k\rightarrow+\infty}I_{k}(x)/k^{n}$. First, we assume that $x\leq d$. Then we can expand:
$I_{k}(x)$ $=$ $\sum_{y=0}^{xk-1}\sum_{m=0}^{\lfloor y/\ell\rfloor}\frac{(\ell m)^{d-3}+O((\ell m)^{d-4})}{(d-3)!}\frac{(\lfloor y/\ell\rfloor-m)^{n-d}+O((\lfloor y/\ell\rfloor-m)^{n-d-1})}{(n-d)!}(kd-\ell m+1)$
$=$ $\left[\frac{k^{n}}{(d-3)!(n-d)!}\sum_{y=0}^{xk-1}\frac{1}{k}\sum_{m=0}^{\lfloor y/\ell\rfloor}\frac{1}{k}(\ell\frac{m}{k})^{d-3}(\ell^{-1}\frac{y}{k}-\frac{m}{k})^{n-d}(d-\ell\frac{m}{k})\right]+O(k^{n-1}).$
As $k\rightarrow+\infty$, we see that $I_{k}(x)=b_{1}k^{n}/n!+O(k^{n-1})$ with $b_{1}$ given by the integral:
(19) $\frac{n!}{(d-3)!(n-d)!}\int_{0}^{x}{\rm d}t\int_{0}^{t/\ell}(\ell s)^{d-3}(\ell^{-1}t-s)^{n-d}(d-\ell s){\rm d}s$
$=$ $\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}\int^{x}_{0}{\rm d}t\int_{0}^{t}\left(z^{d-3}(d-z)(t-z)^{n-d}\right)\,{\rm d}z$
$=$ $\ell^{-(n-d+1)}\frac{n!}{(d-3)!(n-d+1)!}\int_{0}^{x}z^{d-3}(d-z)(x-z)^{n-d+1}\,{\rm d}z$
$=$ $\ell^{-(n-d+1)}x^{n-1}\left(dn-(d-2)x\right)\eqqcolon\phi(x).$
It is easy to see that $\phi(x)$ is an increasing function when $x\in[0,\frac{(n-1)d}{(d-2)}]$ with $\phi(d)=\ell^{-(n-d+1)}d^{n}(n-d+2)$. In particular, $\phi(x)$ is strictly increasing when $x\in[0,d]$ since $d\leq n+1$.
Next, consider the case when $x\geqslant d$. Then we need to split the sum in (18) into two parts:
$\sum_{y=0}^{dk}\sum_{i=0}^{\lfloor y/\ell\rfloor}a_{y,i}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd-\ell i))+\sum_{y=dk+1}^{xk-1}\sum_{i=0}^{\lfloor dk/\ell\rfloor}a_{y,i}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd-i\ell)).$
Similar calculation as above shows that this sum equals $b_{2}k^{n}/n!+O(k^{n-1})$ with $b_{2}$ equal to $C\cdot\mathbf{I}$ where $C=\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}$ and $\mathbf{I}$ is equal to:
$\int_{0}^{d}{\rm d}t\int_{0}^{t}z^{d-3}(t-z)^{n-d}(d-z){\rm d}z+\int_{d}^{x}{\rm d}t\int_{0}^{d}z^{d-3}(t-z)^{n-d}(d-z){\rm d}z$
$=$ $C^{-1}\phi(d)+\int_{0}^{d}z^{d-3}\frac{1}{n-d+1}\left((x-z)^{n-d+1}-(d-z)^{n-d+1}\right)(d-z){\rm d}z$
$=$ $\frac{1}{n-d+1}\int_{0}^{d}z^{d-3}(d-z)(x-z)^{n-d+1}{\rm d}z.$
We denote $V=L^{n}=(-K_{X})^{n}$. So we get the estimate ${\rm vol}(L-xE)\geq V-\phi(x)$ for any $x\geq 0$ where
(20) $\phi(x)=\left\{\begin{array}[]{ll}\ell^{-(n-d+1)}x^{n-1}(dn-(d-2)x)&\text{if }\;0\leq x\leq d\\ \frac{\ell^{-(n-d+1)}n!}{(d-3)!(n-d+1)!}\int_{0}^{d}z^{d-3}(d-z)(x-z)^{n-d+1}\,{\rm d}z&\text{if }\;x\geq d.\end{array}\right.$
Note that the function $\phi=\phi(x)$ belongs to the class $\mathscr{F}$ considered in Lemma 2.4. In other words, it is a continuous, piecewise smooth, strictly increasing function of $x\in[0,\infty)$ satisfying ${\rm lim}_{x\rightarrow 0}\phi(x)=0$ and ${\rm lim}_{x\rightarrow+\infty}\phi(x)=+\infty$. Set
[…]
Set $T={\rm sup}\{x:V-\phi(x)=0\}$. If $T\leq d$, then we get $V=\phi(T)\leq\phi(d)=\ell^{-(n-d+1)}d^{n}(n-d+2)\leq d^{n}(n-d+2)$. Note that the number $d^{n}(n-d+2)$ is nothing but the volume of the Fano hypersurface in ${\mathbb{P}}^{n+1}$ of degree $n-d+2$ which is always less than $2n^{n}$ if $22<d<n+$ 2.
So we can assume $T\geq d$. The equation $\Phi(T)=(T-A)\phi(T)$ from (4) with $A=A_{X}(E)=(d-2)+(n-d+1)\ell$ becomes:
$\frac{1}{n-d+2}\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+2}\,{\rm d}z=(T-(d-2)-(n-d+1)\ell)\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+1}\,{\rm d}z.$
Even though the explicit solution of $T$ is not available, as $\ell\rightarrow+\infty$, the above equation can be written as:
$\frac{1}{n-d+2}T^{n-d+2}(1+O(T^{-1}))=(T-(n-d+1)\ell+O(1))T^{n-d+1}(1+O(T^{-1})).$
So we get the asymptotic expression for its solution: $T=T(\ell)=(n-d+2)\ell+O(1)$. So we can calculate that, up to $O(\ell^{-1})$, $\phi(T)$ is given by:
[…]
$=$ $\ell^{-(n-d+1)}\frac{n!}{(d-3)!(n-d+1)!}\left(\int_{0}^{1}t^{d-3}(1-t){\rm d}t\right)\cdot d^{d-1}\ell^{n-d+1}(n-d+2)^{n-d+1}$
$=$ $(n-d+2)^{n-d+1}\frac{n!}{(d-1)!(n-d+1)!}d^{d-1}={\rm vol}({\mathbb{P}}^{n-d+1}\times{\mathbb{P}}^{d-1}).$
Letting $\ell\rightarrow+\infty$, by Corollary 2.5, we have thus proved ${\rm vol}(X)=(-K_{X})^{n}\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ and complete the proof of Theorem 1.4. ∎
Example 4.1.
By using binomial expansion, one can get the following combinatorial expressions for $x\geq d$:
$\phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=d-1}^{n}\binom{n}{j}\cdot(j-d+2)\cdot d^{j}\cdot(x-d)^{n-j},$
$\Phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=d-1}^{n+1}\binom{n+1}{j}\frac{j-d+2}{n+1}d^{j}\cdot(x-d)^{n+1-j}.$
When $d=n+1$, the above expression becomes: $\phi(x)=(n+1)^{n}$ (for $x\geq d$) which implies the volume estimate ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$.
When $d=n$, the above expressions become:
$\phi(x)$ $=$ $\ell^{-1}\left(n\cdot n^{n-1}\cdot(x-n)+2n^{n}\right)=\ell^{-1}n^{n}(x+2-n),$
$\Phi(x)$ $=$ $\frac{\ell^{-1}}{n+1}\left(\binom{n+1}{2}n^{n-1}(x-n)^{2}+(n+1)2n^{n}(x-n)+3n^{n+1}\right)$
$=$ $\frac{\ell^{-1}n^{n}}{2}\left(x^{2}-2(n-2)x+\frac{n}{(n+1)}(n-1)(n-2)\right).$
The equation $(T-(n-2)-\ell)\phi(T)=\Phi(T)$ is asymptotically given by:
$(T-\ell+O(1))T(1+O(T^{-1}))=\frac{1}{2}T^{2}(1+O(T^{-1}))$
which implies $T=2\ell+O(1)$. So $\phi(T)=\ell^{-1}n^{n}(2\ell+O(1)+2-n)\longrightarrow 2n^{n}$ as $\ell\rightarrow+\infty$ and we get an estimate: ${\rm vol}(X)\leq 2n^{n}={\rm vol}({\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1})={\rm vol}(Q^{n})$.
Proof of Theorem 1.1.
We use the notation $l_{X}$ from (6)-(7). If $l_{X}=n+1$, then as mentioned in Example 2.9 $X$ must be ${\mathbb{P}}^{n}$, and if $l_{X}=n$ then $X$ is either $Q^{n}$ or the blowup of ${\mathbb{P}}^{n}$ along a smooth subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ in a hyperplane. When $d_{Y}\geq 2$, by Lemma 4.2 below we know $(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}<2n^{n}$. When $d_{Y}=1$, by Lemma 4.3 below we know that ${\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is not a K-semistable Fano manifold, so it is excluded. Thus we verified the Theorem in the case $l_{X}\geq n$. When $l_{X}<n$, there are minimal rational curves of degree $d\in\{2,\dots,n-1\}$. For $12<d<n-$ 1, by Theorem 1.4 and Lemma 3.5, ${\rm vol}(X)\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})<2n^{n}$. When $d=2$, by Theorem 3.1 we get ${\rm vol}(X)\leq 2n^{n}$ with equality if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
Lemma 4.2.
Suppose $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\cdots,n\}$ that is contained in a hyperplane (See Example 2.9). Then
$(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}=\left\{\begin{array}[]{ll}2n^{n}&\text{if }d_{Y}=1\\ \frac{d_{Y}n^{n}-(n+1-d_{Y})^{n}}{d_{Y}-1}=:V_{d_{Y}}&\text{if }2\leq d_{Y}\leq n.\end{array}\right.$
Proof.
Denote $d=d_{Y}$. The normal bundle of $Y$ is given by $N_{Y/{\mathbb{P}}^{n}}={\mathcal{O}}_{Y}(1)\oplus{\mathcal{O}}_{Y}(d)$. Let $H=\pi^{*}{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ be the pullback of a hyperplane section of ${\mathbb{P}}^{n}$. Then,
$(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}$ $=$ $((n+1)H-E)^{n}=(n+1)^{n}+\sum_{k=2}^{n}\binom{n}{k}(n+1)^{n-k}\cdot(-1)^{k}(H^{n-k}\cdot E^{k})$
$=$ $(n+1)^{n}-\sum_{k=2}^{n}\binom{n}{k}(n+1)^{n-k}\cdot(s_{k-2}(N_{Y/{\mathbb{P}}^{n}})\cdot H^{n-k}|_{A})$
$=$ $(n+1)^{n}-\sum_{k=2}^{n}\binom{n}{k}(n+1)^{n-k}\cdot(-1)^{k-2}\left(\sum_{i=0}^{k-2}d^{i}\right)\cdot d,$
where $s_{k-2}(N_{Y/{\mathbb{P}}^{n}})$ denote the $(k-2)$-th Segre class of normal bundle $N_{Y/{\mathbb{P}}^{n}}$. If $d=1$, then $\sum_{i=0}^{k-2}d^{i}=k-1$. By the combinatorial identity $\binom{n}{k}(k-1)=n\binom{n-1}{k-1}-\binom{n}{k}$,
$(-K_{{\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}})^{n}$ $=$ $(n+1)^{n}-\sum_{k=2}^{n}(-1)^{k}\left[n\binom{n-1}{k-1}-\binom{n}{k}\right](n+1)^{n-k}=2n^{n}.$
If $d\geqslant 2$, then $\sum_{i=0}^{k-2}d^{i}=(d^{k-1}-1)/(d-1)$. Then,
$(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}$ $=$ $(n+1)^{n}-\sum_{k=2}^{n}(-1)^{k}\binom{n}{k}\cdot\left(\frac{d^{k}-d}{d-1}\right)(n+1)^{n-k}$
$=$ $\frac{1}{d-1}\left(dn^{n}-(n+1-d)^{n}\right).$
Note that by L’Hospital’s rule ${\rm lim}_{d_{Y}\rightarrow 1}V_{d_{Y}}=2n^{n}$ and it is easy to check that $V_{d_{Y}}<2n^{n}$ for $d_{Y}\geq 2$. ∎
Lemma 4.3.
Suppose $\pi\colon X\rightarrow{\mathbb{P}}^{n}$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y\cong{\mathbb{P}}^{n-2}$ of degree $d_{Y}=1$ that is contained in a hyperplane, then $X$ is K-unstable Fano manifold.
Proof.
Let $E=\pi^{-1}({\mathbb{P}}^{n-2})$ be the exceptional divisor. And let $H=\pi^{*}{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ be the pullback of a hyperplane section of ${\mathbb{P}}^{n}$. Note that the nef cone is $\text{Nef}(X)={\mathbb{R}}_{\geqslant 0}[H]+{\mathbb{R}}_{\geqslant 0}[H-E]$. Then the divisor $-K_{X}-tE=(n+1)H-(1+t)E$ is nef if and only if $0\leqslant t\leqslant n$. When $0\leqslant t\leqslant n$, by the similar computation as in Lemma 4.2,
${\rm vol}_{X}(-K_{X}-tE)$ $=$ $((n+1)H-(1+t)E)^{n}=(n+1)^{n}-\sum_{k=2}^{n}\binom{n}{k}(n+1)^{n-k}(-1-t)^{k}(k-1)$
$=$ $(n-t)^{n-1}(2n+(n-1)t).$
We see the pseudo-effective threshold $T_{X}(E)$ equals to the nef threshold $n$. Then,
$S(-K_{X};E)=\frac{1}{(-K_{X})^{n}}\int_{0}^{n}{\rm vol}_{X}(-K_{X}-tE){\rm d}t=1+\frac{n-1}{2(n+1)}.$
So $A_{X}(E)-S(-K_{X};E)=1-(1+\frac{n-1}{2(n+1)})=-\frac{n-1}{2(n+1)}<0$. By the valuative criterion (Theorem 2.1), we conclude that $X$ is K-unstable and destabilized by the exceptional divisor $E$. ∎
Remark 4.4.
Note that $X={\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is a toric Fano variety, so one can also prove lemma 4.3 via toric geometry. The minimal generators of the fan $\Sigma\subseteq{\mathbb{R}}^{n}$ is given by $e_{1}=(1,0\cdots,0),\cdots,e_{n}=(0,\cdots,0,1),e_{0}=(-1,-1,\cdots,-1),f=(1,1,0,\cdots,0)$. Here the ray generated by $f$ corresponds to the exceptional divisor $E=\pi^{-1}({\mathbb{P}}^{n-2})$ by the cone-orbit correspondence. Then the moment polytope with respect to $-K_{X}$ is given by
$P_{X}=\{(x_{1},\cdots,x_{n})\in{\mathbb{R}}^{n}\mid x_{1}+x_{2}\geq-1,\,\sum_{i=1}^{n}x_{i}\leq 1,\,x_{j}\geq-1\,\,\text{for}\,\,j=1,\cdots,n\}.$
By calculus, one can compute the barycenter ${\rm Bc}(P)=(\frac{n-1}{4(n+1)},\frac{n-1}{4(n+1)},-\frac{1}{2(n+1)},\cdots,-\frac{1}{2(n+1)})\in{\mathbb{R}}^{n}$, which does not coincide with the origin. Thus, $X$ is K-unstable, so this gives another proof of Lemma 4.3. In fact, by the formula $\delta(X,-K_{X})={\rm min}_{\rho\subset\Sigma(1)}\frac{1}{\langle{\rm Bc}(P),u_{\rho}\rangle+1}$ in [BJ20, Section 7], we know $\delta(X)=\frac{2n+2}{3n+1}<1$ and it is minimized at $\rho=f=(1,1,0\cdots,0)$. So the exceptional divisor $E$ in Lemma 4.3 is indeed the $\delta$-minimizer of $X$.
As an immediate application of Theorem 1.1, we prove the ODP conjecture (Conjecture 2.3) for Fano cones over K-semistable Fano manifolds.
Theorem 4.5.
Suppose $L=r^{-1}(-K_{X})$ is an ample line bundle over a $n$-dimensional smooth K-semistable Fano manifold $X$ for some $r\in{\mathbb{N}}^{*}$, then ODP conjecture (Conjecture 2.3) holds for the affine cone ${\mathcal{C}}(X,L)\coloneqq{\rm Spec}(\oplus_{k=0}^{+\infty}H^{0}(X,kL))$ with the cone vertex $o$.
Proof.
We denote $\pi\colon{\rm Bl}_{o}{\mathcal{C}}\rightarrow{\mathcal{C}}$ be the blowup of ${\mathcal{C}}$ along the cone vertex $o$. By the result of [Li17, LL19], we know $X$ is K-semistable if and only if the minimizer of normalized volume of $({\mathcal{C}},o)$ is attained by the canonical valuation $v={\rm ord}_{E}$, which is the divisorial valuation associated to the exceptional divisor $E=\pi^{-1}(o)\cong X$. From [Kol13, Section 3.1], we know
$K_{{\rm Bl}_{o}{\mathcal{C}}}+(1-r)E\sim_{{\mathbb{Q}}}0\sim_{{\mathbb{Q}}}\pi^{*}(K_{{\mathcal{C}}}).$
Then $A_{{\mathcal{C}}}({\rm ord}_{E})=1+{\rm ord}_{E}(K_{{\rm Bl}_{o}{\mathcal{C}}}-\pi^{*}K_{{\mathcal{C}}})=1+(r-1)=r$ and ${\rm vol}({\rm ord}_{E})=L^{n}$. Then
${\widehat{\rm vol}}({\mathcal{C}},o)=A_{{\mathcal{C}}}({\rm ord}_{E})^{n+1}\cdot{\rm vol}(v)=r^{n+1}\cdot L^{n}=r(-K_{X})^{n}.$
Let $i(X)\coloneqq{\rm max}\{r\in{\mathbb{N}}^{*}\mid-K_{X}\sim_{{\mathbb{Q}}}rL\,\text{for some ample line bundle }\,L\}$ be the Fano index of $X$. Then the above argument shows that ${\widehat{\rm vol}}({\mathcal{C}},o)\leqslant i(X)\cdot(-K_{X})^{n}$. We know that for smooth Fano manifold, the Fano index always satisfies $i(X)\leqslant n+1$ and $i(X)=n+1$ if and only if $X\cong{\mathbb{P}}^{n}$ and $i(X)=n$ if and only if $X\cong Q\subset{\mathbb{P}}^{n+1}$ is a smooth quadric hypersurface ([KO73]). If $X\cong{\mathbb{P}}^{n}$ and $r=i({\mathbb{P}}^{n})=n+1$, then $({\mathcal{C}}({\mathbb{P}}^{n},{\mathcal{O}}(1)),o)\cong({\mathbb{C}}^{n},0)$ is a smooth point, which is excluded in the ODP conjecture. If $r<i({\mathbb{P}}^{n})=n+1$, then there exists some $p\in{\mathbb{Z}}_{>1}$ such that $rp=i({\mathbb{P}}^{n})=n+1$. Then ${\widehat{\rm vol}}({\mathcal{C}}({\mathbb{P}}^{n},L),o)=r(-K_{{\mathbb{P}}^{n}})^{n}=(n+1)^{n+1}/p<2n^{n+1}$. So we can assume $X\ncong{\mathbb{P}}^{n}$. Then by Theorem 1.1, we have $(-K_{X})^{n}\leqslant 2n^{n}$ and the equality holds if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q$. Combined with Kobayashi-Ochiai’s result $i(X)\leqslant n$, we get
${\widehat{\rm vol}}({\mathcal{C}},o)\leqslant i(X)\cdot(-K_{X})^{n}\leqslant 2n^{n+1}.$
The equality holds if and only if $i(X)=n$ and $(-K_{X})^{n}=2n^{n}$, which implies $({\mathcal{C}},o)\cong({\mathcal{C}}(Q,{\mathcal{O}}_{Q}(1)),o)$ is the ODP singularity (the case when $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ is excluded since the Fano index $i({\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1})=2$). ∎
4.2. Examples and questions
Example 4.6.
Our consideration of weighted blowup is motivated by a special example of a weighted blowup over the quadric hypersurface: $X=Q^{n}=\{Z_{0}Z_{1}+Z_{2}Z_{3}+Z_{4}^{2}+\cdots+Z_{n+1}^{2}=0\}\subset{\mathbb{P}}^{n+1}$ for $n\geq 3$. Fix a minimal rational curve which is the line
$C\cong{\mathbb{P}}^{1}=\{Z_{1}=Z_{3}=Z_{4}=\cdots=0\}=\{[t:0:s:0:0:\cdots:0]\colon[t,s]\in{\mathbb{P}}^{1}\}\subset{\mathbb{P}}^{n+1}.$
On the affine chart $U_{0}=\{Z_{0}\neq 0\}$, set $u_{i}=z_{i}/z_{0}$ with $i\neq 0$. Then the equations of $X$ and $C$ become:
$X:$ $u_{1}+u_{2}u_{3}+u_{4}^{2}+\cdots+u_{n+1}^{2}=0\Longrightarrow u_{1}=-u_{2}u_{3}-u_{4}^{2}-\cdots-u_{n+2}^{2},$
$C:$ $u_{3}=u_{4}=\cdots=u_{n+1}=0.$
In particular, we choose coordinates $\mathbf{u}=\{u_{2},u_{3},u_{4},\cdots,u_{n+1}\}$ as local coordinates on $X$. Similarly, on the affine chart $U_{1}=\{Z_{2}\neq 0\}$, if we set $v_{j}=z_{j}/z_{2}$ with $j\neq 2$. The local coordinate is given by $\mathbf{v}=(v_{0},v_{1},v_{4}\cdots,v_{n+1})$. The equations of $X$, $C$ are given by:
$X:$ $v_{0}v_{1}+v_{3}+v_{4}^{2}+\cdots+v_{n+1}^{2}=0\Longrightarrow v_{3}=-v_{0}v_{1}-v_{4}^{2}-\cdots-v_{n+1}^{2},$
$C:$ $v_{1}=v_{4}=\cdots=v_{n+1}=0.$
The conormal bundle of $C\setminus\{x\}$ is generated by
$\{dv_{1}=-du_{3},dv_{4}=u_{2}^{-1}du_{4},\cdots,dv_{n+1}=u_{2}^{-1}du_{n+1}\}$
which shows in particular that $N^{\vee}_{C/X}={\mathcal{O}}\oplus{\mathcal{O}}(-1)^{\oplus n-2}$.
Consider the ${\mathbb{C}}^{*}$ action on ${\mathbb{P}}^{n+1}$ with weights $(0,2,0,2,\underbrace{1,\dots,1}_{n-3})$ on the homogeneous coordinates. The ${\mathbb{C}}^{*}$ action preserves the quadric hypersurface and gives rise to a divisorial valuation whose center on $X$ is the line $C$. This divisorial valuation can also be obtained via a weighted blowup along the line $C$. Precisely, in the $\mathbf{u}$-coordinates, we perform the weighted blow along $C\cap U_{0}=\{u_{3}=u_{4}=\cdots=u_{n+1}=0\}$ with weights $(0,2,1,\dots,1)$. In the $\mathbf{v}$ coordinates, we also have the weights $(0,2,1,\dots,1)$. The weighted blowup $\mu:\hat{X}\rightarrow X$ in this example is globally defined and the exceptional divisor $E$ is isomorphis to a weighted projective ${\mathbb{P}}(2,\underbrace{1,\dots,1}_{n-2})$-bundle over ${\mathbb{P}}^{1}$. We can calculate the volume function ${\rm vol}(-K_{X}-xE)=n^{n}{\rm vol}(H-\xi E)$ where $\xi=\frac{x}{n}$. Interestingly the situation turns out to be different for $n=3$ and $n\geq 4$. When $n=3$, the nef threshold (Seshadri constant) of $H$ with respect to $E$ coincides with the pseudo-effective threshold and is equal to 2. We have the expression:
${\rm vol}(-K_{X}-xE)=27\cdot{\rm vol}(H-\xi E)=27\cdot\left(2-\frac{3}{2}\xi+\frac{1}{2}\xi^{2}\right),\quad 0\leq\xi\leq 2.$
When $n\geq 4$, the nef threshold of $H$ with respect to $E$ is 1 while the pseudoeffective threshold of $H$ with respect to $E$ is 2. The volume function consists of two smooth pieces:
(22) ${\rm vol}(-K_{X}-xE)$ $=$ $n^{n}{\rm vol}(H-\xi E)$
(24) $=$ $n^{n}\cdot\left\{\begin{array}[]{ll}2-\frac{n}{2}\xi^{n-1}+\frac{n-2}{2}\xi^{n},&\xi=\frac{x}{n}\in[0,1]\\ \frac{n}{2}(2-\xi)^{n-1}-\frac{n-2}{2}(2-\xi)^{n},&\xi=\frac{x}{n}\in[1,2].\end{array}\right.$
In fact, this expression holds for any $n\geq 3$, since when $n=3$ the second piece and the first piece connect smoothly and become one whole smooth piece. To get the above formula, we can first calculate the volume function before the nef threshold by calculating the intersection:
$(\mu^{*}H-\xi E)^{n}$ $=$ $\sum_{k=0}^{n}\binom{n}{k}\mu^{*}H^{n-k}\cdot\xi^{k}(-E)^{k}$
$=$ $H^{n}+n\xi^{n-1}\pi^{*}H\cdot\xi^{n-1}(-E)^{n-1}+\xi^{n}(-E)^{n}$
$=$ $2-\frac{n}{2}\xi^{n-1}+\xi^{n}\frac{n-2}{2}$
where we used a weighted analogue of the usual intersection formula via Segre classes for projective bundles. It is not surprising that the expression of this piece coincides with the piece in the expression of (20) when $x=n\xi\leq n$. The second piece is more difficult to get. One way to get it is to use the fact that the measure $-\frac{1}{n!}d{\rm vol}(H-tE)$ is the Duistermaat-Heckman measure of the previously mentioned Hamiltonian $S^{1}$ action (see [BHJ17]) and then use the symmetry of its density function around $\xi=1$ in our case. We can also directly calculate the Duistermaat-Heckman measure of $S^{1}$-action by using a well-established localization/Fourier formula ([BGV92, 7.4]) and for the interest of the reader we will provide such a derivation in appendix.
Example 4.7.
Let $X\subset{\mathbb{P}}^{n+1}$ be a Fano hypersurface of degree $1\leq b\leq n$. Then the minimal rational curve is of splitting type ${\mathcal{O}}(1)^{n-b}\oplus{\mathcal{O}}^{b-1}$ so that $d=n-b+2$ (see [Kol96, Exercise V.4.4]). The volume of $X$ is given by $V_{d}=d^{n}(n-d+2)=(n+2-b)^{n}\cdot b$. One can easily verify that $V_{d}$ is strictly smaller than ${\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ except when $b=1$ or $b=2$.
Question 4.8.
Is there a K-semistable Fano manifold $X$ with ${\rm vol}(X)={\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ that contains a minimal rational curve of degree $13\leq d\leq n-$ 1 but is not isomorphic to ${\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1}$?
Our results imply that when $d=n$, there is just one such $X$, i.e. the quadric hypersurface $Q^{n}$, while when $d=2$, there is no such $X$. Moreover, such an $X$ must satisfy $l_{X}=d$ (see (7)).
We end this paper by finding the minimal possible anti-canonical volume of $n$-dimensional (K-semistable) Fano manifolds. It is not a well-posed problem for general ${\mathbb{Q}}$-Fano varieties. Even assuming K-semistablity, there exists a sequence of K-semistable ${\mathbb{Q}}$-Fano varieties with volume $(-K_{X})^{n}$ tends to zero (See [Jia20, Example 1.4(2)]), so it is in general not possible to determine an optimal positive lower bound. When we restrict to $n$-dimensional smooth Fano manifold $X$, the volume $(-K_{X})^{n}$ is always a positive integer, so a priori $(-K_{X})^{n}\geqslant 1$.
Example 4.9.
Suppose $n\geqslant 2$ is an integer. Assume $a_{0}\leqslant a_{1}\leqslant\cdots\leqslant a_{n+1}$ are integers and we consider a general degree $d$ hypersurface $X_{d}$ of weighted projective space ${\mathbb{P}}(a_{0},a_{1}\cdots,a_{n+1})$. Note that $X_{d}$ is a Fano variety if $d<\sum_{i=0}^{n+1}a_{i}$. By the adjunction formula
$(-K_{X})^{n}=\frac{d\cdot(\sum_{i=0}^{n+1}a_{i}-d)^{n}}{\prod_{i=0}^{n+1}a_{i}}.$
If $a_{0}=\cdots=a_{n-1}=1,\,a_{n}=2,\,a_{n+1}=n+1$ and $d=2n+2$. Namely, $X_{d}$ is a general degree $22n+$ 2 hypersurface in ${\mathbb{P}}(1^{n},2,n+1)$. Then $(-K_{X_{2n+2}})^{n}=1$. By [IF00, Thm 8.1], we know that $X_{2n+2}$ is smooth when $n$ is even. By [ST24, Thm 1.3], we know $\delta(X_{2n+2})\geqslant(n+1)/2>1$, so $X_{2n+2}$ is a smooth K-stable Fano manifold with minimal anti-canonical volume $(-K_{X_{2n+2}})^{n}=1$ for all even $n$. When $n$ is odd, the general hypersurface $X_{2n+2}\subset{\mathbb{P}}(1^{n},2,n+1)$ is singular. (After communicating with K. Fujita, we find that this type of example already appears in [KSC04, Theorem 5.22].)
Example 4.10.
Let $Y$ be the double cover of ${\mathbb{P}}^{n}$ ramified along a degree $2n$ smooth hypersurface $D$. Then by Hurwitz’s formula, $K_{Y}=\phi^{*}(K_{{\mathbb{P}}^{n}}+\frac{1}{2}D)$. So $(-K_{Y})^{n}=\deg(\phi)\cdot(-K_{{\mathbb{P}}^{n}}-\frac{1}{2}D)^{n}=2$. By the result of [LZ22, Zhu21], we know K-stability of $Y$ is equivalent to the K-stability of the log Fano pair $({\mathbb{P}}^{n},\frac{1}{2}D)$. Since we assume $D$ is a smooth hypersurface, then the log Fano pair $({\mathbb{P}}^{n},(\frac{n+1}{2n}-\varepsilon)D)$ is klt, so it is uniformly K-stable by [ADL23, Theorem 2.10] for $0<\varepsilon\ll 1$. Since $D\sim_{{\mathbb{Q}}}-K_{{\mathbb{P}}^{n}}$, by the interpolation of K-stability ([ADL24, Prop 2.13]), we conclude that $({\mathbb{P}}^{n},\frac{1}{2}D)$ is K-stable. Thus, $Y$ is a smooth K-stable Fano manifold with volume $(-K_{Y})^{n}=2$ for all dimensions $n\geqslant 2$. Note that $Y$ contains a degree 2 free-immersed rational curve with $n-1$ nodes (see [Kol96, IV.2.12.3]).
We indeed have the following:
Proposition 4.11.
The minimum of anticanonical volumes of $n$-dimensional Fano manifolds is equal to 1 if $n$ is even and is equal to 2 if $n$ is odd. The minimum volume is obtained by K-stable Fano manifolds.
This follows from Example 4.9, Example 4.10 and following interesting fact pointed out to us by K. Fujita:
Fact 1.
([KSC04, Solution to Exercise 5.23, Page 217]) Let $X$ be a smooth projective variety of odd dimension $n$. Then the self-intersection number $(-K_{X})^{n}$ is even.
Question 4.12.
Classify all Fano manifolds that obtain the minimal volume.
Appendix: A proof of the formula (22)
We calculate the Duistermaat-Heckman measure via the localization method. Let $\omega=\omega_{\mathrm{FS}}$ denote the restriction of the Fubini-Study metric of ${\mathbb{P}}^{n+1}$ on $X=Q^{n}=\{Z_{0}Z_{1}+Z_{2}Z_{3}+Z_{4}^{2}+\cdots+Z_{n+1}^{2}=0\}\subset{\mathbb{P}}^{n+1}$ for $n\geq 3$. The ${\mathbb{C}}^{*}$ action on $X$ is given by: for any $\lambda\in{\mathbb{C}}^{*}$,
$\lambda\circ(Z_{0},Z_{1},Z_{2},Z_{3},Z_{4},\dots,Z_{n+1})=(Z_{0},\lambda^{2}Z_{1},Z_{2},\lambda^{2}Z_{3},\lambda Z_{4},\dots,\lambda Z_{n+1}).$
The corresponding $S^{1}$ action is a Hamiltonian with moment map given by:
$\mu([Z])=\frac{2|Z_{1}|^{2}+2|Z_{3}|^{2}+|Z_{4}|^{2}+\cdots+|Z_{n+1}|^{2}}{|Z_{0}|^{2}+|Z_{1}|^{2}+\cdots+|Z_{n+1}|^{2}}.$
We introduce the equivariant volume functional:
$V(\theta)=\int_{X}e^{i\mu(z)\xi}\frac{\omega^{n}}{n!}=\int_{\mathbb{R}}e^{i\theta\xi}\,\mathrm{DH}(\xi)$
where $\mathrm{DH}(\xi)=\mu_{*}\frac{\omega^{n}}{n!}$ is the Duistermaat-Heckman measure. As a consequence, the density function of $\mathrm{DH}(\xi)$ is the Fourier transform of $V(\theta)$.
The localization formula calculates the integral $V(\theta)$ via data on the fixed points sets (including the values of moment map and equivariant Euler curvature forms of normal bundles) which consists of three components in this case:
$F_{0}=\mu^{-1}(0)\cong{\mathbb{P}}^{1}:[*,0,*,0,\dots,0],$
$F_{1}=\mu^{-1}(2)\cong{\mathbb{P}}^{1}:[0,*,0,*,0,\dots,0],$
$F_{2}=\mu^{-1}(1)\cong Q^{n-4}:[0,0,0,0,Z_{4},\dots,Z_{n+1}]\text{ with }Z_{4}^{2}+\cdots Z_{n+1}^{2}=0.$
The contribution from $F_{0}$ is given by:
$V_{0}(\theta):=\int_{{\mathbb{P}}^{1}}\frac{e^{\omega}}{((\omega-i\theta))^{n-2}(-i2\theta)}=\frac{(-1)^{n+1}2\pi}{2}((i\theta)^{1-n}+(n-2)(i\theta)^{-n}).$
We have the following useful formula: for any $a\in{\mathbb{R}}$,
$\frac{1}{2\pi}\int_{\mathbb{R}}\frac{e^{ia\theta\xi}}{(i\theta)^{k}}=\frac{(a-\xi)^{k-1}}{2(k-1)!}\mathrm{Sign}(a-\xi).$
So we can calculate the Fourier transform of $V_{0}(\theta)$:
$\frac{1}{2\pi}\int_{{\mathbb{R}}}V_{0}(\theta)e^{-i\theta\xi}{\rm d}\theta$ $=$ $\frac{1}{4(n-1)!}((n-2)\xi^{n-1}-(n-1)\xi^{n-2})\mathrm{Sign}(-\xi)$
=:absent:=:= : $\frac{f(\xi)}{2}\mathrm{Sign}(-\xi).$
The contribution from $F_{2}$ is given by:
$V_{2}(\theta):=\int_{{\mathbb{P}}^{1}}\frac{e^{(\omega+2i\theta)}}{((\omega+i\theta))^{n-2}(i2\theta)}=\frac{2\pi}{2}e^{2i\theta}((i\theta)^{1-n}+(2-n)(i\theta)^{-n}).$
Similar to above, the Fourier transform of $V_{2}(\theta)$ is equal to:
$\frac{1}{2\pi}\int_{{\mathbb{R}}}V_{2}(\theta)e^{-i\theta\xi}{\rm d}\theta$ $=$ $\frac{1}{4(n-1)!}(-(n-2)(2-\xi)^{n-1}+(n-1)(2-\xi)^{n-2})\mathrm{Sign}(2-\xi)$
=:absent:=:= : $\frac{g(\xi)}{2}\mathrm{Sign}(2-\xi).$
The contribution from $F_{1}$ is given by:
$V_{1}(\theta)$ ::=:= $\int_{Q^{n-4}}\frac{e^{(\omega+i\theta)}}{(\omega+i\theta)^{2}(\omega-i\theta)^{2}}=\int_{Q^{n-4}}\frac{e^{\omega}e^{i\theta}}{(\omega^{2}+\theta^{2})^{2}}$
$=$ $\sum_{k=0}^{\lfloor\frac{n-4}{2}\rfloor}\frac{(k+1)}{(n-4-2k)!}\frac{e^{i\theta}}{(i\theta)^{2k+4}}.$
The Fourier transform of $V_{1}(\theta)$ is equal to:
$\frac{1}{2\pi}\int_{{\mathbb{R}}}V_{1}(\theta)e^{-i\theta\xi}{\rm d}\theta$ $=$ $\sum_{k=0}^{\lfloor\frac{n-4}{2}\rfloor}\frac{(k+1)}{2(n-4-2k)!(2k+3)!}(1-\xi)^{2k+3}\mathrm{Sign}(1-\xi).$
We claim that this is equal to $-\frac{f(\xi)+g(\xi)}{2}\mathrm{Sign}(1-\xi)$ so that the Fourier transform is $V(\theta)$ is equal to:
$\rho(\xi)$ $=$ $\frac{f(\xi)}{2}\mathrm{Sign}(-\xi)+\frac{g(\xi)}{2}\mathrm{Sign}(2-\xi)-\frac{f(\xi)+g(\xi)}{2}\mathrm{Sign}(1-\xi)$
$=$ $\left\{\begin{array}[]{ll}-f(\xi)&0\leq\xi\leq 1\\ g(\xi)&1\leq\xi\leq 2\\ 0&\mathrm{otherwise}.\end{array}\right.$
It is immediate to check that $\rho(\xi)=-\frac{1}{n!}\frac{d}{d\xi}{\rm vol}(H-\xi E)$ by using the expression from (22), which verifies the formula (22) after integration. To verify the claim, we calculate:
$\xi^{n-1}-(2-\xi)^{n-1}$ $=$ $(1-(1-\xi))^{n-1}-(1+(1-\xi))^{n-1}=\sum_{k=0}^{n-1}\binom{n-1}{k}((-1)^{k}-1)(1-\xi)^{k}$
$=$ $-\sum_{j=0}^{\lfloor\frac{n-2}{2}\rfloor}\binom{n-1}{2j+1}(1-\xi)^{2j+1},$
and finally verify the claimed identity:
$-2(n-1)!(f(\xi)+g(\xi))$ $=$ $(n-2)(\xi^{n-1}-(2-\xi)^{n-1})-(n-1)(\xi^{n-2}-(2-\xi)^{n-2}),$
$=$ $\sum_{j=0}^{\lfloor\frac{n-2}{2}\rfloor}(n-2)\binom{n-1}{2j+1}(1-\xi)^{2j+1}-\sum_{j=0}^{\lfloor\frac{n-3}{2}\rfloor}(n-1)\binom{n-2}{2j+1}(1-\xi)^{2j+1}$
$=$ $\sum_{j=0}^{\lfloor\frac{n-2}{2}\rfloor}\frac{(n-1)!2j}{(2j+1)!(n-2-2j)!}(1-\xi)^{2j+1}$
$=$ $2(n-1)!\sum_{k=0}^{\lfloor\frac{n-4}{2}\rfloor}\frac{k+1}{(n-4-2k)!(2k+3)!}(1-\xi)^{2k+3}.$

### v2

@@
 Abstract.
-We prove that the volume of an $n$-dimensional K-semistable Fano manifold that is not ${\mathbb{P}}^{n}$ is at most $2n^{n}$.
-Moreover, the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$.
+We prove that the anti-canonical volume of an $n$-dimensional K-semistable Fano manifold that is not ${\mathbb{P}}^{n}$ is at most $2n^{n}$.
+Moreover, the volume is equal to $2n^{n}$ if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$.
 Our proof is based on a new connection between K-semistability and minimal rational curves.
-More generally, we show that the volume of a K-semistable Fano manifold with a minimal rational curve of degree $d$ is bounded above by the volume of ${\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1}$.
 1. Introduction
@@
 For a Fano variety $X$, its (anticanonical) volume ${\rm vol}(X)$ is defined to be the self-intersection number $(-K_{X})^{n}$.
-It is known that the volume of an $n-$dimensional Fano manifold can exceed the volume of ${\mathbb{P}}^{n}$, even among toric Fano manifolds (see [Deb01, Section 5.11]).
-In a major advance in the study of K-stability, Fujita [Fuj18] proved that the volume of an $n$-dimensional K-semistable Fano manifolds $X$ satisfies ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$ and the first equality holds if and only if $X\cong{\mathbb{P}}^{n}$.
+It is known that the volume of an $n$-dimensional Fano manifold can exceed the volume of the complex projective space ${\mathbb{P}}^{n}$, even among toric Fano manifolds (see [Deb01, Section 5.11]).
+In a major advance in the study of K-stability, Fujita [Fuj18] proved that the volume of an $n$-dimensional K-semistable Fano manifold $X$ satisfies ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$ and the first equality holds if and only if $X\cong{\mathbb{P}}^{n}$.
 It was later generalized to possibly singular $\mathbb{Q}$}-Fano varieties by Liu [Liu18].
@@
 Theorem 1.1.
-Any K-semistable Fano manifold $X$ that is not ${\mathbb{P}}^{n}$ satisfies ${\rm vol}(X)=(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$.
-In particular, this holds for any Kähler-Einstein metric manifold with positive Ricci curvature.
-Like the result of Fujita, the last statement could be seen as a result in Kähler differential geometry.
+Any K-semistable Fano manifold $X$ that is not ${\mathbb{P}}^{n}$ satisfies $(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$.
+In particular, this holds for any Kähler manifold admitting Kähler-Einstein metric with positive Ricci curvature.
+Like the result of Fujita ([Fuj18]), the last statement could be seen as a result in Kähler geometry.
 However, our proof uses purely algebraic geometry.
-The above result seems to be first conjectured in [AIM20, Problem 2.6], and is stated as a folklore conjecture in [Zhu25, Conjecture 6.8] which highlights the interesting (but also mysterious) feature that there are two Fano manifolds with second largest volume.
+The above result seems to be first conjectured in [AIM20, Problem 2.6], and is stated as a folklore conjecture in [Zhu25, Conjecture 6.8] which highlights the interesting (but also mysterious) feature that there are two Fano manifolds with the second largest volume.
 Our proof will indeed give a satisfactory explanation of this feature by connecting it to the theory of minimal rational curves on Fano manifolds.
 There is a local analogue of this conjecture sometimes called the ODP conjecture (see Conjecture 2.3) for the volume of klt singularity (see [SS17, LX19]).
-The ODP conjecture would imply that the volume of singular K-semistable Fano varieties is strictly less than $2n^{n}$ (by Liu’s local-to-global volume comparison, see Remark 3.9).
+The ODP conjecture would imply that the volume of singular K-semistable Fano varieties is strictly less than $2n^{n}$ (by Fujita-Liu’s local-to-global volume comparison, see Remark 3.7).
 Conversely, using Theorem 1.1, we immediately verify the ODP conjecture for the Fano cone over a K-semistable smooth Fano manifold (see Theorem 4.15).
 The toric version of Theorem 1.1 has its own interest and is called the “gap hypothesis” in [AB24, Conjecture 3.10].
-It implies the sharp bound of the canonical height of canonical model of toric Fano varieties over ${\rm Spec}\,{\mathbb{Z}}$ by [AB24, Lemma 3.8].
+It implies the sharp bound of a canonical height of canonical model of toric Fano varieties over $\rm Spec$$\mathbb{Z}$} by [AB24, Lemma 3.8].
 Moreover, in the toric case, the ODP conjecture is solved by Moraga-Süß’s result ([MS24]) using convex geometric methods.
 So combined with Theorem 1.1, we indeed have the sharp upper bound for all possibly singular K-semistable toric Fano varieties.
-Theorem 1.2.
-The second largest volume of $n$-dimensional K-semistable toric Fano varieties is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
+Theorem 1.2 (See Theorem 3.6).
+Let $X$ be an $n$-dimensional K-semistable toric Fano variety that is not ${\mathbb{P}}^{n}$.
+Then $(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
 We remark that Theorem 1.2 was proved in [AB24] for smooth toric Fano manifolds of dimension $n\leqslant 6$ by using the classification of [Obr07] and also for certain singular toric Fano varieties by convex geometric arguments (see [AB24, Lemma 3.9]).
 Furthermore, using the well-known one-to-one correspondence between Gorenstein toric Fano varieties and reflexive lattice polytopes in ${\mathbb{R}}^{n}$, Theorem 1.2 immediately implies a convex geometric statement.
-Corollary 1.3.
-Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with barycenter at $0\in{\mathbb{R}}^{n}$.
-Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $[0,2]\times(n\Delta_{n-1})$.
-We will prove Theorem 1.1 by exploring a new connection between K-stability and minimal rational curves.
-Roughly speaking, we are going to estimate the volume by using the valuative criterion to test K-semistability via valuations associated to weighted blowups along minimal rational curves.
-An interesting new phenomenon we find is that the volume estimates become sharp when we let certain weights go to infinity, and we indeed get the following general result, which to our knowledge is the first result connecting explicitly the theory of K-stability and the theory of minimal rational curves on Fano manifolds.
+Corollary 1.3 (See Corollary 3.9).
+Suppose $P\subseteq{\mathbb{R}}^{n}$ is an $n$-dimensional reflexive lattice polytope with its barycenter at $0\in{\mathbb{R}}^{n}$.
+Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$.
+Then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $0,2]\times(n\Delta_{n-1})$.
+We will prove Theorem 1.1 by exploring a new connection between K-stability and minimal rational curves. We use the following invariant
+$l_{X}={\rm min}\{(-K_{X}\cdot C)|\,C\subset X\,\text{is a minimal rational curve on}\,X\}\in\{2,3,\cdots,\dim X+1\}.$
+For the case of $l_{X}=\dim X+1$ and $l_{X}=\dim X$, we can use the classification result of [CMSB02, Miy04, CD15, DH17] to verify Theorem 1.1.
+At the other extreme, if $l_{X}=2$ and when the minimal rational curve is embedded, then the minimal rational curve has a trivial normal bundle.
+In this special case Theorem 1.1 is a consequence of the following generalization of a result of K.
+Fujita ([Fuj18]).
 Theorem 1.4.
-Assume that a K-semistable Fano manifold $X$ admits a minimal rational curve $f:{\mathbb{P}}^{1}\rightarrow X$ such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ for $12\leq d\leq n+$ 1.
-Then we have ${\rm vol}(X)\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$.
-The smooth toric case is special in the sense that there are always smoothly embedded projective spaces with trivial normal bundles (see [CFH14, Ara06]).
-So the toric Fano case can also be considered as an example of the following direct generalization of Fujita’s result.
-Theorem 1.5.
-Let $X$ be an $n$-dimensional K-semistable Fano manifold, and $Z\subset X$ a codimension $r$ smooth complex submanifold.
+Let $X$ be an $n$-dimensional K-semistable Fano manifold, and $Z\subset X$ be a codimension $r$ smooth complex submanifold.
 Assume that the normal bundle of $Z$ inside $X$ is trivial and set $d=(-K_{X})^{n-r}\cdot Z$.
@@
 (1) $(-K_{X})^{n}\leqslant(r+1)^{r}\cdot\binom{n}{r}\cdot d=(-K_{{\mathbb{P}}^{r}\times Z})^{n}.$
-Moreover, the equality holds if and only if $X$ is biholomorphic to ${\mathbb{P}}^{r}\times Z$. In particular $Z$ is also a K-semistable Fano manifold.
+Moreover, the equality holds if and only if $X$ is biholomorphic to ${\mathbb{P}}^{r}\times Z$. In particular in the equality case $Z$ is also a K-semistable Fano manifold.
+To estimate the volume in much more involved cases when $3\leq d=l_{X}\leq n-1$, we are going to use the valuative criterion to test K-semistability via special valuations centered at minimal rational curves.
+While the proof of Theorem 1.4 uses ordinary blowup of submanifolds, the proof of Theorem 1.1 needs weighted blowups (see Remark 4.11).
+In particular, we will define and analyze a particular weighted blowup of type $({1}^{\oplus(d-2)},2^{\oplus(n-d+1)})$ that is adapted to the splitting of normal bundle over ${\mathbb{P}}^{1}$ (see (12)).
+The analysis will also allow us to overcome the difficulty that the minimal rational curve may have singularities.
 We sketch the organization of this paper.
 In the next section, we recall basic knowledge about some key concepts used in later sections: K-stability, Seshadri constant, and minimal rational curves.
-In section 3, we prove Theorem 1.5, and Theorem 1.1 in the case the Fano manifold contains a minimal rational curve with a trivial normal bundle.
+In section 3, we prove Theorem 1.4 which implies Theorem 1.1 in the special case when the Fano manifold contains an embedded minimal rational curve with a trivial normal bundle.
 We also deduce Theorem 1.2 and Corollary 1.3.
-In section 4, we prove Theorem 1.4 to achieve the full proof of Theorem 1.1.
+In section 4, we give the full proof of Theorem 1.1.
 We end the paper by giving some related examples and finding (K-semistable) Fano manifolds with minimal volumes.
@@
 We thank Robert Berman, Kento Fujita, Yuchen Liu, Chenyang Xu and Ziquan Zhuang for their interest and helpful comments, and Kento Fujita for pointing to us the Fact 1.
+We are especially grateful to Ziquan Zhuang for kindly pointing out some error in our estimates that appeared in the first version of this preprint (see Remark 4.4) and very helpful suggestions.
 This work is carried out during M.
@@
 Unless otherwise specified, all varieties are assumed to be normal and projective.
-A ${\mathbb{Q}}$-Fano variety $X$ is a variety with at worst klt singularities such that the anti-canonical divisor $-K_{X}$ is an ample ${\mathbb{Q}}$-Cartier divisor.
+A $\mathbb{Q}$}-Fano variety $X$ is a variety with at worst klt singularities such that the anti-canonical divisor $K_{X}$ is an ample $\mathbb{Q}$}-Cartier divisor.
 A singularity $x\in X$ consists of a variety $X$ and a closed point $x\in X$.
 A singularity $x\in X$ is called klt if $X$ is klt near a neighborhood of $x$.
-A ${\mathbb{R}}$-valuation over singularity $x$ is a valuation $v\colon K(X)^{*}\rightarrow{\mathbb{R}}$ center at $x$ (namely, for all $f\in{\mathcal{O}}_{X,x}$, we have $v(f)\geqslant 0$ and $v(f)>0$ if and only if $f\in\mathfrak{m}_{x}$) and $v|_{{\mathbb{C}}^{*}}=0$.
+A $\mathbb{R}$}-valuation over singularity $x$ is a valuation $v\colon K(X)^{*}\rightarrow{\mathbb{R}}$ centered at $x$ (namely, for all $f\in{\mathcal{O}}_{X,x}$, we have $v(f)\geqslant 0$ and $v(f)>0$ if and only if $f\in\mathfrak{m}_{x}$) and $v|_{{\mathbb{C}}^{*}}=0$.
 The set of all such valuations is denoted by ${\rm Val}_{X,x}$.
@@
 K-stability, first introduced by Tian ([Tia97]) and later reformulated algebraically by Donaldson ([Don02]), is an algebro-geometric notion to characterize the existence of Kähler-Einstein metrics on Fano varieties.
-In this subsection, we recall some notions in K-stability theory that are relevant to our paper, and refer to [Xu24] for a detailed exposition of K-stability theory.
+In this subsection, we recall some notions in K-stability theory that are relevant to our paper, and refer to [Xu25] for a detailed exposition of K-stability theory.
 We say that a prime divisor $E$ is over $X$ if there exists a proper birational morphism $\mu\colon Y\rightarrow X$ such that $Y$ is normal and $E$ is a prime divisor on $Y$.
@@
 Define the $S$-invariant
-$S(E):=S(-K_{X};E)\coloneqq\frac{1}{(-K_{X})^{n}}\int_{0}^{\infty}{\rm vol}(-K_{X})-xE)\,{\rm d}x,$
+$S(E):=S(-K_{X};E)\coloneqq\frac{1}{(-K_{X})^{n}}\int_{0}^{\infty}{\rm vol}(-K_{X}-xE)\,{\rm d}x,$
 where, for the simplicity of notation, we just write ${\rm vol}(-K_{X}-xE)$ for ${\rm vol}(\mu^{*}(-K_{X})-xE)$. We have the following valuative criterion for K-semistability,
@@
 [Fuj19, Li17] A Fano variety $X$ is K-semistable if and only if $A(E)-S(E)\geqslant 0$ for every divisor $E$ over $X$.
-Another equivalent way to characterize K-semistability is via the $\delta$-invariant (also known as stability threshold), we recall $\delta(X)\coloneqq{\rm inf}_{E/X}\frac{A(E)}{S(E)}$ where the infinum is taking over all divisors $E$ over $X$ ([FO18, BJ20]).
+Another equivalent way to characterize K-semistability is via the $\delta$-invariant (also known as the stability threshold), we recall $\delta(X)\coloneqq{\rm inf}_{E/X}\frac{A(E)}{S(E)}$ where the infinum is taking over all divisors $E$ over $X$ ([FO18, BJ20]).
 Then $X$ is K-semistable if and only if $\delta(X)\geqslant 1$ by the valuative criterion.
@@
 Conjecture 2.3.
-(ODP conjecture) The second largest local volume of an $n$-dimensional klt singularity is $2(n-1)^{n}$, with equality if and only if $x\in X$ is an ordinary double point.
+(ODP conjecture) The second largest local volume of an $n$-dimensional klt singularity is $2(n-1)^{n}$.
+Moreover, the local volume is equal to $2(n-1)^{n}$ if and only if $x\in X$ is an ordinary double point.
 For later purposes, we formulate a calculus lemma and an estimate of volume.
@@
 If moreover the first strict inequality holds for $V\geq V_{1}$, then $F^{\phi}(V)>F^{\psi}(V)$ for $V\geq V_{1}$.
-As a consequence, we get $(F^{\phi})^{-1}(A)<(F^{\psi})^{-1}(A)$ for any $A\in(0,\infty)$ with strict inequality if $A>F^{\psi}(V_{1})$.
+As a consequence, we get $(F^{\phi})^{-1}(A)\leq(F^{\psi})^{-1}(A)$ for any $A\in(0,\infty)$ with a strict inequality if $A>F^{\psi}(V_{1})$.
 Proof.
-First assume that $\phi(x)$ is smooth and strictly increasing on $[0,\infty)$.
+First assume that $\phi(x)$ is smooth and strictly increasing on $0,\infty)$.
 Then its inverse function $\phi^{-1}(x)$ is also smooth, strictly increasing and satisfies (2).
@@
 So we conclude that $(F^{\psi})^{-1}(A)>(F^{\phi})^{-1}(A)$ when $A>F^{\psi}(V_{1})$.
-If $\phi$ is only piecewise smooth, we can just carry out the above argument piecewise on each interval of smoothness. The same argument applies to the second statement too. ∎
+If $\phi$ is only piecewise smooth, we can just carry out the above argument piecewise on each interval of smoothness. The same argument applies to prove the second statement too. ∎
 Corollary 2.5.
@@
 (3) ${\rm vol}(L-xE)\geqslant V-\phi(x).$
-There exists a unique solution $T=T(A_{X}(E))$ to the equation
+Then there exists a unique solution $T=T(A_{X}(E))$ to the equation
 (4) $(T-A_{X}(E))\phi(T)=\Phi(T)$
@@
 We have an estimate:
-[…]
+$0$ $\leqslant$ $A_{X}(E)-\frac{1}{V}\int_{0}^{+\infty}{\rm vol}(L-xE){\rm d}x\leqslant A_{X}(E)-\frac{1}{V}\int_{0}^{\phi^{-1}(V)}(V-\phi(x)){\rm d}x$
 $=$ $A_{X}(E)-F^{\phi}(V).$
-By the above lemma, $F^{\phi}(V)$ is a strictly increasing function that diverges to +$\infty$+ ∞ as $V\rightarrow+\infty$.
+By the above lemma, $F^{\phi}(V)$ is a strictly increasing function that diverges to $\infty$ as $V\rightarrow+\infty$.
 So there exists a unique $V^{*}$ that satisfies $A(E)=F^{\phi}(V^{*})$ which is equivalent to the equality:
@@
 Example 2.7.
-Assume that $\phi(x)=2nx^{n-1}$ and $A(E)=n-1$. The $\Phi(x)=2x^{n}$. The equation (4) has the solution $T=n$ and $\phi(T)=2n^{n}$.
+Assume that $\phi(x)=2nx^{n-1}$ and A(E)=n1A(E)=n-1. The $\Phi(x)=2x^{n}$. The equation (4) has the solution $T=n$ and $\phi(T)=2n^{n}$.
 2.2. Seshadri Constant
@@
 Since $\hat{X}$ has klt singularites, then it has rational singularities ([KM98, Theorem 5.22]), so in particular it is Cohen-Macaulay.
-Since $B=-K_{\hat{X}}+(r-1-\epsilon)E\sim_{\Phi,{\mathbb{Q}}}0$, we get $-K_{\hat{X}}\sim_{\Phi,{\mathbb{Q}}}\lambda E$ where $\lambda-r+1+\epsilon>1$.
+Since $B=-K_{\hat{X}}+(r-1-\epsilon)E\sim_{\Phi,{\mathbb{Q}}}0$, we get $K_{\hat{X}}\sim_{\Phi,{\mathbb{Q}}}\lambda E$ where $\lambda-r+1+\epsilon>1$.
 By [LZ18, Lemma 8], we have $\Phi\colon\hat{X}\rightarrow Y$ is not birational.
@@
 So we may assume ${\mathcal{K}}={\mathcal{O}}_{Y}$, then ${\mathcal{L}}={\mathcal{O}}_{Y}(-1)\cong p_{2}^{*}{\mathcal{O}}_{{\mathbb{P}}^{r-1}}(-1)$ with the projection $p_{2}\colon Y\rightarrow{\mathbb{P}}^{r-1}$ and we have the following short exact sequence
-$0\rightarrow{\mathcal{O}}_{Y}\rightarrow{\mathcal{E}}\rightarrow{\mathcal{O}}_{Y}(-1)\rightarrow 0$
+$0\rightarrow{\mathcal{O}}_{Y}\rightarrow{\mathcal{E}}\rightarrow{\mathcal{O}}_{Y}(-1)\rightarrow 0.$
 By [Har77, Prop III.6.3], we have
@@
 2.3. Minimal rational curves
-Since the celebrated work of Mori introducing the bend-and-break method for constructing rational curves on Fano manifolds ([Mor79]), the theory of rational curves has been developed extensively and has found numerous applications in algebraic geometry, especially in the classification of Fano manifolds with special properties.
+Since the celebrated work of Mori that introduces the bend-and-break method for constructing rational curves on Fano manifolds ([Mor79]), the theory of rational curves has been developed extensively and has found numerous applications in algebraic geometry, especially in the classification of Fano manifolds with special properties.
 Here we recall some basic knowledge of rational curves and refer to [Kol96] for detailed expositions.
@@
 A free rational curve is called minimal (or standard) if
-$f^{*}TX={\mathcal{O}}_{{\mathbb{P}}^{1}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)^{d-2}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus(n-d+1)}.$
-Such a morphism $f$ must be an immersion and its image $f({\mathbb{P}}^{1})$ has at worst nodal singularities. We will also be interested in its normal bundle:
+$f^{*}TX={\mathcal{O}}_{{\mathbb{P}}^{1}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}^{\oplus(n-d+1)}.$
+Such a morphism $f$ must be an immersion. We will also be interested in its normal bundle:
 $N_{f/X}:=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
@@
 An irreducible component $\mathcal{M}$} of $\mathrm{RatCurves}^{n}(X)$ is referred to as a family of rational curves on $X$.
-The anticanonical degree $\deg({\mathcal{M}})$ of the family ${\mathcal{M}}$ is defined to be $-K_{X}\cdot C$ for any curve $C$ belonging to the family.
+The anticanonical degree $\deg({\mathcal{M}})$ of the family $\mathcal{M}$} is defined to be $K_{X}\cdot C$ for any curve $C$ belonging to the family.
 This family $\mathcal{M}$} is equipped with a ${\mathbb{P}}^{1}$-bundle $p:{\mathcal{U}}\rightarrow{\mathcal{M}}$ and an evaluation morphism $q:{\mathcal{U}}\rightarrow X$.
@@
 compact).
-Note that ${\mathcal{M}}$ is a dominating family of rational curves on $X$ that has a minimal anticanonical degree, then ${\mathcal{M}}$ is a dominating locally unsplit family.
+Note that if $\mathcal{M}$} is a dominating family of rational curves on $X$ that has a minimal anticanonical degree, then $\mathcal{M}$} is a dominating locally unsplit family.
 But not all dominating locally unsplit family has a minimal anticanonical degree.
@@
 Indeed, if a general free rational curve $C=[f]$ in ${\mathcal{M}}_{x}$ is not minimal, then there are at least two ${\mathcal{O}}(2)$ summands in the splitting of $f^{*}TX$.
-We can then fix two points on the curve $C$ and bend-and-break rational curve into a non-integral (reducible) curve which would contradict the locally unsplit property.
+We can then fix two points on the curve $C$ and bend-and-break the rational curve into a non-integral (reducible) curve which would contradict the locally unsplit property.
 Following [Miy04, CD15], we set
-(6) $l_{X}:={\rm min}\{\deg{\mathcal{M}};{\mathcal{M}}\text{ is a dominating locally unsplit family of rational curves on }X\}.$
-By the above discussion, it is easy to see that (see [CD15, Remark 4.2]):
+[…]
+By the above discussion, we see that (see [CD15, Remark 4.2]):
 (7) $l_{X}$ $=$ ${\rm min}\{-K_{X}\cdot C;C\subset X\text{ is a free rational curve on }X\}$
 $=$ ${\rm min}\{-K_{X}\cdot C;C\subset X\text{ is a minimal rational curve on }X\}.$
-Example 2.9.
-To prove our main results, we will need the following important results on classification of Fano manifolds with dominating locally unsplit families of large anticanonical degrees.
+Theorem 2.9.
+Using the above notation, we have the following important classificaiton results:
   1. (1)
@@
   2. (2)
-    ([Miy04, CD15, DH17]) $l_{X}=n$ if and only if $X\cong Q^{n}$ or $X$ is the blowup of ${\mathbb{P}}^{n}$ along subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ that is contained in a hyperplane.
+    ([Miy04, CD15, DH17]) $l_{X}=n$ if and only if $X\cong Q^{n}$ or $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ that is contained in a hyperplane.
 Example 2.10 ([HM03]).
@@
   3. (3)
-    for a general point $x\in X$, there exists a rational curve which has degree 2 with respect to $K_{X}^{-1}$.
+    for a general point $x\in X$, there exists a rational curve that has degree 2 with respect to $K_{X}^{-1}$.
 There are many examples of such Fano manifolds.
@@
 Hypersurfaces of ${\mathbb{P}}^{n+1}$ of degree $n$ or $n+1$ are also examples of such Fano manifolds when $n\geqslant 3$.
-3. Warm-up: case of trivial normal bundles
-In this section, we prove the following theorem, which is a special case of Theorem 1.1.
-Theorem 3.1.
-Let $X$ be a K-semistable Fano manifold that contains a rational curve with a trivial normal bundle.
-Then $(-K_{X})^{n}\leqslant 2n^{n}$ and the equality holds if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
-The rest of this section is devoted to the proof of Theorem 3.1. We start with a proposition that proves a more general version of estimate (1) by incorporating the $\delta$-invariant.
+3. Submanifolds with normal bundles
+In this section, we will prove Theorem 1.4 by generalizing the methods from [Fuj18, LZ18], and apply it in the toric case to prove Theorem 1.2.
+We start with a proposition that proves a more general version of the estimate (1) by incorporating the $\delta$-invariant.
 Proposition 3.1.
-Assume $X$ is a $n$-dimensional Fano manifold, and $Z\subset X$ is a codimension $r$ non-singular subvariety with trivial normal bundle $N_{Z/X}={\mathcal{O}}_{Z}^{\oplus r}$, and we set $d=(-K_{X})^{n-r}\cdot Z$, then
+Let $X$ be a $n$-dimensional Fano manifold, and $Z\subset X$ be a codimension $r$ non-singular subvariety with trivial normal bundle $N_{Z/X}={\mathcal{O}}_{Z}^{\oplus r}$.
+Set $d=(-K_{X})^{n-r}\cdot Z$.
+Then we have the estimate:
 $(-K_{X})^{n}\leqslant\delta(X)^{-r}\cdot(r+1)^{r}\binom{n}{r}\cdot d=\delta(X)^{-r}\cdot(-K_{{\mathbb{P}}^{r}\times Z})^{n}.$
@@
 For $j\in{\mathbb{Z}}_{\geqslant 0}$, we use the exact sequence
-$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0$
+$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0.$
 And $H^{1}(X,kL\otimes({\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}))=\oplus H^{1}(Z,-kK_{Z})=0$ from Kodaira vanishing and the assumption that normal bundle $N_{Z/X}$ is trivial.
@@
 Proof.
-The argument is similar with [Fuj18, Theorem 2.3(2)]. We denote
+The argument is similar to the proof of [Fuj18, Theorem 2.3(2)]. We denote
 (11) $\gamma=\gamma_{Z}(L)\coloneqq{\rm max}\{t\in{\mathbb{R}}_{\geqslant 0}\mid x\in\Lambda_{Z}(L)\,\,\text{for all}\,x\in[0,t]\}.$
@@
 Then the exact sequence is given by
-$0\rightarrow H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow H^{0}(X,kL)\rightarrow H^{0}(X,kL\otimes{\mathcal{O}}_{kx_{t}Z})\rightarrow H^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow 0$
+$0\rightarrow H^{0}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow H^{0}(X,kL)\rightarrow H^{0}(X,kL\otimes{\mathcal{O}}_{kx_{t}Z})\rightarrow H^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})\rightarrow 0.$
 And $H^{i}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})=H^{i-1}(X,kL\otimes({\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{kx_{t}}))$ for $i\geqslant 2$.
@@
 Then, $h^{1}(X,kL\otimes{\mathcal{I}}_{Z}^{kx_{t}})=o(k^{n})$. On the other hand, for $j\in{\mathbb{Z}}_{\geqslant 0}$, by the exact sequence
-$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0$
+$0\rightarrow{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j+1}\rightarrow{\mathcal{O}}_{X}/{\mathcal{I}}_{Z}^{j}\rightarrow 0.$
 We have $H^{i}(X,kL\otimes{\mathcal{I}}_{Z}^{j}/{\mathcal{I}}_{Z}^{j+1})=0$ for $i\geqslant 1$ and $j\geqslant 1$ since $N_{Z/X}$ is trivial. So we get
@@
 Proof.
-When $X$ is a smooth toric Fano manifold, by the work of Chen-Fu-Hwang [CFH14, Corollary 2.5] which is partly based on the work of Araujo [Ara06], we know that there exists submanifold $Z\cong{\mathbb{P}}^{n-r}$ that is contained in $X$ with trivial normal bundle.
+When $X$ is a smooth toric Fano manifold, by the work of Chen-Fu-Hwang [CFH14, Corollary 2.5] which is partly based on the work of Araujo [Ara06], we know that there exists a submanifold $Z\cong{\mathbb{P}}^{n-r}$ for $0\leq r\leq n-1$ that is contained in $X$ with a trivial normal bundle.
 By Proposition 3.1 and $\delta(X)\geqslant 1$, we have $(-K_{X})^{n}\leqslant(-K_{{\mathbb{P}}^{r}\times Z})^{n}=(-K_{{\mathbb{P}}^{r}\times{\mathbb{P}}^{n-r}})^{n}=\binom{n}{r}\cdot(r+1)^{r}\cdot(n-r+1)^{n-r}=:c_{r}$.
-By the following lemma 3.5, for $12\leq r\leq n-$ 1, $c_{r}\leq 2n^{n}$ with equality holds if and only if $r=2$ or $r=n-1$ in which case $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
+By the following lemma 3.4, for $1\leq r\leq n-1$, $c_{r}\leq 2n^{n}$ with equality holds if and only if $r=1$ or r=n1r=n-1 and in both cases $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
 ∎
@@
 $\frac{c_{r+1}}{c_{r}}$ $=$ $\left(\frac{r+2}{r+1}\right)^{r+1}\cdot\left(\frac{n-r}{n-r+1}\right)^{n-r}=a_{r+1}/a_{n-r}$
-[…]
-Proposition 3.6.
-Let $X$ be a K-semistable Fano manifold.
-Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersed rational curve with a non-empty subset of nodal points such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}^{n-1}$.
-Then we have an estimate of the volume $(-K_{X})^{n}<2n^{n}$.
-As a consequence, if there exists a minimal rational curve with a trivial normal bundle on $X$ and $(-K_{X})^{n}=2n^{n}$, then the minimal rational curve has no nodal points and $X$ must be isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
-Proof.
-The last statement follows from the strict volume inequality and Theorem 1.5.
-So we just need to prove the inequality $(-K_{X})^{n}<2n^{n}$.
-Denote the ideal sheaf of the nodal curve $C=f({\mathbb{P}}^{1})$ in $X$ by ${\mathcal{I}}={\mathcal{I}}_{C}$.
-Consider the blow up $\mu:\tilde{X}\rightarrow X$ along $C$.
-By using the same argument as in the proof of Proposition 3.2, we get the estimate:
-${\rm vol}({\mu}^{*}(-K_{X})-xE)$$\geqslant$ $(-K_{X})^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{h^{0}(X,L^{k}\otimes{\mathcal{O}}_{kxC})}{k^{n}/n!}$
-$\geqslant$ $(-K_{X})^{n}-\mathop{{\rm lim\,sup}}_{k\rightarrow+\infty}\frac{\mathop{\sum}_{y=0}^{xk-1}h^{0}(C,kL\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})}{k^{n}/n!}.$
-Since $C$ is a local complete intersection, its conormal sheaf $N^{\vee}_{C/X}={\mathcal{I}}/{\mathcal{I}}^{2}$ is locally free and ${\mathcal{I}}^{y}/{\mathcal{I}}^{y+1}={\rm Sym}^{y}(N_{C/X}^{\vee})$.
-So we get:
-$h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})=h^{0}(C,L^{k}\otimes{\rm Sym}^{y}({\mathcal{I}}/{\mathcal{I}}^{2}))\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(dk)\otimes{\rm Sym}^{y}(f^{*}({\mathcal{I}}/{\mathcal{I}}^{2}))).$
-There is an inclusion ${\mathcal{F}}:=f^{*}({\mathcal{I}}/{\mathcal{I}}^{2})\hookrightarrow N_{f/X}^{\vee}=(f^{*}TX/T{\mathbb{P}}^{1})^{\vee}$.
-We know that $\deg(N^{\vee}_{f/X})=2-d$ and $\deg({\mathcal{F}})=\deg(K_{X})-\deg\omega_{C}=-d-2m+2$ where $d=-K_{X}\cdot C$.
-Since $N^{\vee}_{f/X}$ is trivial, $d=2$ and ${\mathcal{F}}={\mathcal{O}}(-a_{1})\oplus{\mathcal{O}}(-a_{2})\oplus\cdots\oplus{\mathcal{O}}(-a_{n-1})$ with $a_{1}\geq a_{2}\geq\cdots\geq a_{n-1}\geq 0$ and $\sum_{i=1}^{n-1}a_{i}=2m$.
-It is then easy to get the estimate:
-(12) $h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-2)\oplus{\mathcal{O}}^{\oplus(n-2)}))$
-or
-(13) $h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leqslant h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-1)^{\oplus 2}\oplus{\mathcal{O}}^{\oplus(n-3)})).$
-For the first case (12), we get:
-(14) $\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leq\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-2i+1)a_{y,i}$
-where $a_{y,i}=\binom{n-3+y-i}{n-3}$.
-If $x\leqslant 1$, then by using the explicit expression of binomial coefficient, the right-hand-side of (14) is equal to $b_{1}k^{n}/n!+O(k^{n-1})$ where $b_{1}$ is given by the following integral (see (19) for a derivation of a more complicated formula):
-$b_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-2z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z=\frac{2n!}{(n-2)!}\int_{0}^{x}(1-z)(x-z)^{n-2}{\rm d}z$
-$=$ $2nx^{n-1}-2x^{n}<2nx^{n-1}.$
-When $x>1$, the right-hand-side of (14) splits into the sum: $\sum_{y=0}^{k}\sum_{i=0}^{y}+\sum_{y=k+1}^{xk-1}\sum_{i=0}^{k}$ and the right-hand-side of (14) is equal to $b_{2}k^{n}/n!+O(k^{n-1})$ with
-$b_{2}$ $=$ $(2n-2)+n!\int_{1}^{x}{\rm d}t\int_{0}^{1}(2-2z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z$
-$=$ $(2n-2)+\frac{2n!}{(n-2)!}\int_{0}^{1}(1-z)((x-z)^{n-2}-(1-z)^{n-2}){\rm d}z$
-$=$ $\frac{2n!}{(n-2)!}\int_{0}^{1}u(x-1+u)^{n-2}{\rm d}u+(2n-2)-(2n-2)$
-$=$ $2\displaystyle 2(nx^{n-1}-x^{n}+(x-1)^{n})<2(n-1)x^{n-1}+$ 2 .
-We set $\psi=2nx^{n-1}$ and
-(15) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 1\\ 2(n-1)x^{n-1}+2&\text{if }\;x>1,\end{array}\right.$
-so that $\phi(x)<\psi(x)$.
-It is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and the strict inequality holds when $V>2n$.
-So we can apply Lemma 2.4 to conclude that $F^{\phi}(V)\geq F^{\psi}(V)$ and the strict inequality holds for $V>2n$.
-From Lemma 2.4 and Corollary 2.5, we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$.
-In the second case (13), we get:
-$\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}^{y}/{\mathcal{I}}^{y+1})\leq\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-i+1)(i+1)\binom{n-4+y-i}{n-4}.$
-Estimating as before, when $x\leqslant 2$, the right-hand-side of (14) is equal to $c_{1}k^{n}/n!+O(k^{n-1})$. The coefficient $c_{1}$ is equal to:
-$c_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-z)z\frac{(t-z)^{n-4}}{(n-4)!}{\rm d}z$
-$=$ $\frac{n!}{(n-3)!}\int_{0}^{x}(2-z)z(x-z)^{n-3}{\rm d}z=2x^{n-1}(n-x)<2nx^{n-1}.$
-When $x\geqslant 2$, the right-hand-side of (14) splits into the sum: $\sum_{y=0}^{2k}\sum_{i=0}^{y}+\sum_{y=2k+1}^{xk-1}\sum_{i=0}^{2k}$ and the right-hand-side of (14) is equal to $c_{2}k^{n}/n!+O(k^{n-1})$ with the coefficient $c_{2}$ is equal to:
-$c_{2}$ $=$ $2^{n}(n-2)+n!\int_{2}^{x}{\rm d}t\int_{0}^{2}(2-z)z\frac{(t-z)^{n-4}}{(n-4)!}{\rm d}z$
-$=$ $2^{n}(n-2)+\frac{n!}{(n-3)!}\int_{0}^{2}(2-z)z((x-z)^{n-3}-(2-z)^{n-3}){\rm d}z$
-$=$ $\frac{n!}{(n-3)!}\int_{0}^{2}(2-z)z(x-z)^{n-3}{\rm d}z$
-$=$ $2nx^{n-1}-2x^{n}+(x-2)^{n}+n(x-2)^{n-1}<(2n-1)x^{n-1}$
-$<$ $(2n-1)x^{n-1}+2^{n-1}.$
-We set $\psi(x)=2nx^{n-1}$ and
-(16) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 2\\ (2n-1)x^{n-1}+2^{n-1}&\text{if }\;x\geq 2.\end{array}\right.$
-It is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and strict inequality holds for $V>2^{n}\cdot n$.
-Then by Lemma 2.4 we get $F^{\phi}(V)\geq F^{\psi}(V)$ with strict inequality when $V>2^{n}\cdot n$.
-So we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$.
-∎
-Proof of Theorem 3.1.
-If $X$ contains a rational curve with non-empty nodal points, then by Proposition 3.6, we have $(-K_{X})^{n}<2n^{n}$.
-So it suffices to consider the case when $C$ is an embedded rational curve, in which case the result follows immediately from Theorem 1.5.
+where $a_{k}=(\frac{k+1}{k})^{k}=(1+k^{-1})^{k}$ is an strict increasing sequence for $k\geq 1$ (that converges to $e$ as $k\rightarrow+\infty$).
+So we get $c_{r+1}/{c_{r}}<1$ when r+1<nrr+1<n-r (or equivalently $2r+1<n)$.
+The statement then follows easily.
 ∎
 Proposition 3.5.
-Suppose $X$ is a singular $n$-dimensional K-semistable toric ${\mathbb{Q}}$-Fano variety.
+Let $X$ be a singular $n$-dimensional K-semistable toric $\mathbb{Q}$}-Fano variety.
 When $n=2$, then $(-K_{X})^{2}\leqslant\frac{9}{2}$.
@@
 Proof.
-Suppose $x\in X$ lies in the singular locus and $(X,x)$ is a toric singularity.
-It was proved in [MS24, Theorem 2] that if $(X,x)$ is a $n$-dimensional ${\mathbb{Q}}$-Gorenstein toric singularity, then
+Suppose $x\in X$ lies in the singular locus and $(X,x)$ is a toric singularity. It was proved in [MS24, Theorem 2] that if $(X,x)$ is a $n$-dimensional $\mathbb{Q}$}-Gorenstein toric singularity, then
   1. (1)
@@
 Theorem 3.6.
-The second largest volume of $n$-dimensional K-semistable toric ${\mathbb{Q}}$-Fano varieties is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
+The second largest volume of $n$-dimensional K-semistable toric $\mathbb{Q}$}-Fano varieties is $2n^{n}$ and it is obtained only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
 Proof.
@@
 Since the inequality is strict, it is expected that the equality of Theorem 1.1 would hold only if $X$ is a smooth Fano manifold.
-We conclude this section by a convex-geometric application.
+As mentioned in the Introduction, Theorem 3.6 finds an application in Arakelov geometry on the sharp bound of height of arithmetic Fano varieties whose complexification is K-semistable.
+We refer our readers to [AB24, Ber25] for more detailed explanation and possible arithmetic applications.
+Theorem 3.8.
+Andreasson-Berman’s conjecture holds for the toric canonical model of a toric Fano varieties.
+Namely, for a $n$-dimensional K-semistable toric Fano variety and let $\mathcal{X}$} be the canonical model of $X$ over $\mathbb{Z}$}, then for any volume-normalized continous metric $\left\|\cdot\right\|$ on $K_{X}$ with positive curvature current, we have $h_{\left\|\cdot\right\|}({\mathcal{X}},-K_{{\mathcal{X}}})\leq h_{\left\|\cdot\right\|_{\text{FS}}}({\mathbb{P}}^{n}_{{\mathbb{Z}}},-K_{{\mathbb{P}}^{n}_{{\mathbb{Z}}}})$ where $K_{{\mathbb{P}}^{n}_{{\mathbb{C}}}}$ is endowed with the volume normalized Fubini-Study metric.
+Proof.
+It follows directly from Theorem 3.6 and [AB24, Lemma 3.8]. ∎
+We conclude this section with another application in convex geometry.
 Corollary 3.9.
-Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with barycenter at $0\in{\mathbb{R}}^{n}$.
-Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $[0,2]\times(n\Delta_{n-1})$.
+Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with its barycenter at $0\in{\mathbb{R}}^{n}$.
+Assume that $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $0,2]\times(n\Delta_{n-1})$.
 Proof.
@@
 4. Volume estimates and minimal rational curves
-4.1. Proof of Theorem 1.4 and Theorem 1.1
 The calculation in the previous section is based on the test of K-stability via standard blow-ups.
-In this section, we prove Theorem 1.4 and Theorem 1.1 by considering the test of K-stability via weighted blow-ups along (Zariski open sets of) minimal rational curves.
-Proof of Theorem 1.4.
-Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ be minimal rational curve, or equivalently an immersed rational curve that satisfies $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ with $12\leq d\leq n+$ 1 and $d=f^{*}(-K_{X})\cdot{\mathbb{P}}^{1}=(-K_{X})\cdot f({\mathbb{P}}^{1})$.
+In this section, we prove Theorem 1.1 by carrying out a test of K-stability via a particular weighted blow-ups along minimal rational curves (see Example (4.16) for the motivating example).
+We also need to overcome the difficulty of possible singularities on the rational curve.
+4.1. Two divisorial valuations: $v_{1}$ and $v_{2}$
+Let $f:{\mathbb{P}}^{1}\rightarrow X$ be a minimal rational curve.
+In other words, it is an immersed rational curve that satisfies $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ with $2\leq d\leq n+1$ and $d=f^{*}(-K_{X})\cdot{\mathbb{P}}^{1}=(-K_{X})\cdot f({\mathbb{P}}^{1})$.
 The (relative) normal bundle has the splitting:
-$N_{f/X}=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
+(12) $N_{f/X}=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
 Correspondingly, we have the splitting of the conormal bundle:
 $N^{\vee}_{f/X}={\mathcal{O}}(-1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
-Set $Z=f({\mathbb{P}}^{1})$ which is an irreducible curve with possible nodal singularities.
-Fix a regular point $x\in Z$, there exists a Zariski open neighborhood $U$ and coordinates $\{z_{i};i=1,\dots,n\}$ such that
+Set $Z=f({\mathbb{P}}^{1})$ which is an irreducible curve with possibly a finite number of singularities $Z^{\mathrm{sing}}=\{p_{1},\dots,p_{k}\}$ and $f:{\mathbb{P}}^{1}\rightarrow Z$ is the normalization of $Z$.
+Set $f^{-1}(p_{i})=\{t_{i}^{j};1\leq j\leq m_{i}\}$ to be the inverse image of the singularity $p_{i}$ under $f$ so that $f^{-1}(Z^{\mathrm{sing}})=\cup_{i}f^{-1}(p_{i})$.
+For any point $t\in{\mathbb{P}}^{1}$, we first construct a complex analytic coordinate chart near $f(t)\in X$ in the following way, which is adapted to the splitting (12).
+Choose a disk $B_{t}:=B_{t}(\epsilon)\subset{\mathbb{P}}^{1}$ of radius $\epsilon>0$ (with respect to the round spherical metric) centered at $t$, such that $f:B_{t}\rightarrow X$ is an embedding.
+Since $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersion and $f^{-1}(Z^{\mathrm{sing}})$ is a finite set, it is easy to see that we can choose $\epsilon$ sufficiently small such that $B_{t}$ exists for all $t\in{\mathbb{P}}^{1}$.
+Moreover, by choosing a uniform $0<\epsilon\ll 1$, we can choose an open neighborhood $U=U_{t}$ of $f(t)$ and complex analytic coordinates ${\mathbf{z}}={\mathbf{z}}_{t}:=\{z_{1},\dots,z_{n}\}$ on $U$ such that
   1. (1)
-    $Z\cap U=\{z_{1}=\cdots=z_{n-1}=0\}$;
+    $z_{i}(f(t))=0$ for $i=1,\dots,n$; ${\mathbf{z}}(U_{t})=\{|z_{i}|<\epsilon;i=1,\dots,n\}$;
   2. (2)
-    For any $21\leq i\leq d-$ 2, ${\rm d}z_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}$ over $Z\cap U$;
+    $f(B_{t})\cap U_{t}=\{z_{1}=\cdots=z_{n-1}=0\}$;
   3. (3)
-    For any $d-1\leq j\leq n-1$, ${\rm d}z_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}$ over $Z\cap U$.
-For any $\ell\in{\mathbb{Z}}_{>0}$, we can define a divisorial valuation $v={\rm ord}_{E}$: for any $f\in{\mathcal{O}}(U)$,
-(17) $v(f)={\rm min}\{|I|+\ell|J|;f=\sum_{I,J}{a_{IJ}}z^{\prime I}z^{\prime\prime J},a_{IJ}=a_{IJ}(z_{n})\neq 0\}$
-where $z^{\prime}=\{z_{1},\dots,z_{d-2}\}$ (resp.
-$z^{\prime\prime}=\{z_{d-1},\cdots,z_{n-1}\}$), and $z^{\prime I}=z_{1}^{i_{1}}\cdots z_{d-2}^{i_{d-2}}$ for $I=(i_{1},\dots,i_{d-2})\in{\mathbb{N}}^{d-2}$ with $|I|=i_{1}+\cdots+i_{d-2}$ (resp.
-$z^{\prime\prime J}=z_{d-1}^{j_{1}}\cdots z_{n-1}^{j_{n-d+1}}$ for $J=(j_{1},\cdots,j_{n-d+1})\in{\mathbb{N}}^{n-d+1}$ with $|J|=j_{1}+\cdots+j_{n-d+1}$).
-When restricted to $U$, the valuation $v$ corresponds to the exceptional divisor $E$ of the weighted blowup of $X\cap U$ along $Z\cap U$ with weights $(\underbrace{1,\dots,1}_{d-2},\underbrace{\ell,\cdots,\ell}_{n-d+1})$ (See [QR12] for more a general set-up of weighted blowups).
-Set ${\mathcal{I}}_{x}={\mathcal{I}}_{x}({\rm ord}_{E})=\{f\in{\mathcal{O}}_{X};v(f)\geq x\}$.
-Then we have the exact sequence:
+    For any $1\leq i\leq d-2$, $f^{*}{\rm d}z_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}\hookrightarrow N^{\vee}_{f/X}$ over $B_{t}$;
+  4. (4)
+    For any $d-1\leq j\leq n-1$, $f^{*}{\rm d}z_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}\hookrightarrow N^{\vee}_{f/X}$ over $B_{t}$.
+We now define a divisorial valuation $v={\rm ord}_{E}$ whose center over $X$ is $Z$.
+Pick any rational function $\psi$ on $X$.
+We assume that $\psi$ is defined on a Zariski open set $W$ such that $W\cap Z\neq\emptyset$.
+Pick any point $t\in f^{-1}(W\cap Z)$.
+The restriction $\psi$ to $U=U_{t}$ becomes holomorphic near $0\in U$.
+Expand the restriction $f|U$ near $0\in U$:
+(13) $\psi({\mathbf{z}})=\sum_{I,J}{b_{IJ}}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J}$
+where ${\mathbf{z}}^{\prime}=\{z_{1},\dots,z_{d-2}\}$ (resp.
+${\mathbf{z}}^{\prime\prime}=\{z_{d-1},\cdots,z_{n-1}\}$), and ${\mathbf{z}}^{\prime I}=z_{1}^{i_{1}}\cdots z_{d-2}^{i_{d-2}}$ for $I=(i_{1},\dots,i_{d-2})\in{\mathbb{N}}^{d-2}$ with $|I|=i_{1}+\cdots+i_{d-2}$ (resp.
+${\mathbf{z}}^{\prime\prime J}=z_{d-1}^{j_{1}}\cdots z_{n-1}^{j_{n-d+1}}$ for $J=(j_{1},\cdots,j_{n-d+1})\in{\mathbb{N}}^{n-d+1}$ with $|J|=j_{1}+\cdots+j_{n-d+1}$).
+We then define: for $\ell=1$ or $\ell=2$
+(14) $v_{\ell}(\psi)={\rm min}\{|I|+\ell|J|;\psi=\sum_{I,J}{b_{IJ}}{\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J},b_{IJ}=b_{IJ}(z_{n})\not\equiv 0\}.$
+We claim that this is independent of $t\in f^{-1}(W\cap Z)$.
+Since $Z\setminus W$ is a finite set, $W\cap Z$ is connected.
+So we just need to verify the claim locally.
+Assume that $t^{\prime}$ is near $t$ so that $x^{\prime}=f(t^{\prime})$ is near $x=f(t)$ and we can find a sequence of points $\{t_{i};i=1,\dots,N\}$ and small discs $B_{t_{i}}(\epsilon_{1})\subset B_{t_{i}}(\epsilon)=B_{t_{i}}$ with $1<\epsilon_{1}\ll\epsilon$ such that $B_{t_{0}}(\epsilon_{1})=B_{t}(\epsilon_{1})$, $B_{t_{N}}(\epsilon_{1})=B_{t^{\prime}}(\epsilon_{1})$ and $B_{t_{i}}(\epsilon_{1})\cap B_{t_{i+1}}(\epsilon_{1})\supset B_{t_{i}}(\epsilon_{1}/3)\cup B_{t_{i+1}}(\epsilon_{1}/3)$.
+In particular, this implies such that $U_{t_{i}}\cap U_{t_{i+1}}\neq\emptyset$.
+For each $1\leq i\leq N-1$, the transition function on $U_{t_{i}}\cap U_{t_{i+1}}$ from $(U_{t_{i}},{\mathbf{z}}={\mathbf{z}}_{t_{i}})$ and $(U_{t_{i+1}},\hat{{\mathbf{z}}}:={\mathbf{z}}_{t_{i+1}})$ is of the form:
+(15) $z_{k}$ $=$ $a_{k}(\hat{z}_{n})\hat{z}_{k}+R_{2},1\leq k\leq n-1,\quad z_{n}=g_{n}(\hat{z}_{n})+R_{1}$
+where $a_{k}(\hat{z}_{n})$ is a non-zero holomorphic transition function of ${\mathcal{O}}_{{\mathbb{P}}^{1}}(-1)$ for $1\leq k\leq d-2$ and of ${\mathcal{O}}_{{\mathbb{P}}^{1}}$ for $d-1\leq k\leq n-1$, $g_{n}(\hat{z}_{n})$ is an invertible holomorphic function and $R_{2}$ (resp.
+$R_{1}$) are holomorphic functions whose expansion consists only of terms of degree at least 2 (resp.
+at least 1) in $\hat{z}_{1},\dots,\hat{z}_{n-1}$.
+Assume $\psi|_{U_{t_{i}}}$ has the expansion (13) centered at $t_{i}\in U_{t_{i}}$ with all terms of weighted degree $|I|+\ell|J|$ at least $y$, i.e.
+the right-hand-side of (14) is equal to $y$.
+Then
+$\psi({\mathbf{z}})$ $=$ $\sum_{|I|+\ell|J|\geq y}b_{IJ}(g_{n}(\hat{z}_{n})+R_{1})(\mathbf{a}^{\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime}+R_{2})^{I}(\mathbf{a}^{\prime\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime\prime}+R_{2})^{J}$
+$=$ $\sum_{|I|+\ell|J|=y}b_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime})^{I}(\mathbf{a}^{\prime\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime\prime})^{J}+\sum_{|I|+\ell|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime})^{I}R^{\prime J}_{2}+R^{w}_{y+1}$
+$=$ $\sum_{K,L}\hat{b}_{KL}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime K}\hat{{\mathbf{z}}}^{\prime\prime L}$
+where $(\mathbf{a}^{\prime}\hat{\mathbf{z}}^{\prime})^{I}=\prod_{k=1}^{d-2}a_{k}^{i_{k}}z_{k}^{i_{k}}$, $(\mathbf{a}^{\prime\prime}\hat{\mathbf{z}}^{\prime\prime})^{J}=\prod_{k=d-1}^{n-1}a_{k}^{j_{k}}z_{k}^{j_{k}}$, $R^{\prime}_{2}$ consists of terms of degree at least $2$ in $(\hat{z}_{1},\dots,\hat{z}_{d-2})$, $R^{w}_{y+1}$ consists of terms of weighted degree $|I|+\ell|J|$ at least $y+1$ and the last expression is the expansion of $f|U_{t_{i+1}}$ near the center $t_{i+1}\in U_{t_{i+1}}$.
+It is then easy to see that for $\ell=1$ or $\ell=2$,
+${\rm min}\{|K|+\ell|L|;\psi=\sum_{K,L}\hat{b}_{KL}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime K}\hat{{\mathbf{z}}}^{\prime\prime L},\hat{b}_{KL}\not\equiv 0\}$
+is at least the quantity on the right-hand-side of (14).
+Since we can switch the role of $t_{i}$ and $t_{i+1}$, we know that the two quantities are actually equal.
+By passing from $t_{0}=t$ to $t_{N}=t^{\prime}$, we conclude that for $\ell=1,2$, $v_{\ell}(f)$ does not depend on $t$ and $t^{\prime}$.
+Remark 4.1.
+Definition of quasi-monomial valuations by using analytic expansions under complex analytic coordinates has also recently appeared in [MP25, 2.3.1].
+In particular, a similar discussion about the independence of local coordinate charts is contained in [MP25, Proposition 2.3.4].
+When restricted to each $U=U_{t}$, the valuation $v_{\ell}$ corresponds to the exceptional divisor $E$ of the weighted blowup of $U_{t}$ along $f(B_{t})\cap U_{t}$ with weights
+$(1^{\oplus(d-2)},\ell^{\oplus(n-d+1)})=\left\{\begin{array}[]{ll}(\underbrace{1,\dots,1}_{n-1})&\text{for }\ell=1,\\ (\underbrace{1,\dots,1}_{d-2},\underbrace{2,\cdots,2}_{n-d+1})&\text{for }\ell=2\end{array}\right$.
+(See [QR12] for more a general set-up of weighted blowups). In fact, for $\ell=1$ the valuation $v_{1}$ corresponds to the symbolic blowup of $X$ along $Z$.
+Fix $\ell\in\{1,2\}$ and set ${\mathcal{I}}_{\lambda}={\mathcal{I}}_{\lambda}(v_{\ell})=\{f\in{\mathcal{O}}_{X};v_{\ell}(f)\geq\lambda\}$.
+Choose $\lambda=xk$ for $x\in{\mathbb{Q}}_{>0}$.
+Then similar to the proof of Proposition 3.1, we have the exact sequence:
 $0\rightarrow H^{0}(X,L^{k}\otimes{\mathcal{I}}_{xk})\rightarrow H^{0}(X,L^{k})\rightarrow H^{0}({X},L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk})\rightarrow\cdots,$
@@
 (16) $h^{0}(X,L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk})$ $\leq$ $\sum_{y=0}^{xk-1}h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1}).$
-For any $[s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, we know that with respect to a local trivialization of $L^{k}$, the section $s$ is represented locally over $U$ by a holomorphic function with the following expression modulo terms with $|I|+\ell|J|>y$,
-$s$ $=$ $\sum_{|I|+\ell|J|=y}a_{IJ}z^{\prime I}z^{\prime\prime J}=\sum_{k=0}^{\lfloor y/\ell\rfloor}\sum_{|I|=y-\ell k,|J|=k}a_{IJ}z^{\prime I}z^{\prime\prime J}$
-$=$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor+m(\ell-1)\\ |I|=y-\ell\lfloor y/\ell\rfloor+m\ell\\ |J|=\lfloor y/\ell\rfloor-m\end{subarray}}a_{IJ}z^{\prime I}z^{\prime\prime J}.$
-Assume that $W$ is another Zariski open set such that $W\cap Z$ is non-empty and has at most one nodal point.
-In particular, there are at most two irreducible components of $W\cap Z$.
-Fix one irreducible component $Z_{1}$ which is smooth, we can choose coordinates (after possibly shrinking $W$) such that
-  1. (1)
-    $W\cap Z_{1}=\{w_{1}=\cdots=w_{n-1}=0\}$
-  2. (2)
-    For any $21\leq i\leq d-$ 2, ${\rm d}w_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}$ over $W\cap Z_{1}$.
-  3. (3)
-    For any $d-1\leq j\leq n-1$, ${\rm d}w_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}$ over $W\cap Z_{1}$.
-Near $U\cap W\cap Z$, the transition functions from $\mathbf{z}$ and $\mathbf{w}$ must be of the form:
-$z_{k}=a_{k}(w_{n})w_{k}+R_{2},\quad 1\leq k\leq n-1$
-where $a_{k}(w_{n})$ is a transition function of ${\mathcal{O}}_{{\mathbb{P}}^{1}}(-1)$ for $21\leq k\leq d-$ 2 and of ${\mathcal{O}}_{{\mathbb{P}}^{1}}$ for $d-1\leq k\leq n-1$, and $R_{2}$ consists of terms of degree at least 2 in $w_{1},\dots,w_{n-1}$.
-So near $W\cap Z_{1}$, we have $s=\sum_{|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor}b_{IJ}(w_{n})w^{\prime I}w^{\prime\prime J}+R$ where $R$ are terms of (unweighted) total degree in $(w^{\prime},w^{\prime\prime})$ strictly greater than $y-(\ell-1)\lfloor y/\ell\rfloor$.
-We can deal with the other possible component of $W\cap Z$ using the same argument.
-So we see that $s$ induces a well-defined section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor}({\mathcal{O}}^{\oplus(n-d+1)}))$.
-This defines a linear map of vector spaces:
-$H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\rightarrow H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor}({\mathcal{O}}^{\oplus(n-d+1)})),$
-whose kernel consists of elements $s\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ that has expansion ($m\geq 1$ now):
-$s=\sum_{m=1}^{\lfloor y/\ell\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-(\ell-1)\lfloor y/\ell\rfloor+m(\ell-1)\\ |I|=y-\ell\lfloor y/\ell\rfloor+m\ell\\ |J|=\lfloor y/\ell\rfloor-m\end{subarray}}a_{IJ}z^{\prime I}z^{\prime\prime J},$
-which by the same argument as above induces a section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor+\ell}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-1}({\mathcal{O}}^{\oplus(n-d+1)}))$.
+4.2. Singular minimal rational curves with trivial normal bundles
+Our goal in this subsection is to prove the following gap result when there is an immersed rational curve with a trivial bundle and a singular image set.
+Proposition 4.2.
+Let $X$ be a K-semistable Fano manifold.
+Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersed rational curve with a non-empty subset of singular points such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}^{n-1}$.
+Then we have an estimate of the volume $(-K_{X})^{n}<2n^{n}$.
+As a consequence, if there exists a minimal rational curve with a trivial normal bundle on $X$ and $(-K_{X})^{n}=2n^{n}$, then the minimal rational curve has no singular points and $X$ must be isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
+The rest of this subsection is devoted to proving the above proposition.
+The last statement follows from the strict volume inequality and Theorem 1.4.
+So we just need to prove the inequality $(-K_{X})^{n}<2n^{n}$ when $f$ is not an embedding.
+We use the same notation as the previous subsection but now specialize to the case when $\ell=1$ and $d=2$.
+For any $s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, choose a local generator $\sigma_{t}$ of $L^{k}$ over $U_{t}$, the section $[s]$ is represented uniquely over $U_{t}$ by a holomorphic function with the following expression (note that, since $d=2$, we get d2=0d-2=0 so that $(z_{1},\dots,z_{n-1},z_{n})=({\mathbf{z}}^{\prime\prime},z_{n})$):
+$\frac{s_{t}}{\sigma_{t}}$ $=$ $\sum_{|J|=y}b^{t}_{J}(z_{n}){\mathbf{z}}^{\prime\prime J}$
+which we call the initial term representation of $s$ over $U_{t}$. Set
+$\gamma_{t}([s])=\sigma_{t}\cdot\sum_{|J|=y}b^{t}_{J}(z_{n})({\rm d}{\mathbf{z}}^{\prime\prime})^{\otimes J}$
+which defines a holomorphic section of $L^{k}\otimes\mathrm{Sym}^{y}({\mathcal{O}}^{\oplus(n-1)})$ over $B_{t}$. It is easy to check that $\gamma_{t}([s])$ patches together to define a global section
+$\gamma([s])\in H^{0}({\mathbb{P}}^{1},L^{k}\otimes\mathrm{Sym}^{y}({\mathcal{O}}^{\oplus n-1}))=H^{0}({\mathbb{P}}^{1},L^{k})\otimes{\mathbb{C}}^{\binom{n+y-2}{y}}$.
+In other words, we know that for each $J=(j_{1},\dots,j_{n-1})$ with $|J|=j_{1}+\cdots+j_{n-1}=y$, $b_{J}^{t}(z_{n})({\rm d}{\mathbf{z}}^{\prime\prime})^{\otimes J}$ patches together to become a section $s_{J}$ in $H^{0}({\mathbb{P}}^{1},L^{k})$.
+Now assume that $t\neq t^{\prime}\in{\mathbb{P}}^{1}$ satisfy $f(t)=f(t^{\prime})=x\in f({\mathbb{P}}^{1})=Z\subset X$.
+Let $C=f(B_{t})$ and $C^{\prime}=f(B_{t^{\prime}})$ be the two distinct locally smooth branches of $Z$ that pass through $x$.
+Let $(U=U_{t},{\mathbf{z}}={\mathbf{z}}_{t})$ and $(\hat{U}=U_{t^{\prime}},\hat{{\mathbf{z}}}=z_{t^{\prime}})$ be coordinate charts centered at $t$ and $t^{\prime}$ respectively in Subsection 4.1.
+For each $1\leq k\leq n$, one can write:
+$z_{k}$ $=$ $g_{k1}(\hat{z}_{n})\hat{z}_{1}+\cdots+g_{k,{n-1}}(\hat{z}_{n})\hat{z}_{n-1}+\hat{z}_{n}^{m_{k}}h_{k}(\hat{z}_{n})+R_{2},$
+where $g_{kj}(\hat{z}_{n})$ and $h_{k}(\hat{z}_{n})$ are holomorphic functions of $\hat{z}_{n}$ with $h_{k}(0)\neq 0$ if $m_{k}>0$, and $R_{2}$ consists of terms of degree at least $2$ in $\hat{z}_{1},\dots,\hat{z}_{n-1}$.
+Without the loss of generality, assume that $m_{1}\geq m_{2}\geq\cdots\geq m_{n-1}\geq 0$.
+Then we must have $m_{1}>0$.
+Otherwise, we have $C^{\prime}=\{\hat{z}_{1}=\cdots=\hat{z}_{n-1}=0\}\subseteq C=\{z_{1}=\cdots=z_{n-1}=0\}$ which is not possible.
+Assume that $1\leq q\leq n-1$ satisfies $m_{q}>0$ and $m_{q+1}=0$ if q<n1q<n-1.
+For any $q+1\leq k\leq n-1$, we have ${\rm d}z_{k}|_{T_{x}X}=\sum_{j=1}^{n-1}g_{kj}(0)d\hat{z}_{j}|_{T_{x}X}$ which implies that the matrix $(g_{kj}(0))^{q+1\leq k\leq n-1}_{1\leq j\leq n-1}$ has rank equal to n1qn-1-q.
+For the section $s_{J}$ obtained from above, we get:
+$b^{t}_{J}(z_{n})z_{1}^{j_{1}}z_{2}^{j_{2}}\cdots z_{n-1}^{j_{n-1}}$ $=$ $b^{t}_{J}(z_{n})(\prod_{k=1}^{q}\hat{z}_{k}^{m_{k}}h_{k}(\hat{z}_{n})^{j_{k}}+R_{1})\cdot\prod_{k=q+1}^{n-1}\left(\sum_{j=1}^{n-1}g_{kj}(\hat{z}_{n})\hat{z}_{j}+R_{2}\right)^{j_{k}}$
+which should be of the form $b_{J}^{t^{\prime}}(\hat{z}_{n})\hat{z}_{1}^{j_{1}}\hat{z}_{2}^{j_{2}}\cdots\hat{z}_{n-1}^{j_{n-1}}+R_{|J|+1}$, since by construction $f^{*}{\rm d}z_{j}=f^{*}d\hat{z}_{j}$ on $U_{t}\cap U_{t^{\prime}}$ which is the restrction to $U_{t}\cap U_{t^{\prime}}$ of a global non-zero section of the $j$-th factor of ${\mathcal{O}}^{\oplus(n-1)}=N_{f/X}^{\vee}$.
+This forces $b^{t}_{J}(z_{n})$ to vanish at $z_{n}=0$ up to order at least $j_{1}+\cdots+j_{q}\geq j_{1}$, and moreover $\hat{z}_{n}^{m_{n}}h_{n}(\hat{z}_{n})\equiv 0$.
+So we see that $s_{J}\in H^{0}({\mathbb{P}}^{1},{\mathcal{O}}(-1)^{j_{1}}\otimes{\mathcal{O}}^{j_{2}}\otimes\cdots\otimes{\mathcal{O}}^{j_{n-1}})$.
+So we get the estimate:
+(17) $h^{0}(C,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\leq h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes\mathrm{Sym}^{y}({\mathcal{O}}(-1)\oplus{\mathcal{O}}^{(n-2)})).$
+By (16) and (17), we get:
+(18) $\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ $\leq$ $\sum_{y=0}^{xk-1}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-1)\oplus{\mathcal{O}}^{\oplus(n-2)}))$
+$\leq$ $\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-i+1)a_{y,i}$
+where $a_{y,i}=\binom{n-3+y-i}{n-3}$.
+If $x\leqslant 2$, then by using the explicit expression of binomial coefficient, the right-hand-side of (18) is equal to $b_{1}k^{n}/n!+O(k^{n-1})$ where $b_{1}$ is given by the following integral (see (22) for a derivation of a more complicated formula):
+$b_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z=\frac{n!}{(n-2)!}\int_{0}^{x}(2-z)(x-z)^{n-2}{\rm d}z$
+$=$ $2nx^{n-1}-x^{n}<2nx^{n-1}.$
+When $x>2$, the right-hand-side of (18) splits into the sum: $\sum_{y=0}^{2k}\sum_{i=0}^{y}+\sum_{y=2k+1}^{xk-1}\sum_{i=0}^{2k}$ and the right-hand-side of (18) is equal to $b^{\prime}_{1}k^{n}/n!+O(k^{n-1})$ with
+$b^{\prime}_{1}$ $=$ $n!\int_{0}^{2}{\rm d}t\int_{0}^{t}(2-z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z+n!\int_{2}^{x}{\rm d}t\int_{0}^{2}(2-z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z$
+$=$ $\frac{n!}{(n-2)!}\int_{0}^{2}(2-z)(x-z)^{n-2}{\rm d}z$
+$=$ $2nx^{n-1}-x^{n}+(x-2)^{n}<(2n-1)x^{n-1}+2^{n-1}.$
+Set $\psi(x)=2nx^{n-1}$ and
+(19) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 2\\ (2n-1)x^{n-1}+2^{n-1}&\text{if }\;x\geq 2,\end{array}\right$.
+so that $\phi(x)<\psi(x)$ with strict inequality when $x>2$.
+Moreover it is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and the strict inequality holds when $V>2n$.
+So we can apply Lemma 2.4 to conclude that $F^{\phi}(V)\geq F^{\psi}(V)$ and the strict inequality holds for $V>2n$.
+From Lemma 2.4 and Corollary 2.5, we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$ (see Example 2.7).
+This completes the proof of Proposition 4.2.
+4.3. Minimal rational curves with degree $d\geq 3$
+In this section, we are going to prove the following result which resolves a large part of Theorem 1.1.
+Recall the invariant $l_{X}$ from (7), whose definition is the smallest degree of minimal rational curves on $X$.
+Proposition 4.3.
+Let $X$ be a K-semistable Fano manifold with $3\leq l_{X}\leq n-1$. Then ${\rm vol}(X)<2n^{n}$.
+We use the notation from subsection 4.1 and specialize to the case when $\ell=2$ and $3\leq d\leq n+1$.
+For any $s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, choose a local generator $\sigma_{t}$ of $L^{k}$ over $U_{t}$, the section $[s]$ is represented uniquely over $U_{t}$ by a holomorphic function with the following expression:
+$\frac{s_{t}}{\sigma_{t}}$ $=$ $\sum_{|I|+2|J|=y}b^{t}_{IJ}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J}=\sum_{m=0}^{\lfloor y/2\rfloor}\sum_{\begin{subarray}{c}|I|+|J|=y-\lfloor y/2\rfloor+m\\ |I|=y-2\lfloor y/2\rfloor+2m\\ |J|=\lfloor y/2\rfloor-m\end{subarray}}b_{IJ}^{t}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J}$
+which we call the initial term representation of $s$ over $U_{t}$. We set
+$\gamma_{t}([s]):=\sigma_{t}\cdot\sum_{\begin{subarray}{c}|I|=y-2\lfloor y/2\rfloor\\ |J|=\lfloor y/2\rfloor\\ |I|+|J|=y-\lfloor y/2\rfloor\end{subarray}}b^{t}_{IJ}(z_{n})({\rm d}{\mathbf{z}}^{\prime})^{\otimes I}({\rm d}{\mathbf{z}}^{\prime\prime})^{\otimes J}$
+which defines a local section of $L^{k}\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor}({\mathcal{O}}^{\oplus(n-d+1)}))$ over $B_{t}$.
+We claim that this section is well-defined over the whole ${\mathbb{P}}^{1}$, i.e.
+it is independent of $t\in{\mathbb{P}}^{1}$.
+For any other $t^{\prime}\in{\mathbb{P}}^{1}$, choose a sequence $\{t_{i};i=0,\dots,N\}$ and $B_{t_{i}}(\epsilon^{\prime})$ as before.
+We use the transition function (15) to get
+$\frac{s_{t}}{\sigma_{t}}$ $=$ $\sum_{|I|+2|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n})+R_{1})(\mathbf{a}^{\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime}+R_{2})^{I}(\mathbf{a}^{\prime\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime\prime}+R_{2})^{J}$
+$=$ $\sum_{|I|+2|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime})^{I}(\mathbf{a}^{\prime\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime\prime})^{J}+\sum_{|I|+2|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(z_{n})\hat{{\mathbf{z}}})^{\prime I}R^{\prime J}_{2}+R^{w}_{y+1}$
+$=:$ $\frac{s_{t^{\prime}}}{\sigma_{t^{\prime}}}\frac{\sigma_{t^{\prime}}}{\sigma_{t}}+\tilde{R}^{w}_{y+1}$
+where $R^{\prime}_{2}$ is quadratic in $\hat{z}_{1},\dots,\hat{z}_{d-1}$, $R^{w}_{y+1}$ consists of terms of weighted degree $2|I|+J\geq y+1$ and $\tilde{R}^{w}_{y+1}$ consists of terms with $2|I|+J\geq y$ and $|I|+|J|>y-2\lfloor y/2\rfloor$.
+As a consequence, we get:
+$\gamma_{t^{\prime}}([s])$ $=$ $\sigma_{t}\cdot\sum_{\begin{subarray}{c}|I|=y-2\lfloor y/2\rfloor\\ |J|=\lfloor y/2\rfloor\end{subarray}}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(a^{\prime}(z_{n}){\rm d}\hat{{\mathbf{z}}}^{\prime})^{\otimes I}(a^{\prime\prime}(z_{n})\hat{{\mathbf{z}}}^{\prime\prime})^{\otimes J}$
+which coincides with $\gamma_{t}([s])$ on $B_{t}\cap B_{t^{\prime}}$.
+So we get a linear map of vector spaces:
+$\gamma:H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\rightarrow H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor}({\mathcal{O}}^{\oplus(n-d+1)})),$
+whose kernel consists of elements $s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ that has the initial term representation given by ($m\geq 1$ now):
+$\frac{s_{t}}{\sigma_{t}}=\sum_{m=1}^{\lfloor y/2\rfloor}\quad\sum_{\begin{subarray}{c}|I|+|J|=y-\lfloor y/2\rfloor+m\\ |I|=y-2\lfloor y/2\rfloor+2m\\ |J|=\lfloor y/2\rfloor-m\end{subarray}}b_{IJ}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J},$
+which by the same argument as above induces a section in $H^{0}({\mathbb{P}}^{1},L^{k}\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor+2}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor-1}({\mathcal{O}}^{\oplus(n-d+1)}))$.
 So we inductively get the estimate:
-$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$$\leq$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{y-\ell\lfloor y/\ell\rfloor+m\ell}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)}))$
-$\leq$ $\sum_{m=0}^{\lfloor y/\ell\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{\ell m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/\ell\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
-For the last inequality, we used the facts that $y-\ell\lfloor y/\ell\rfloor\geq 0$ and ${\mathcal{O}}(-1)$ is negative. Then the right-hand-side of (18) is bounded above by
+$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$
+$\leq$ $\sum_{m=0}^{\lfloor y/2\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{y-2\lfloor y/2\rfloor+2m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)}))$
+(21) $\leq$ $\sum_{m=0}^{\lfloor y/2\rfloor}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{2m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{\lfloor y/2\rfloor-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
+For the last inequality, we used the facts that $y-2\lfloor y/2\rfloor\geq 0$ and ${\mathcal{O}}(-1)$ is negative.
+Remark 4.4.
+In the first version of this preprint, we defined valuations $v_{\ell}$ for any $\ell\geq 1\in{\mathbb{N}}$ and claimed that the above estimate is true for any $\ell\geq 1\in{\mathbb{N}}$.
+Ziquan Zhuang pointed out that for $\ell\geq 3$, this is not true.
+The reason is that while $v_{1}$ and $v_{2}$ are well-defined globally, the valuation $v_{\ell}$ for $\ell\geq 3$ depends on a collection of coordinate charts that only cover a dense open subset $f({\mathbb{P}}^{1})$ (the complement of finite many points), which causes the failure of the above estimate.
+For $v_{1}$ and $v_{2}$ the coordinates charts we constructed cover the whole $f({\mathbb{P}}^{1})$ and there is no such an issue.
+By a similar argument in Subsection 4.1, a sufficient condition for $v_{\ell}$ to be defined globally is that the transition functions between coordinate charts covering the whole $f({\mathbb{P}}^{1})$ are of the form:
+$z_{k}=a_{k}(\hat{z}_{n})\hat{z}_{k}+R_{\ell},1\leq k\leq n-1;z_{n}=g_{n}(\hat{z}_{n})+R_{\ell-1}$
+where $R_{\ell}$ (resp.
+$R_{\ell-1}$) consists of terms of degree at least $\ell$ (resp.
+at least $\ell-1$) in $\hat{z}_{1},\dots,\hat{z}_{n-1}$.
+When the minimal rational curve is embedded, this condition is satisfied exactly when the rational curve is $(\ell-2)$-comfortably embedded in the sense of Abate-Bracci-Tovena (see [ABT09, Remark 3.3]).
+The latter condition is a refinement and a slight weakening of Grauert’s $(\ell-1)$-linearizability condition.
+From this point of view, the valuation $v_{1}$, $v_{2}$ are always globally defined since any minimal rational curve is always $1$-linearizable, and in particular (trivially) 0-comfortably embedded.
+By the similar (and simpler) argument as above, we get the similar estimate for the valuation $v_{1}$ (i.e $v_{\ell}$ from section 4.1 with $\ell=1$):
+$h^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}(v_{1})/{\mathcal{I}}_{y+1}(v_{1}))$ $\leq$ $\sum_{m=0}^{y}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd)\otimes{\rm Sym}^{m}({\mathcal{O}}(-1)^{\oplus(d-2)})\otimes{\rm Sym}^{y-m}({\mathcal{O}}^{\oplus(n-d+1)})).$
+By (16) and (21), for each $v_{\ell}$ with $\ell\in\{1,2\}$, we get that $h^{0}(X,L^{k}\otimes{\mathcal{O}}_{X}/{\mathcal{I}}_{xk}(v_{\ell}))$ is bounded from above by:  [^1] Our main interest will be the case $\ell=2$.
+We include the parameter $\ell=1$ for comparison later.
 $I_{k}(x):=\sum_{y=0}^{xk-1}\sum_{m=0}^{\lfloor y/\ell\rfloor}a_{y,m}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd-\ell m)),$
-where
+where the coefficients
 $a_{y,m}=\binom{d-3+\ell m}{d-3}\binom{n-d+\lfloor y/\ell\rfloor-m}{n-d}$.
-We will calculate the leading coefficient $b_{1}={\rm lim}_{k\rightarrow+\infty}I_{k}(x)/k^{n}$. First, we assume that $x\leq d$. Then we can expand:
+We will calculate the leading coefficient $b_{1}={\rm lim}_{k\rightarrow+\infty}(I_{k}(x)\cdot n!)/k^{n}$. First, we assume that $x\leq d$. Then we can expand:
 $I_{k}(x)$ $=$ $\sum_{y=0}^{xk-1}\sum_{m=0}^{\lfloor y/\ell\rfloor}\frac{(\ell m)^{d-3}+O((\ell m)^{d-4})}{(d-3)!}\frac{(\lfloor y/\ell\rfloor-m)^{n-d}+O((\lfloor y/\ell\rfloor-m)^{n-d-1})}{(n-d)!}(kd-\ell m+1)$
@@
 (22) $\frac{n!}{(d-3)!(n-d)!}\int_{0}^{x}{\rm d}t\int_{0}^{t/\ell}(\ell s)^{d-3}(\ell^{-1}t-s)^{n-d}(d-\ell s){\rm d}s$
-$=$ $\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}\int^{x}_{0}{\rm d}t\int_{0}^{t}\left(z^{d-3}(d-z)(t-z)^{n-d}\right)\,{\rm d}z$
+$=$ $\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}\int^{x}_{0}{\rm d}t\int_{0}^{t}z^{d-3}(d-z)(t-z)^{n-d}\,{\rm d}z$
 $=$ $\ell^{-(n-d+1)}\frac{n!}{(d-3)!(n-d+1)!}\int_{0}^{x}z^{d-3}(d-z)(x-z)^{n-d+1}\,{\rm d}z$
@@
 $\sum_{y=0}^{dk}\sum_{i=0}^{\lfloor y/\ell\rfloor}a_{y,i}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd-\ell i))+\sum_{y=dk+1}^{xk-1}\sum_{i=0}^{\lfloor dk/\ell\rfloor}a_{y,i}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(kd-i\ell)).$
-Similar calculation as above shows that this sum equals $b_{2}k^{n}/n!+O(k^{n-1})$ with $b_{2}$ equal to $C\cdot\mathbf{I}$ where $C=\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}$ and $\mathbf{I}$ is equal to:
+Similar calculation as above shows that this sum equals $b^{\prime}_{1}k^{n}/n!+O(k^{n-1})$ with $b^{\prime}_{1}$ equal to $C\cdot\mathbf{I}$ where $C=\frac{n!}{(d-3)!(n-d)!}\ell^{-(n-d+1)}$ and $\mathbf{I}$ is equal to:
 $\int_{0}^{d}{\rm d}t\int_{0}^{t}z^{d-3}(t-z)^{n-d}(d-z){\rm d}z+\int_{d}^{x}{\rm d}t\int_{0}^{d}z^{d-3}(t-z)^{n-d}(d-z){\rm d}z$
@@
 Set
-[…]
+(26) $\Phi(x)=\int_{0}^{x}\phi(t)dt=\left\{\begin{array}[]{ll}\frac{\ell^{-(n-d+1)}x^{n}}{n+1}(d(n+1)-(d-2)x)&0\leq x\leq d\\ \frac{\ell^{-(n-d+1)}n!}{(d-3)!(n-d+2)!}\int_{0}^{\rm d}z^{d-3}(d-z)(x-z)^{n-d+2}{\rm d}z&x\geq d.\end{array}\right$.
+By using the binomial expansion:
+$(x-z)^{n-d+1}=(x-d+d-z)^{n-d+1}=\sum_{j=0}^{n-d+1}\binom{n-d+1}{j}(d-z)^{j}(x-d)^{n-1+d-j}.$
+it is straightforward to get the following combinatorial expressions when $x\geq d$:
+(27) $\phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=0}^{n-d+1}\binom{n}{j}\cdot(n-d+2-j)d^{n-j}(x-d)^{j},$
+(28) $\Phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=0}^{n-d+2}\binom{n+1}{j}\frac{n-d+3-j}{n+1}d^{n+1-j}(x-d)^{j}$
 Set $T={\rm sup}\{x:V-\phi(x)=0\}$.
 If $T\leq d$, then we get $V=\phi(T)\leq\phi(d)=\ell^{-(n-d+1)}d^{n}(n-d+2)\leq d^{n}(n-d+2)$.
-Note that the number $d^{n}(n-d+2)$ is nothing but the volume of the Fano hypersurface in ${\mathbb{P}}^{n+1}$ of degree $n-d+2$ which is always less than $2n^{n}$ if $22<d<n+$ 2.
-So we can assume $T\geq d$. The equation $\Phi(T)=(T-A)\phi(T)$ from (4) with $A=A_{X}(E)=(d-2)+(n-d+1)\ell$ becomes:
-$\frac{1}{n-d+2}\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+2}\,{\rm d}z=(T-(d-2)-(n-d+1)\ell)\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+1}\,{\rm d}z.$
-Even though the explicit solution of $T$ is not available, as $\ell\rightarrow+\infty$, the above equation can be written as:
-$\frac{1}{n-d+2}T^{n-d+2}(1+O(T^{-1}))=(T-(n-d+1)\ell+O(1))T^{n-d+1}(1+O(T^{-1})).$
-So we get the asymptotic expression for its solution: $T=T(\ell)=(n-d+2)\ell+O(1)$. So we can calculate that, up to $O(\ell^{-1})$, $\phi(T)$ is given by:
-[…]
-$=$ $\ell^{-(n-d+1)}\frac{n!}{(d-3)!(n-d+1)!}\left(\int_{0}^{1}t^{d-3}(1-t){\rm d}t\right)\cdot d^{d-1}\ell^{n-d+1}(n-d+2)^{n-d+1}$
-$=$ $(n-d+2)^{n-d+1}\frac{n!}{(d-1)!(n-d+1)!}d^{d-1}={\rm vol}({\mathbb{P}}^{n-d+1}\times{\mathbb{P}}^{d-1}).$
-Letting $\ell\rightarrow+\infty$, by Corollary 2.5, we have thus proved ${\rm vol}(X)=(-K_{X})^{n}\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ and complete the proof of Theorem 1.4. ∎
+Note that the number $d^{n}(n-d+2)$ is nothing but the volume of the Fano hypersurface in ${\mathbb{P}}^{n+1}$ of degree nd+2n-d+2 which is always less than $2n^{n}$ if $2<d<n+2$.
+Remark 4.5.
+Setting $d=2$ in the formula (27)-(28), we also recover functions from Example 2.7 (the $\ell^{1-n}$ factor comes the fact that here we are using the $\ell^{\oplus(n-1)}$ blowup):
+$\phi(x)=2n\cdot\ell^{1-n}x^{n-1},\quad\Phi(x)=2\ell^{1-n}x^{n}.$
+So we can assume $T\geq d$. Now we set $\ell=2$ and $A=A_{X}(v_{2})=(d-2)+2(n-d+1)=2n-d$. The equation $\Phi(T)=(T-A)\phi(T)$ from (4) becomes:
+(29) $\frac{1}{n-d+2}\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+2}\,{\rm d}z=(T-(2n-d))\int_{0}^{d}z^{d-3}(d-z)(T-z)^{n-d+1}\,{\rm d}z.$
+By Corollary 2.5, there exists a unique solution $T=T(d,n)$ to the above equation.
 Example 4.6.
-By using binomial expansion, one can get the following combinatorial expressions for $x\geq d$:
-$\phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=d-1}^{n}\binom{n}{j}\cdot(j-d+2)\cdot d^{j}\cdot(x-d)^{n-j},$
-$\Phi(x)$ $=$ $\ell^{-(n-d+1)}\sum_{j=d-1}^{n+1}\binom{n+1}{j}\frac{j-d+2}{n+1}d^{j}\cdot(x-d)^{n+1-j}.$
-When $d=n+1$, the above expression becomes: $\phi(x)=(n+1)^{n}$ (for $x\geq d$) which implies the volume estimate ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$.
-When $d=n$, the above expressions become:
-$\phi(x)$ $=$ $\ell^{-1}\left(n\cdot n^{n-1}\cdot(x-n)+2n^{n}\right)=\ell^{-1}n^{n}(x+2-n),$
-$\Phi(x)$ $=$ $\frac{\ell^{-1}}{n+1}\left(\binom{n+1}{2}n^{n-1}(x-n)^{2}+(n+1)2n^{n}(x-n)+3n^{n+1}\right)$
-$=$ $\frac{\ell^{-1}n^{n}}{2}\left(x^{2}-2(n-2)x+\frac{n}{(n+1)}(n-1)(n-2)\right).$
-The equation $(T-(n-2)-\ell)\phi(T)=\Phi(T)$ is asymptotically given by:
-$(T-\ell+O(1))T(1+O(T^{-1}))=\frac{1}{2}T^{2}(1+O(T^{-1}))$
-which implies $T=2\ell+O(1)$.
-So $\phi(T)=\ell^{-1}n^{n}(2\ell+O(1)+2-n)\longrightarrow 2n^{n}$ as $\ell\rightarrow+\infty$ and we get an estimate: ${\rm vol}(X)\leq 2n^{n}={\rm vol}({\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1})={\rm vol}(Q^{n})$.
+When $d=n+1$, the expression (23) becomes: $\phi(x)=(n+1)nx^{n-1}-(n-1)x^{n}$ if $x\leq n+1$ and $\phi(x)=(n+1)^{n}$ (for $x\geq n+1$).
+This implies the (un-interesting) volume estimate ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$.
+Example 4.7.
+When $d=n$, the expressions (27)-(28) become:
+(30) $\phi(x)$ $=$ $\ell^{-1}\left(n\cdot n^{n-1}\cdot(x-n)+2n^{n}\right)=\ell^{-1}n^{n}(x+2-n),$
+$\Phi(x)$ $=$ $\frac{\ell^{-1}n^{n}}{2}\left(x^{2}-2(n-2)x+\frac{n}{(n+1)}(n-1)(n-2)\right).$
+When $d=n$ and for the case of $(1^{\oplus(d-2)},\ell^{\oplus(n-d+1)})$-weighted blowup, equation $(T-A(v_{\ell}))\phi(T)=\Phi(T)$ is equivalent to $T^{2}-2(n-2+\ell)T+\left(\frac{n(n-2)(n+3)}{2(n+1)}+(2\ell-4)\right)=0$.
+One can solve
+$T=\left\{\begin{array}[]{ll}(n-1)+\sqrt{\frac{3(n-1)}{n+1}}&\text{for }\ell=1;\\ n+\sqrt{\frac{6n}{n+1}}&\text{for }\ell=2.\end{array}\right$.
+So we get the volume upper bound ${\rm vol}(X)\leq\phi_{\ell}(T)$ which is given by:
+$\phi_{\ell}(T)=\left\{\begin{array}[]{ll}n^{n}(T-n+2)=n^{n}(1+\sqrt{\frac{3(n-1)}{n+1}})&\text{for }\ell=1\\ \frac{1}{2}n^{n}(T-n+2)=n^{n}\left(1+\sqrt{\frac{3n}{2(n+1)}}\right)&\text{for }\ell=2.\end{array}\right$.
+Note that when $n=2$, $\phi_{1}(T)=\phi_{2}(T)={\rm vol}(Q)=8$.
+When $n\geq 3$, $\phi_{1}(T)>\phi_{2}(T)>2n^{n}={\rm vol}(Q)$.
+To get the sharp upper bound of ${\rm vol}(X)=(-K_{X})^{n}$ for all $d=n\geq 3$, we are going to use the classification results stated in Theorem 2.9.
+By Corollary 2.5, Proposition 4.3 is equivalent to the following proposition.
+Proposition 4.8.
+Let $3\leq d\leq n-1$ and $T$ be the solution to equation (29).
+We have $\phi(T)<2n^{n}$ where $\phi$ was defined in (23).
+As a consequence, we know that if $X$ is a K-semistable Fano manifold $3\leq l_{X}=d\leq n-2$, then ${\rm vol}(X)<2n^{n}$.
+Remark 4.9.
+Example 4.7 shows that for $d=n$ the estimate $\phi(T)<2n^{n}$ is not true.
+Surprisingly, Proposition 4.8 implies that the $(1^{\oplus(d-2)},2^{\oplus(n-d+1)})$ weighted blow-up is sufficient to establish the sharp estimate for all $3\leq d\leq n-1$.
+Even though the statement appears elementary, due to the complicated expansion from (27)-(28) it seems not so easy to find a simple proof for all pairs of positive integers $(d,n)$ satisfying $3\leq d\leq n-1$.
+Motivated by some numerical calculations, we managed to get the desired estimate using some tools from calculus and probability theory.
+See Figure 1 for the numerical illustration for $n=7$ and $2\leq d\leq 7$.
+The rest of this subsection is devoted to explain the proof of Proposition 4.8.
+We first consider the case d=n1d=n-1.
+Note that equation $(T-A_{X}(E))\phi(T)=\Phi(T)$ appeared in Corollary 2.5 is a degree nd+2n-d+2 polynomial equation.
+When d=n1d=n-1, the corresponding equation is a cubic equation in $T$ for which we can explicitly compute the root $T$ using Cardano’s formula.
+Proposition 4.10.
+Proposition 4.8 holds when d=n1d=n-1.
+Proof.
+When d=n1d=n-1 and $\ell=2$, the expressions from (27)-(28) become:
+$\phi(x)$ $=$ $\frac{1}{2^{2}}\left(\frac{n(n-1)^{n-1}}{2}(x-n+1)^{2}+2n(n-1)^{n-1}(x-n+1)+3(n-1)^{n}\right),$
+$\Phi(x)$ $=$ $\frac{1}{2^{2}}(\frac{n(n-1)^{n-1}}{6}(x-n+1)^{3}+n(n-1)^{n-1}(x-n+1)^{2}$
+$\,+3(n-1)^{n}(x-n+1)+\frac{4}{n+1}(n-1)^{n+1}).$
+The equation $(T-A(E))\phi(T)=\Phi(T)$ is equivalent to $\tau^{3}-12\tau-\frac{6(n-1)(5n+1)}{n(n+1)}=0$ where $\tau=T-n+1$. By Cardano’s formula,
+$\tau=\sqrt[3]{-\frac{q}{2}+\sqrt{\frac{q^{2}}{4}-64}}+\sqrt[3]{-\frac{q}{2}-\sqrt{\frac{q^{2}}{4}-64}},$
+where $q=-\frac{6(n-1)(5n+1)}{n(n+1)}$.
+Hence, ${\rm vol}(X)\leq\phi(T)=\frac{n(n-1)^{n-1}}{8}((\tau+2)^{2}+2-\frac{6}{n})\eqqcolon F(n)$.
+We claim that for $n\geq 4$, we have $F(n)<2n^{n}={\rm vol}(Q)$.
+When $n\geq 4$, note that the sequence $\tau(n)$ is monotone increasing with respect to $n$ and ${\rm lim}_{n\rightarrow+\infty}\tau(n)=\sqrt[3]{15+\sqrt{161}}+\sqrt[3]{15-\sqrt{161}}\eqqcolon\tau_{\infty}$.
+So, in particular, we have $\tau\leq\tau_{\infty}$.
+We further consider
+$\frac{F(n)}{2n^{n}}$ $\leq$ $\frac{n(n-1)^{n-1}}{8\cdot 2n^{n}}\left((\tau_{\infty}+2)^{2}+2-\frac{6}{n}\right)$
+$\leq$ $\frac{1}{16\left(1+\frac{1}{n-1}\right)^{n-1}}\left(\left(\sqrt[3]{15+\sqrt{161}}+\sqrt[3]{15-\sqrt{161}}+2\right)^{2}+2\right)\eqqcolon\gamma(n).$
+Note that $\gamma(n)$ is monotone decreasing with respect to $n$ and ${\rm lim}_{n\rightarrow\infty}\gamma(n)=\frac{1}{16e}(\sqrt[3]{15+\sqrt{161}}+\sqrt[3]{15-\sqrt{161}}+2)^{2}<1$.
+So $\gamma(n)<1$ when $n$ is sufficiently large.
+By numerical method, we can show that $\gamma(n)<1$ when $n\geq 19$.
+For the case when $4\leq n\leq 18$, by direct computation, one can check that $F(n)<2n^{n}$.
+∎
+Remark 4.11.
+If we use the valuation $v_{1}$ in the case of d=n1d=n-1, we can do a similar computation as in the proof of Proposition 4.10, then
+$\tau=T-n+1=-1+\sqrt[3]{-1+\frac{q}{2}+\sqrt{\frac{q^{2}}{4}-q}}+\sqrt[3]{-1+\frac{q}{2}-\sqrt{\frac{q^{2}}{4}-q}}$
+where $q=\frac{12(n-1)^{2}}{n(n+1)}$.
+In this case, $\phi(T)=\frac{n(n-1)^{n-1}}{2}\tau^{2}+2n(n-1)^{n-1}\tau+3(n-1)^{n}>2n^{n}={\rm vol}(Q)$ when $n\geq 4$.
+This shows that the use of $(1^{\oplus(d-2)},2^{\oplus(n-d+1)})$-weighted blowup is crucial for the proof of Proposition 4.3 and hence for Theorem 1.1.
+Completion of proof of Proposition 4.8.
+Because of Proposition 4.10, we just need to prove Proposition 4.8 when $3\leq d\leq n-2$. For simplicity, set
+$\Psi(x)$ $=$ $(x-(2n-d))\phi(x)-\Phi(x).$
+Note that $\Psi^{\prime}(x)=(x-(2n-d))\phi^{\prime}(x)$ since $\phi(x)=\Phi^{\prime}(x)$. Since $\phi^{\prime}(x)>0$ for $x>0$, we have the two useful facts:
+  • $\Psi(x)$ is negative for $x\in(0,T)$ and positive for $x\in(T,+\infty)$, where $T$ is the unique solution to $\Psi(T)=0$ (see Corollary 2.5).
+  • x>2nd+1x>2n-d+1, $\Psi^{\prime}(x)>\phi^{\prime}(x)$.
+Now suppose there exists $x_{1}\geq 2n-d+1$ such that
+$2n^{n}-\phi(x_{1})+\Psi(x_{1})>0,$
+or equivalently $\phi(x_{1})<2n^{n}-\Psi(x_{1})$. If $\Psi(x_{1})\geq 0$, then $T\leq x_{1}$ and $\phi(T)\leq\phi(x_{1})<2n^{n}$. If $\Psi(x_{1})<0$, then $T>x_{1}$ and since $\Psi(T)=0$,
+$\phi(T)$ $=$ $\phi(x_{1})+\int_{x_{1}}^{T}\phi^{\prime}(x){\rm d}x<\phi(x_{1})+\int_{x_{1}}^{T}\Psi^{\prime}(x){\rm d}x$
+$=$ $\phi(x_{1})+\Psi(T)-\Psi(x_{1})<2n^{n}.$
+By this simple calculus argument, we see that the proposition 4.8 is proved once we prove (31) for some $x_{1}\geq 2n-d+1$.
+Now the main claim is that (31) indeed holds when $3\leq d\leq n-2$ and $x_{1}=2n-d+2$.
+In other words, we need to show that for any $3\leq d\leq n-2$, the following inequality is true:
+(31) $\phi(2n-d+2)-\Psi(2n-d+2)=\Phi(2n-d+2)-\phi(2n-d+2)<2n^{n}.$
+Since the proof depends on long estimates, we put the proof in the Appendix A. ∎
+4.4. Completion of the proof of Theorem 1.1
 Proof of Theorem 1.1.
 We use the notation $l_{X}$ from (6)-(7).
-If $l_{X}=n+1$, then as mentioned in Example 2.9 $X$ must be ${\mathbb{P}}^{n}$, and if $l_{X}=n$ then $X$ is either $Q^{n}$ or the blowup of ${\mathbb{P}}^{n}$ along a smooth subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ in a hyperplane.
+If $l_{X}=n+1$, then by Theorem 2.9, $X$ must be ${\mathbb{P}}^{n}$, and if $l_{X}=n$ then $X$ is either $Q^{n}$ or the blowup of ${\mathbb{P}}^{n}$ along a smooth codimension two subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ in a hyperplane.
 When $d_{Y}\geq 2$, by Lemma 4.12 below we know $(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}<2n^{n}$.
 When $d_{Y}=1$, by Lemma 4.13 below we know that ${\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is not a K-semistable Fano manifold, so it is excluded.
-Thus we verified the Theorem in the case $l_{X}\geq n$.
+Thus, we verified the Theorem 1.1 in the case $l_{X}\geq n$.
 When $l_{X}<n$, there are minimal rational curves of degree $d\in\{2,\dots,n-1\}$.
-For $12<d<n-$ 1, by Theorem 1.4 and Lemma 3.5, ${\rm vol}(X)\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})<2n^{n}$.
-When $d=2$, by Theorem 3.1 we get ${\rm vol}(X)\leq 2n^{n}$ with equality if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
+For $3\leq d\leq n-1$, by Proposition 4.8, ${\rm vol}(X)<2n^{n}$.
+When $d=2$, by Theorem 1.4 and Proposition 4.2 we get ${\rm vol}(X)\leq 2n^{n}$ with equality if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
 ∎
 Lemma 4.12.
-Suppose $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\cdots,n\}$ that is contained in a hyperplane (See Example 2.9). Then
+Suppose $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\cdots,n\}$ that is contained in a hyperplane (See Theorem 2.9). Then
 $(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}=\left\{\begin{array}[]{ll}2n^{n}&\text{if }d_{Y}=1\\ \frac{d_{Y}n^{n}-(n+1-d_{Y})^{n}}{d_{Y}-1}=:V_{d_{Y}}&\text{if }2\leq d_{Y}\leq n.\end{array}\right$.
@@
 $=$ $(n+1)^{n}-\sum_{k=2}^{n}\binom{n}{k}(n+1)^{n-k}\cdot(-1)^{k-2}\left(\sum_{i=0}^{k-2}d^{i}\right)\cdot d,$
-where $s_{k-2}(N_{Y/{\mathbb{P}}^{n}})$ denote the $(k-2)$-th Segre class of normal bundle $N_{Y/{\mathbb{P}}^{n}}$.
+where $s_{k-2}(N_{Y/{\mathbb{P}}^{n}})$ denote the (k2)(k-2)-th Segre class of normal bundle $N_{Y/{\mathbb{P}}^{n}}$.
 If $d=1$, then $\sum_{i=0}^{k-2}d^{i}=k-1$.
@@
 Lemma 4.13.
-Suppose $\pi\colon X\rightarrow{\mathbb{P}}^{n}$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y\cong{\mathbb{P}}^{n-2}$ of degree $d_{Y}=1$ that is contained in a hyperplane, then $X$ is K-unstable Fano manifold.
+Suppose $\pi\colon X\rightarrow{\mathbb{P}}^{n}$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y\cong{\mathbb{P}}^{n-2}$ of degree $d_{Y}=1$ contained in a hyperplane of ${\mathbb{P}}^{n}$, then $X$ is K-unstable.
 Proof.
@@
 Note that the nef cone is $\text{Nef}(X)={\mathbb{R}}_{\geqslant 0}[H]+{\mathbb{R}}_{\geqslant 0}[H-E]$.
-Then the divisor $-K_{X}-tE=(n+1)H-(1+t)E$ is nef if and only if $0\leqslant t\leqslant n$.
+Then the divisor $K_{X}-tE=(n+1)H-(1+t)E$ is nef if and only if $0\leqslant t\leqslant n$.
 When $0\leqslant t\leqslant n$, by the similar computation as in Lemma 4.12,
@@
 Here the ray generated by $f$ corresponds to the exceptional divisor $E=\pi^{-1}({\mathbb{P}}^{n-2})$ by the cone-orbit correspondence.
-Then the moment polytope with respect to $-K_{X}$ is given by
+Then the moment polytope with respect to $K_{X}$ is given by
 $P_{X}=\{(x_{1},\cdots,x_{n})\in{\mathbb{R}}^{n}\mid x_{1}+x_{2}\geq-1,\,\sum_{i=1}^{n}x_{i}\leq 1,\,x_{j}\geq-1\,\,\text{for}\,\,j=1,\cdots,n\}.$
@@
 $C:$ $v_{1}=v_{4}=\cdots=v_{n+1}=0.$
+$v_{0}=\frac{Z_{2}}{Z_{2}}=u_{2}^{-1},v_{1}=\frac{u_{1}}{u_{2}}=-u_{3}-\sum_{k=4}^{n+1}\frac{u_{k}^{2}}{u_{2}},v_{k}=\frac{Z_{k}}{Z_{2}}=\frac{Z_{k}}{Z_{0}}\frac{Z_{0}}{Z_{2}}=u_{k}\cdot u_{2}^{-1}$
 The conormal bundle of $C\setminus\{x\}$ is generated by
@@
 In the $\mathbf{v}$ coordinates, we also have the weights $(0,2,1,\dots,1)$.
-The weighted blowup $\mu:\hat{X}\rightarrow X$ in this example is globally defined and the exceptional divisor $E$ is isomorphis to a weighted projective ${\mathbb{P}}(2,\underbrace{1,\dots,1}_{n-2})$-bundle over ${\mathbb{P}}^{1}$.
+We see that the weighted blowup $\mu:\hat{X}\rightarrow X$ is indeed globally defined and the exceptional divisor $E$ is isomorphis to a weighted projective ${\mathbb{P}}(2,\underbrace{1,\dots,1}_{n-2})$-bundle over ${\mathbb{P}}^{1}$.
 We can calculate the volume function ${\rm vol}(-K_{X}-xE)=n^{n}{\rm vol}(H-\xi E)$ where $\xi=\frac{x}{n}$.
@@
 The second piece is more difficult to get.
-One way to get it is to use the fact that the measure $-\frac{1}{n!}d{\rm vol}(H-tE)$ is the Duistermaat-Heckman measure of the previously mentioned Hamiltonian $S^{1}$ action (see [BHJ17]) and then use the symmetry of its density function around $\xi=1$ in our case.
-We can also directly calculate the Duistermaat-Heckman measure of $S^{1}$-action by using a well-established localization/Fourier formula ([BGV92, 7.4]) and for the interest of the reader we will provide such a derivation in appendix.
+One way to get it is to use the fact that the measure $\frac{1}{n!}d{\rm vol}(H-tE)$ is the Duistermaat-Heckman measure of the previously mentioned Hamiltonian $S^{1}$ action (see [BHJ17]) and then use the symmetry of its density function around $\xi=1$ in our case.
+We can also directly calculate the Duistermaat-Heckman measure of $S^{1}$-action by using a well-established localization/Fourier formula ([BGV92, 7.4]) and for the interest of the reader we will provide such a derivation in Appendix B.
+It is straightforward to verify the estimate ${\rm vol}(-K_{X}-xE)\geq V-\phi(x)$ (see (30)) for this example which is reduced to the inequality: for $n\leq x\leq 2n$,
+$F(x):=\frac{n^{2}}{2}(2n-x)^{n-1}-\frac{n-2}{2}(2n-x)^{n}+\frac{1}{2}n^{n}(x+2-n)-2n^{n}\geq 0.$
+This can be easily verified by the following facts: $F(n)=F^{\prime}(n)=0$ and $F^{\prime\prime}(x)=\frac{n}{2}(n-1)(n-2)(2n-x)^{n-3}(x-n)>0$ for $n<x<2n$.
 Example 4.17.
 Let $X\subset{\mathbb{P}}^{n+1}$ be a Fano hypersurface of degree $1\leq b\leq n$.
-Then the minimal rational curve is of splitting type ${\mathcal{O}}(1)^{n-b}\oplus{\mathcal{O}}^{b-1}$ so that $d=n-b+2$ (see [Kol96, Exercise V.4.4]).
+Then the minimal rational curve is of splitting type ${\mathcal{O}}(1)^{n-b}\oplus{\mathcal{O}}^{b-1}$ so that d=nb+2d=n-b+2 (see [Kol96, Exercise V.4.4]).
 The volume of $X$ is given by $V_{d}=d^{n}(n-d+2)=(n+2-b)^{n}\cdot b$.
 One can easily verify that $V_{d}$ is strictly smaller than ${\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ except when $b=1$ or $b=2$.
-Question 4.8.
-Is there a K-semistable Fano manifold $X$ with ${\rm vol}(X)={\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ that contains a minimal rational curve of degree $13\leq d\leq n-$ 1 but is not isomorphic to ${\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1}$?
-Our results imply that when $d=n$, there is just one such $X$, i.e. the quadric hypersurface $Q^{n}$, while when $d=2$, there is no such $X$. Moreover, such an $X$ must satisfy $l_{X}=d$ (see (7)).
+Question 4.18.
+Assume that a K-semistable Fano manifold $X$ contains a minimal rational curve of degree $3\leq d\leq n-1$ in $X$.
+Is it true that ${\rm vol}(X)=(-K_{X})^{n}\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$?
+In the first version, this was claimed as a result.
+However due to the issue discussed in Remark 4.4, the question is left un-answered.
+On the other hand, our current results imply that the answer is affirmative if we instead assume $d\in\{2,n,n+1\}$.
 We end this paper by finding the minimal possible anti-canonical volume of $n$-dimensional (K-semistable) Fano manifolds.
@@
 If $a_{0}=\cdots=a_{n-1}=1,\,a_{n}=2,\,a_{n+1}=n+1$ and $d=2n+2$.
-Namely, $X_{d}$ is a general degree $22n+$ 2 hypersurface in ${\mathbb{P}}(1^{n},2,n+1)$.
+Namely, $X_{d}$ is a general degree $2n+2$ hypersurface in ${\mathbb{P}}(1^{n},2,n+1)$.
 Then $(-K_{X_{2n+2}})^{n}=1$.
@@
 So $(-K_{Y})^{n}=\deg(\phi)\cdot(-K_{{\mathbb{P}}^{n}}-\frac{1}{2}D)^{n}=2$.
-By the result of [LZ22, Zhu21], we know K-stability of $Y$ is equivalent to the K-stability of the log Fano pair $({\mathbb{P}}^{n},\frac{1}{2}D)$.
+It was first shown in [AGP06, Theorem 3.2] that $Y$ admits Kähler-Einstein metric, hence K-stable (See also [Der16] and [AZ22, Corollary 4.9(3)] for a purely algebraic proof of K-stable).
+Here we illustrate a different proof based on the result of [LZ22, Zhu21], we know K-stability of $Y$ is equivalent to the K-stability of the log Fano pair $({\mathbb{P}}^{n},\frac{1}{2}D)$.
 Since we assume $D$ is a smooth hypersurface, then the log Fano pair $({\mathbb{P}}^{n},(\frac{n+1}{2n}-\varepsilon)D)$ is klt, so it is uniformly K-stable by [ADL23, Theorem 2.10] for $0<\varepsilon\ll 1$.
@@
 Thus, $Y$ is a smooth K-stable Fano manifold with volume $(-K_{Y})^{n}=2$ for all dimensions $n\geqslant 2$.
-Note that $Y$ contains a degree 2 free-immersed rational curve with $n-1$ nodes (see [Kol96, IV.2.12.3]).
+Note that $Y$ contains a degree 2 free-immersed rational curve with n1n-1 nodes (see [Kol96, IV.2.12.3]).
 We indeed have the following:
@@
 ([KSC04, Solution to Exercise 5.23, Page 217]) Let $X$ be a smooth projective variety of odd dimension $n$. Then the self-intersection number $(-K_{X})^{n}$ is even.
-Question 4.12.
+Question 4.22.
 Classify all Fano manifolds that obtain the minimal volume.
-Appendix: A proof of the formula (22)
+Appendix A Proof of the estimate (31)
+The following graphs illustrate the result in Proposition 4.8.
+(a) d=7
+(b) d=6
+(c) d=5
+(d) d=4
+(e) d=3
+(f) d=2
+Figure 1. $n=7,7\geq d\geq 2$
+The solid curve (resp.
+the dashed curve) is the graph of $f(y):=\phi(y)/(2n^{n})-1$ (resp.
+$g(y):=\Psi(y)/(2n^{n})$) where y=xdy=x-d.
+The first graph illustrates Example 4.7.
+The second figure illustrates Proposition 4.10.
+The 3rd to 5th figures illustrates that for $d\in\{5,4,3\}$ at $y_{1}=2(n-d+1)=x_{1}-d\in\{6,8,10\}$, $g^{\prime}(y_{1})>f^{\prime}(y_{1})$ and $g(y_{1})>f(y_{1})$ as claimed in the proof of Proposition 4.8.
+The last figure illustrates Remark 4.5.
+The graphs are generated by the following codes.
+1L=2;n=7;
+[…]
+[…]
+[…]
+[…]
+[…]
+Proof of the estimate (31).
+For the simplicity of notation, we set r=n+1dr=n+1-d so that $r+d=n+1$. We will prove that (31) holds for $x_{1}=2r+d=2n-d+2$.
+(Case I: Both $r$ and $d$ are big) By using the expression (23)-(26), we can calculate:
+(35) $S=S(r,d)=\Phi(x_{1})-\phi(x_{1})$
+$=$ $\frac{2^{-(n-d+1)}n!}{(d-3)!(n-d+2)!}\int_{0}^{d}z^{d-3}(d-z)(x_{1}-z)^{n-d+1}(n-z){\rm d}z$
+(36) $=$ $\frac{1}{2^{r}(r+1)}\sum_{j=0}^{r}(r+1-j)\cdot\left[n(n-j+1)-d(d-2)\right]\frac{n!}{j!(n+1-j)!}d^{n-j}(2r)^{j}$
+(37) $=$ $\frac{(d+2r)^{n+1}\cdot n}{2^{r}(r+1)(n+1)d}\sum_{j=0}^{r}(r+1-j)(r-j+\Delta)\binom{n+1}{j}(1-p)^{n+1-j}p^{j}$
+where
+$\Delta=\frac{d(r+1)}{n},\quad p=\frac{2r}{d+2r},\quad 1-p=\frac{d}{d+2r}.$
+For the identity (36) we use the binomial expansion of $(x_{1}-z)^{r}=(2r+(d-z))^{r}$ and integrate via the definition of Beta functions.
+To estimate the partial binomial sum, we use a method from probability theory called Crámer’s tilt.
+Set
+$q=\frac{r}{n+1}=\frac{r}{r+d}<p=\frac{2r}{2r+d}$
+such that the new expectation of the binomial distribution is $r=(n+1)q$. We can calculate:
+$\frac{(1-p)^{n+1-j}p^{j}}{(1-q)^{n+1-j}q^{j}}$ $=$ $\left(\frac{p(1-q)}{q(1-p)}\right)^{j-(n+1)q}\left(\frac{1-p}{1-q}\left(\frac{p(1-q)}{q(1-p)}\right)^{q}\right)^{n+1}$
+$=$ $\eta^{r-j}\left(\frac{(1-p)^{1-q}p^{q}}{(1-q)^{1-q}q^{q}}\right)^{n+1}$
+where
+$\eta=\frac{q(1-p)}{p(1-q)}=\frac{rd}{2rd}=\frac{1}{2},\quad\frac{(1-p)^{1-q}p^{q}}{(1-q)^{1-q}q^{q}}=2^{q}\frac{r+d}{2r+d}.$
+Substituting into the formula (37) and let j=rkj=r-k, we get
+(38) $S$ $=$ $\frac{(n+1)^{n}n}{(r+1)d}\sum_{k=0}^{r}(k+1)(k+\Delta)\frac{}{}\binom{n+1}{r-k}(1-q)^{n+1-r+k}q^{r-k}2^{-k}$
+(39) $=$ $\frac{(n+1)^{n}n}{(r+1)d}\sum_{k=0}^{r}(k+1)(k+\Delta)2^{-k}a_{k}$
+where $a_{k}=\binom{n+1}{r-k}(1-q)^{n+1-r+k}q^{r-k}$. Note that for $k\geq 1$,
+$\frac{a_{k}}{a_{k-1}}=\frac{r-k+1}{n-r+k+1}\cdot\frac{d}{r}\leq\frac{d}{d+k}\leq\frac{d}{d+1}<1.$
+So if we set $\theta_{1}=\frac{d}{2(d+3)}<\frac{1}{2}$ and $\beta\coloneqq\frac{n+1}{dr}\geq\frac{n}{d(r+1)}=\Delta^{-1}$, we get the estimate:
+$\frac{S}{2n^{n}}$ $\leq$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}\sum_{k=0}^{+\infty}(k+1)(k+\Delta)2^{-k}\cdot a_{k}$
+$=$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}a_{0}\left(\Delta+(1+\Delta)\frac{a_{1}}{a_{0}}+\frac{a_{2}}{a_{0}}\sum_{j=0}^{+\infty}(j+3)(j+2+\Delta)2^{-j-2}\frac{a_{2+j}}{a_{2}}\right)$
+$\leq$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}a_{0}\left(\Delta+(1+\Delta)\frac{d}{d+1}+\frac{d^{2}}{4(d+1)(d+2)}\sum_{j=0}^{+\infty}(j+3)(j+2+\Delta)2^{-j}\left(\frac{d}{d+3}\right)^{j}\right)$
+$=$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}a_{0}\left(\alpha_{1}\Delta+\tau_{1}\right)$
+where $\alpha_{1}=1+\frac{d}{d+1}+\frac{d^{2}}{4(d+1)(d+2)}\frac{3-2\theta}{(1-\theta)^{2}}$ and $\tau_{1}=\frac{d}{d+1}+\frac{d^{2}}{2(d+1)(d+2)}\frac{\theta^{2}-3\theta+3}{(1-\theta)^{3}}$.
+By using Robbins’ version of the Stirling approximation:
+$e^{\frac{1}{12n+1}}<\frac{n!}{n^{n}e^{-n}\sqrt{2\pi n}}<e^{\frac{1}{12n}},$
+one can verify that: with $r=(n+1)q$,
+$a_{0}$ $=$ $\binom{n+1}{r}(1-q)^{n+1-r}q^{r}\leq\frac{1}{\sqrt{2\pi(n+1)q(1-q)}}$
+$=$ $\sqrt{\frac{(n+1)}{(2\pi)rd}}=\sqrt{\frac{\beta}{2\pi}}.$
+Thus, when $d\leq r$, we have $\beta=d^{-1}+r^{-1}\leq 2d^{-1}$, then
+$\frac{S}{2n^{n}}\leq\frac{e}{2\sqrt{2\pi}}\sqrt{\beta}\left(\alpha_{1}+\tau_{1}\beta\right)\leq\frac{e}{2\sqrt{2\pi}}\sqrt{(2/d)}\left(\alpha_{1}+\tau_{1}\cdot(2/d)\right)\leq 1$
+provided if $d\geq 9$.
+When $r\leq d$, similarly
+$\frac{S}{2n^{n}}$ $\leq$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}a_{0}\left(\Delta+(1+\Delta)\frac{a_{1}}{a_{0}}+\frac{a_{2}}{a_{0}}\sum_{j=0}^{+\infty}(j+3)(j+2+\Delta)2^{-j-2}\frac{a_{2+j}}{a_{2}}\right)$
+$\leq$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}a_{0}\left(\Delta+(1+\Delta)+\frac{r-1}{r}\sum_{j=0}^{+\infty}(j+3)(j+2+\Delta)2^{-j-2}\left(\frac{r-2}{r}\right)^{j}\right)$
+$=$ $\frac{(n+1)^{n}}{2n^{n}}\Delta^{-1}a_{0}\left(\alpha_{2}\Delta+\tau_{2}\right)$
+where $\theta=\theta(r)=\frac{r-2}{2r}$ and $\alpha_{2}=2+\frac{r-1}{4r}\frac{3-2\theta}{(1-\theta)^{2}},\,\tau_{2}=1+\frac{r-1}{2r}\frac{\theta^{2}-3\theta+3}{(1-\theta)^{3}}$.
+Use $\Delta^{-1}\leq\frac{r}{r+1}\beta$ with $\beta=d^{-1}+r^{-1}\leq 2r^{-1}$.
+We see the inequality
+$\frac{S}{2n^{n}}\leq\frac{e}{2\sqrt{2\pi}}\left(\alpha_{2}+\frac{r}{r+1}\tau_{2}\cdot(2/r)\right)\sqrt{2/r}\leq 1$
+holds provided if $r\geq 11$.
+(Case II: $r$ small and $d$ large) Recall the relation $n+1=r+d$.
+To illustrate our way to estimate, consider the case r=nd+1=3r=n-d+1=3 and $d\geq 3$ so that $n\geq 5$.
+Then we can calculate and estimate:
+$\frac{\Psi(x_{1})-\phi(x_{1})}{2n^{n}}$ $=$ $\frac{(n-2)^{n-2}}{16n^{n}(1+n)}(115n^{3}-126n^{2}-66n+40)$
+$\leq$ $\frac{e^{-2}}{16}\frac{115n^{3}-126n^{2}-66n+40}{(n-2)^{2}(1+n)}$
+$=$ $\frac{115e^{-2}}{16}\left(1+\frac{1}{115}\frac{{219}n^{2}-{66}n-{420}}{(n-2)^{2}(n+1)}\right)$
+$=$ $\frac{115e^{-2}}{16}\left(1+\frac{1}{115}\frac{m}{m+3}(219m^{-1}+810m^{-2}+324m^{-3})\right)$
+where for the first inequality we used $(n-2)^{n}/n^{n}\leq e^{-2}$ and for the last equality we set m=n2m=n-2.
+One can verify numerically that the last expression is less than 1 when $m=n-2\geq 69$.
+So we can just verify numerically for $6\leq n\leq 70$ or equivalently $3\leq d\leq 68$.
+For $r\geq 4$, we use the formula (38) for $S=S(r,d)$ to estimate (note that $\Delta=\frac{d(r+1)}{n}\leq r+1$):
+$\frac{S}{2n^{n}}=\frac{(n+1)^{n}n}{2n^{n}(r+1)d}\sum_{k=0}^{r}(k+1)(k+\Delta)\frac{}{}\binom{n+1}{r-k}(1-q)^{n+1-r+k}q^{r-k}2^{-k}$
+$\leq$ $\frac{(n+1)^{n}}{2n^{n}}\frac{1}{r+1}\left(1+\frac{r-1}{d}\right)\sum_{k=0}^{r}\frac{(k+1)(k+r+1)(n+1)!}{(r-k)!(n+1-(r-k))!}(\frac{d}{n+1})^{d+k}(\frac{r}{n+1})^{r-k}2^{-k}$
+$\leq$ $\frac{e}{2}\frac{1}{r+1}(1+\epsilon)\sum_{k=0}^{r}(k+1)(k+r+1)\frac{1}{(r-k)!}r^{r-k}2^{-k}(1+\frac{r}{d})^{-d}$
+where we set $\epsilon=r/d$. Moreover, we have $1+\epsilon<e^{\epsilon}$ and
+$-d\cdot{\rm log}(1+\epsilon)=-d(\epsilon-\frac{\epsilon^{2}}{2}+\cdots)\leq-d\epsilon+d\frac{\epsilon^{2}}{2}=-r+\frac{r\epsilon}{2}.$
+Then
+$\frac{S}{2n^{n}}$ $\leq$ $\frac{e}{2(r+1)}(1+\epsilon)\sum_{k=0}^{r}\frac{(k+1)(k+r+1)r^{r-k}}{2^{k}(r-k)!}e^{-r+r\epsilon/2}.$
+We define
+(40) $R_{\infty}=R_{\infty}(r)=\frac{e^{1-r}}{2(r+1)}\sum_{k=0}^{r}\frac{(k+1)(k+r+1)r^{r-k}}{2^{k}(r-k)!}={\rm lim}_{d\rightarrow+\infty}\frac{S}{2n^{n}}.$
+We have $S/2n^{n}<R_{\infty}e^{r(r+2)/(2d)}$.
+So $S<2n^{n}$ when $d\geq d(r)=\frac{r(r+2)}{2(-{\rm log}R_{\infty})}$.
+By (Case I), we can assume that $4\leq r\leq 10$.
+By the numerical analysis (we provide the following Wolfram Mathematica code), we can determine the choice of $d(r)$ for each $4\leq r\leq 10$.
+Moreover, by the beginning discussion of (Case II), we can assume $d(3)=68$.
+We list the results in Table 1.
+[…]
+[…]
+[…]
+Table 1. $d(r)$
+$r$ 3 4 5 6 7 8 9 10
+$d(r)$ 68 101 89 91 97 106 117 128
+(Case III: $d$ small and $r$ large) We use the integral formula (35) to estimate. We use
+$2^{-(n-d+1)}(x_{1}-z)^{n-d+1}=2^{-r}(2r+d-z)^{r}=r^{r}(1+\frac{d-z}{2r})^{r}\leq r^{r}e^{(d-z)/2}$
+and $n-z\leq n$ to get:
+$\frac{S}{2n^{n}}$ $=$ $\frac{2^{-(n-d+1)}n!}{2n^{n}(d-3)!(n-d+2)!}\int_{0}^{d}z^{d-3}(d-z)(x_{1}-z)^{n-d+1}(n-z){\rm d}z$
+$\leq$ $\frac{n(n-1)\cdots(n-d+3)}{(d-3)!}\cdot\frac{r^{r}\cdot n}{2n^{n}}\int_{0}^{d}z^{d-3}(d-z)e^{(d-z)/2}{\rm d}z$
+$=$ $\frac{n(n-1)\cdots(n-d+3)}{(d-3)!n^{d-2}}\frac{(n-d+1)^{r}}{2n^{r}}\int_{0}^{d}z^{d-3}(d-z)e^{(d-z)/2}{\rm d}z$
+$\leq$ $(1+\frac{d-1}{r})^{-r}\frac{1}{2(d-3)!}\int_{0}^{d}z^{d-3}(d-z)e^{(d-z)/2}{\rm d}z$
+where $\epsilon=\frac{d-1}{r}$ and
+$-r{\rm log}(1+\epsilon)=-r\sum_{k=0}^{+\infty}(\epsilon-\frac{\epsilon^{2}}{2}+\frac{\epsilon^{3}}{3}+\cdots)\leq-r\epsilon+r\epsilon^{2}/2.$
+Then
+$\frac{S}{2n^{n}}$ $\leq$ $e^{r\epsilon^{2}/2}e^{1-d/2}\frac{1}{2(d-3)!}\int_{0}^{d}z^{d-3}(d-z)e^{-z/2}{\rm d}z$
+We define
+$R_{\infty}(d)$ $=$ $e^{1-\frac{d}{2}}\frac{1}{2(d-3)!}\int_{0}^{d}z^{d-3}(d-z)e^{-z/2}{\rm d}z=\mathop{{\rm lim}}_{r\rightarrow+\infty}\frac{S}{2n^{n}}.$
+Therefore, we have $S/2n^{n}<e^{\frac{(d-1)^{2}}{2r}}R_{\infty}(d)$.
+So for $S<2n^{n}$, it suffices $r\geq r(d):=\frac{(d-1)^{2}}{2(-{\rm log}R_{\infty}(d))}$.
+By Case I, we can assume $3\leq d\leq 8$ and the corresponding $r(d)$ can be solved numerically (see the following Wolfram Mathematica code):
+[…]
+[…]
+[…]
+Table 2. $r(d)$
+$d$ 3 4 5 6 7 8
+$r(d)$ 16 20 27 34 43 52
+(Case IV: Both $d$ and r=n+1dr=n+1-d are small). We can numerically check (see the following Wolfram Mathematica code) that the estimate (31) holds for the following finitely many cases:
+  • $3\leq d\leq 8$ and $3\leq r\leq r(d)-1$.
+  • $4\leq r\leq 10$ and $3\leq d\leq d(r)-1$.
+Both $r(d)$ and $d(r)$ are given by Table 1 and 2.
+1x1 = 2 r + d
+2n = r + d - 1
+[…]
+[…]
+[…]
+[…]
+Appendix B A proof of the formula (32)
 We calculate the Duistermaat-Heckman measure via the localization method.
@@
 $\frac{1}{2\pi}\int_{{\mathbb{R}}}V_{0}(\theta)e^{-i\theta\xi}{\rm d}\theta$ $=$ $\frac{1}{4(n-1)!}((n-2)\xi^{n-1}-(n-1)\xi^{n-2})\mathrm{Sign}(-\xi)$
-=:absent:=:= : $\frac{f(\xi)}{2}\mathrm{Sign}(-\xi).$
+$=:$ $\frac{f(\xi)}{2}\mathrm{Sign}(-\xi).$
 The contribution from $F_{2}$ is given by:
@@
 $\frac{1}{2\pi}\int_{{\mathbb{R}}}V_{2}(\theta)e^{-i\theta\xi}{\rm d}\theta$ $=$ $\frac{1}{4(n-1)!}(-(n-2)(2-\xi)^{n-1}+(n-1)(2-\xi)^{n-2})\mathrm{Sign}(2-\xi)$
-=:absent:=:= : $\frac{g(\xi)}{2}\mathrm{Sign}(2-\xi).$
+$=:$ $\frac{g(\xi)}{2}\mathrm{Sign}(2-\xi).$
 The contribution from $F_{1}$ is given by:
-$V_{1}(\theta)$ ::=:= $\int_{Q^{n-4}}\frac{e^{(\omega+i\theta)}}{(\omega+i\theta)^{2}(\omega-i\theta)^{2}}=\int_{Q^{n-4}}\frac{e^{\omega}e^{i\theta}}{(\omega^{2}+\theta^{2})^{2}}$
+$V_{1}(\theta)$ $:=$ $\int_{Q^{n-4}}\frac{e^{(\omega+i\theta)}}{(\omega+i\theta)^{2}(\omega-i\theta)^{2}}=\int_{Q^{n-4}}\frac{e^{\omega}e^{i\theta}}{(\omega^{2}+\theta^{2})^{2}}$
 $=$ $\sum_{k=0}^{\lfloor\frac{n-4}{2}\rfloor}\frac{(k+1)}{(n-4-2k)!}\frac{e^{i\theta}}{(i\theta)^{2k+4}}.$
@@
 $\frac{1}{2\pi}\int_{{\mathbb{R}}}V_{1}(\theta)e^{-i\theta\xi}{\rm d}\theta$ $=$ $\sum_{k=0}^{\lfloor\frac{n-4}{2}\rfloor}\frac{(k+1)}{2(n-4-2k)!(2k+3)!}(1-\xi)^{2k+3}\mathrm{Sign}(1-\xi).$
-We claim that this is equal to $-\frac{f(\xi)+g(\xi)}{2}\mathrm{Sign}(1-\xi)$ so that the Fourier transform is $V(\theta)$ is equal to:
+We claim that this is equal to $\frac{f(\xi)+g(\xi)}{2}\mathrm{Sign}(1-\xi)$ so that the Fourier transform is $V(\theta)$ is equal to:
 $\rho(\xi)$ $=$ $\frac{f(\xi)}{2}\mathrm{Sign}(-\xi)+\frac{g(\xi)}{2}\mathrm{Sign}(2-\xi)-\frac{f(\xi)+g(\xi)}{2}\mathrm{Sign}(1-\xi)$